### v1

# From Poisson Observations to Fitted Negative Binomial Distribution

Abstract
The Kolmogorov-Smirnov (KS) test has been widely used for testing whether a random sample comes from a specific distribution, possibly with estimated parameters. If the data come from a Poisson distribution, however, one can hardly tell that they do not come from a negative binomial distribution by running a KS test, even with a large sample size. In this paper, we rigorously justify that the KS test statistic converges to zero almost surely, as the sample size goes to infinity. To prove this result, we demonstrate a notable finding that in this case the maximum likelihood estimates (MLE) for the parameters of the negative binomial distribution converge to infinity and one, respectively and almost surely. Our result highlights a potential limitation of the KS test, as well as other tests based on empirical distribution functions (EDF), in efficiently identifying the true underlying distribution. Our findings and justifications also underscore the importance of careful interpretation and further investigation when identifying the most appropriate distributions in practice.
Key words and phrases: Empirical distribution function, KS test, MLE, Negative binomial distribution, Poisson distribution
1 Introduction
The Kolmogorov-Smirnov (KS) test, introduced by Massey Jr (1951), has been widely used as a fundamental tool for assessing whether a given random sample $\{Y_{1},\ldots,Y_{n}\}$ originates from a continuous cumulative distribution function (CDF) $F_{\bm{\theta}}(y)$ characterized by specific model parameter(s) $\bm{\theta}$. The KS test statistic is defined as $D_{n}=\sup_{y}|F_{n}(y)-F_{\bm{\theta}}(y)|$, where $F_{n}(y)=n^{-1}\sum_{i=1}^{n}{\mathbf{1}}_{(-\infty,y]}(Y_{i})$ represents the empirical distribution function (EDF) of the sample. Dimitrova et al. (2020) extended the applicability of the KS test to encompass a broader range of distributions $F_{\bm{\theta}}(y)$, including discrete and mixed distributions with known $\bm{\theta}$. When $\bm{\theta}$ is unknown, Lilliefors (1967, 1969) suggested replacing $\bm{\theta}$ with its estimate $\hat{\bm{\theta}}$ when calculating $D_{n}$, along with adjusted $p$-values or thresholds for test statistics. Aldirawi et al. (2019) proposed a bootstrapped Monte Carlo estimator for calculating adjusted $p$-values under those circumstances. R packages, such as iZID (Wang et al., 2020) and AZIAD (Dousti Mousavi et al., 2023), have been developed for such kind of purposes.
Suppose we have two different families of probability distributions, ${\cal F}_{1}=\{F_{1}(y;\bm{\theta}_{1})\mid\bm{\theta}_{1}\in\bm{\Theta}_{1}\}$ and ${\cal F}_{2}=\{F_{2}(y;\bm{\theta}_{2})\mid\bm{\theta}_{2}\in\bm{\Theta}_{2}\}$, where $\bm{\Theta}_{1}$ and $\bm{\Theta}_{2}$ are the corresponding parameter spaces. Given a random sample $\{Y_{1},\ldots,Y_{n}\}$ from $F_{1}(y;\bm{\theta}_{1})\in{\cal F}_{1}$, can we always tell by a KS test that it does not come from ${\cal F}_{2}$? The answer is no if $F_{1}(y;\bm{\theta}_{1})=F_{2}(y;\bm{\theta}_{2})$ for some $\bm{\theta}_{2}$ and all $y$. For examples, Dousti Mousavi et al. (2023) pointed out that given a random sample from a zero-inflated distribution, such as a zero-inflated Poisson distribution (ZIP), the KS-test can hardly tell that the sample does not come from the corresponding zero-altered distribution, such as a zero-altered Poisson (ZAP, also known as Poisson hurdle or PH) distribution. This is because ZIP can be regarded as a special case of ZAP or PH.
In this paper, we consider a different type of phenomena, Poisson distribution and negative binomial (NB) distribution. Neither of them is a special case of the other, although it is well known that a Possible distribution can be approximated well by a negative binomial distribution. For example, as a direct conclusion of Theorem 5.13.1 in Fisz (1963), $F_{{\rm NB}(\nu,p(\nu))}(k)\rightarrow F_{\lambda}(k)$ as $\nu\rightarrow\infty$ for each $k$, where $p(\nu)=\nu/(\nu+\lambda)$, $F_{{\rm NB}(\nu,p)}$ is the CDF of negative binomial distribution with parameters $\nu>0$ and $p\in(0,1)$, and $F_{\lambda}$ is the CDF of Poisson distribution with parameter $\lambda>0$.
Suppose we have a random sample $\{Y_{1},\ldots,Y_{n}\}$ from a Poisson distribution and let $F_{n}(y)$ be the EDF. Pretending that the sample comes from a negative binomial distribution, we obtain the maximum likelihood estimate (MLE) $\hat{\bm{\theta}}=(\hat{\nu},\hat{p})^{T}$ and let $F_{\hat{\bm{\theta}}}(y)$ be the CDF of negative binomial distribution with parameters $\hat{\bm{\theta}}$. The question is whether we can use the KS test to tell that the underlying distribution is not negative binomial but Poisson.
To see a numerical answer, it is not difficult to simulate random samples from a Poisson distribution, for example, Poisson($\lambda=10$), with various sample sizes. For each sample, we calculate the KS test statistic $D_{n}=\sup_{y}|F_{n}(y)-F_{\hat{\bm{\theta}}}(y)|$ for negative binomial distributions using an algorithm described in Section 4.1 and list them in Table 1. There is a clear trend that $D_{n}$ vanishes as $n$ goes to infinity, which implies that it becomes harder and harder to detect the true distribution based on $D_{n}$
Table 1: KS test statistic for negative binomial distribution given random samples from Poisson($\lambda=10$) with various sample sizes $n$
n $50$ $500$ $5000$ $50000$ $500000$
$\bm{D_{n}=\sup_{y}|F_{n}(y)-F_{\hat{\bm{\theta}}}(y)|}$ 0.0555 0.0220 0.0032 0.0013 0.0004
To answer the question with theoretical justifications, however, it is much more difficult, because there is no simple formula for calculating the MLE for negative binomial distributions with both $\nu$ and $p$ as parameters. It is straightforward to figure out that the MLE of $p$ must satisfy $\hat{p}=\hat{\nu}/(\hat{\nu}+\bar{Y}_{n})$ (see, for example, Bliss and Fisher (1953) or Section 5.8.2 in Johnson et al. (2005)), where $\bar{Y}_{n}$ is the sample mean. The difficulty stays with finding the MLE of $\nu$. Anscombe (1950) conjectured that the likelihood equation of $\nu$ has a unique solution if and only if $n^{-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}>\bar{Y}_{n}$ Since then, a lot of efforts have been made on proving Anscombe’s conjecture (see Wang (1996), Bandara et al. (2019), and references therein). According to Bandara et al. (2019), the conjecture has been proved by Simonsen (1976, 1980). Nevertheless, the question still remains. Even if the condition $n^{-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}>\bar{Y}_{n}$ is satisfied, the MLE of $\nu$ can only be obtained by numerical algorithms, such as the Newton-Raphson method (Bandara et al., 2019). In our problem, since the data come from a Poisson distribution, the probability of $n^{-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}>\bar{Y}_{n}$ is only about $0.5$ or less (see Table 2).
Table 2: $P(n^{-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}>\bar{Y}_{n})$ estimated by $1000$ random samples of size $n$ from Poisson($\lambda$)
n
$\bm{\lambda}$ 50 500 5000 50000 500000
1 0.408 0.500 0.497 0.476 0.484
3 0.441 0.445 0.461 0.503 0.487
5 0.406 0.495 0.490 0.476 0.501
10 0.408 0.487 0.496 0.496 0.469
In this paper, we show in Theorem 3.4 that if $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson($\lambda$) with $\lambda>0$, then for any $\nu_{\rm max}>0$, with probability 1, the MLE of $\nu$ is $\nu_{\rm max}$ for a negative binomial distribution under the constraint $0<\nu\leq\nu_{\rm max}$, as $n$ goes to $\infty$. In Theorem 3.6, we further show that if we adjust the constraint to $0<\nu\leq\nu_{\rm max}(n)$, where $\lim_{n\rightarrow\infty}\nu_{\rm max}(n)=\infty$, then with probability 1, the KS test statistic $D_{n}\rightarrow 0$, as $n$ goes to infinity.
The remaining sections are organized as follows. In Section 2, we review and summarize the results relevant to the MLE for negative binomial distributions, and derive the efficient score function $g(\nu)$, which plays a key role throughout this paper. In Section 3, we obtain the asymptotic properties of the MLE and the KS test statistic for negative binomial distributions given a Poisson random sample. In Section 4, we propose an accurate and robust algorithm for obtaining the MLE for negative binomial distributions under general circumstances, and numerically confirm that the KS test statistic vanishes as the sample size goes to infinity. We conclude in Section 5.
2 Finding MLE for Negative Binomial Distribution
In this section, we consider finding the maximum likelihood estimates (MLE) of $(\nu,p)$ given $Y_{1},\ldots,Y_{n}$ iid $\sim$ NB$(\nu,p)$ under the constraint $\nu\in(0,\nu_{\rm max}]$ and $p\in[0,1]$, where $\nu_{\rm max}>0$ is pre-determined. It is typically the situation when finding the MLE numerically for negative binomial distributions (see, e.g., Dousti Mousavi et al. (2023)).
2.1 Likelihood equation
In this section, we review and summarize the results in the literature on finding the MLE for negative binomial parameters.
Following Gu (2013, Section 5.4.6), a negative binomial distribution with $\bm{\theta}=(\nu,p)^{T}\in(0,\infty)\times(0,1)$ has a probability mass function (PMF) given by
$f_{\bm{\theta}}(y)=\frac{\Gamma(\nu+y)}{\Gamma(\nu)\Gamma(y+1)}p^{\nu}(1-p)^{y},\>y=0,1,\ldots$
with mean $\nu(1-p)/p$ and variance $\nu(1-p)/p^{2}$. For alternative parameterizations, please see Section 5.1 in Johnson et al. (2005) and references therein. Then
$\log f_{\bm{\theta}}(y)=\log\Gamma(\nu+y)-\log\Gamma(\nu)-\log\Gamma(y+1)+\nu\log p+y\log(1-p)$
Taking partial derivatives, we have
$\frac{\partial\log f_{\bm{\theta}}(y)}{\partial\nu}$ $=$ $\Psi(\nu+y)-\Psi(\nu)+\log p$
$\frac{\partial\log f_{\bm{\theta}}(y)}{\partial p}$ $=$ $\frac{\nu}{p}-\frac{y}{1-p}$
where $\Psi(\cdot)=\Gamma^{\prime}(\cdot)/\Gamma(\cdot)$ is known as the digamma function (see, for example, Section 6.3 in Abramowitz and Stegun (1970)).
Suppose $Y_{1},\ldots,Y_{n}$ iid $\sim$ NB$(\nu,p)$. The likelihood $l(\bm{\theta})=\sum_{i=1}^{n}\log f_{\bm{\theta}}(Y_{i})$. Then
$\frac{\partial l(\bm{\theta})}{\partial p}=\frac{n\nu}{p}-\frac{\sum_{i=1}^{n}Y_{i}}{1-p}=\frac{n(\nu+\bar{Y}_{n})}{p(1-p)}\left(\frac{\nu}{\nu+\bar{Y}_{n}}-p\right)\left\{\begin{array}[]{ll}>0&\mbox{ if }p<\frac{\nu}{\nu+\bar{Y}_{n}}\\ =0&\mbox{ if }p=\frac{\nu}{\nu+\bar{Y}_{n}}\\ <0&\mbox{ if }p>\frac{\nu}{\nu+\bar{Y}_{n}}\end{array}\right.$
where $\bar{Y}_{n}=n^{-1}\sum_{i=1}^{n}Y_{i}$ That is, $\partial l(\bm{\theta})/\partial p=0$, known as the likelihood equation of $p$, implies $\hat{p}=\hat{\nu}/(\hat{\nu}+\bar{Y}_{n})$ for the MLE $(\hat{\nu},\hat{p})$.
Example 2.1.
A trivial case of all zerosA special case is $Y_{1}=\cdots=Y_{n}=0$, which is possible in practice especially with a moderate $n$. In this case, $\bar{Y}_{n}=0$ as well. It can be verified that $l(\bm{\theta})=n\nu\log p\leq 0$. To maximize $l(\bm{\theta})$ with $\nu\in(0,\infty)$ and $p\in[0,1]$, we must have $\hat{p}=1$ and $\hat{\nu}$ can be any positive number. normal-$\Box$
As a conclusion, regardless $\bar{Y}_{n}=0$ or $\bar{Y}_{n}>0$, the maximum likelihood estimate of $p$ given $\nu>0$ must take the form as follows
$\hat{p}=\hat{p}(\nu)=\frac{\nu}{\nu+\bar{Y}_{n}}$ (1)
(see, e.g., Bliss and Fisher (1953) or Section 5.8.2 in Johnson et al. (2005)).
To avoid the trivial case as in Example 2.1, from now on, we assume $\bar{Y}_{n}>0$, that is, $Y_{i}>0$ for at least one $i$.
Since the best $p$ must be the function $\hat{p}(\nu)$ of $\nu$, maximizing $l(\bm{\theta})=l(\nu,p)$ is equivalent to maximizing
$h(\nu)=l(\nu,\hat{p}(\nu))$ (2)
$=$ $\sum_{i=1}^{n}\log\Gamma(\nu+Y_{i})-\sum_{i=1}^{n}\log\Gamma(Y_{i}+1)-n\log\Gamma(\nu)$
$-\ n(\nu+\bar{Y}_{n})\log(\nu+\bar{Y}_{n})+n\nu\log\nu+n\bar{Y}_{n}\log\bar{Y}_{n}$
According to (1), $\partial l(\nu,\hat{p}(\nu))/\partial p=0$. Then
$h^{\prime}(\nu)$ $=$ $\frac{\partial l}{\partial\nu}(\nu,\hat{p}(\nu))+\frac{\partial l}{\partial p}(\nu,\hat{p}(\nu))\cdot\hat{p}^{\prime}(\nu)=\frac{\partial l}{\partial\nu}(\nu,\hat{p}(\nu))$
$=$ $\sum_{i=1}^{n}\left[\Psi(\nu+Y_{i})-\Psi(\nu)\right]-n\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right)$
$\stackrel{{\scriptstyle\triangle}}{{=}}$ $ng(\nu)$
where (see also Bliss and Fisher (1953))
$g(\nu)$ $=$ $\frac{1}{n}\sum_{i=1}^{n}\left[\Psi(\nu+Y_{i})-\Psi(\nu)\right]-\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right)$ (3)
$=$ $\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right)$ (4)
by using the fact $\Psi(x+1)-\Psi(x)=x^{-1}$ (see, for example, Formula 6.3.5 in Abramowitz and Stegun (1970)). Here $I$ in (4) is the range (or the collection of distinct values) of $\{Y_{1},\ldots,Y_{n}\}$, and $f_{y}=\#\{1\leq i\leq n\mid Y_{i}=y\}$ is the frequency of $y$ in $\{Y_{1},\ldots,Y_{n}\}$.
In the literature, $g(\nu)$, defined by (3) or (4), is known as the efficient score function of $\nu$ up to a constant $n$ (Bliss and Fisher, 1953). Essentially, $g(\nu)=0$ is the likelihood equation of $\nu$ for negative binomial distributions. If $g(\nu)=0$ has a unique solution $\hat{\nu}\in(0,\infty)$, then $\hat{\nu}$ must be the MLE of $\nu$.
In our notation, Anscombe (1950) conjectured that $g(\nu)=0$ has a unique solution if and only if $n^{-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}>\bar{Y}_{n}$ According to Bandara et al. (2019), Simonsen (1976, 1980) proved Anscombe’s conjecture and obtained the following theorem.
Theorem 2.1 (Simonsen (1976, 1980)).
Denote $M=\max\{Y_{1},\ldots,Y_{n}\}$.
  • (1)
    If $M=1$ or $M\geq 2$ and $n^{-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}\leq\bar{Y}_{n}$, then $g(\nu)=0$ has no solution.
  • (2)
    If $M\geq 2$ and $n^{-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}>\bar{Y}_{n}$, then $g(\nu)=0$ has a unique solution in $(0,\infty)$.
2.2 Efficient score function
According to Theorem 2.1, the efficient score function $g(\nu)$ does not have a solution in $(0,\infty)$ if $M=1$ or $M=2$ and $n^{-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}\leq\bar{Y}_{n}$. In this section, we show that if $g(\nu)=0$ does not have a solution in $(0,\infty)$, then the MLE $\hat{\nu}=\nu_{\rm max}$ under the constraint $\nu\in(0,\nu_{\rm max}]$.
Due to (3) or (4), the first-order derivative of $g(\nu)$ is
$g^{\prime}(\nu)$ $=$ $\frac{1}{n}\sum_{i=1}^{n}\Psi_{1}(\nu+Y_{i})-\Psi_{1}(\nu)+\frac{1}{\nu}-\frac{1}{\nu+\bar{Y}_{n}}$ (5)
$=$ $-\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\left[\frac{1}{\nu^{2}}+\cdots+\frac{1}{(\nu+y-1)^{2}}\right]+\frac{\bar{Y}_{n}}{\nu(\nu+\bar{Y}_{n})}$
where $\Psi_{1}(\cdot)=\Psi^{\prime}(\cdot)$ is known as the trigamma function (see, for example, Section 6.4 in Abramowitz and Stegun (1970)).
Lemma 2.1.
Suppose $\bar{Y}_{n}>0$. Then $\lim_{\nu\rightarrow 0^{+}}g(\nu)=\infty$, $\lim_{\nu\rightarrow\infty}g(\nu)=0$, $\lim_{\nu\rightarrow 0^{+}}g^{\prime}(\nu)=-\infty$, and $\lim_{\nu\rightarrow\infty}g^{\prime}(\nu)=0$.
All proofs are relegated to the Appendix.
Example 2.2.
A trivial case of constant observationsSuppose $Y_{1}=\cdots=Y_{n}=c\geq 1$. Then $\bar{Y}_{n}=c$, $I=\{c\}$, $f_{c}=n$, and thus
$g^{\prime}(\nu)$ $=$ $-\frac{1}{\nu^{2}}-\cdots-\frac{1}{(\nu+c-1)^{2}}+\frac{c}{\nu(\nu+c)}$
$<$ $-\frac{1}{\nu(\nu+1)}-\cdots-\frac{1}{(\nu+c-1)(\nu+c)}+\frac{c}{\nu(\nu+c)}$
$=$ $-\frac{1}{\nu}+\frac{1}{\nu+1}-\cdots-\frac{1}{v+c-1}+\frac{1}{\nu+c}+\frac{1}{\nu}-\frac{1}{\nu+c}$
$=$ 00 0
for all $\nu>0$. According to Lemma 2.1, $\lim_{\nu\rightarrow 0^{+}}g(\nu)=\infty$ and $\lim_{\nu\rightarrow\infty}g(\nu)=0$, we must have $g(\nu)>0$ for all $\nu>0$ since $g(\nu)$ is strictly decreasing for all $\nu>0$ in this case.
According to the definition of $g(\nu)$, it implies that $\partial l(\bm{\theta})/\partial\nu=ng(\nu)>0$ for all $\nu>0$. Then the MLE of $\nu$ must be the maximum value that $\nu$ is allowed to be.
That is, the MLE of $\nu$ under the constraint $\nu\in(0,\nu_{\rm max}]$ is $\hat{\nu}=\nu_{\rm max}$. normal-$\Box$
Example 2.3 provides a case when $g(\nu)=0$ does not have a solution. In practice, to find the MLE of $\nu$, instead of solving $g(\nu)=0$, we may minimize $g^{2}(\nu)$, which always has a solution (see Algorithm 1 in Section 4.1).
Theorem 2.2.
Suppose $\bar{Y}_{n}>0$. If $\min_{\nu\in(0,\nu_{\rm max}]}g^{2}(\nu)>0$, then the MLE of $\nu$ under the constraint $\nu\in(0,\nu_{\rm max}]$ is $\hat{\nu}=\nu_{\rm max}$.
Example 2.3.
Mean is greater than varianceIn this example, we simulate a random sample of $n=5,000$ from a Poisson distribution with $\lambda=5$ and then add a constant $A$ to each of the observations. This results in a data set whose sample mean is greater than the sample variance when $A$ is large enough.
Using a numerical algorithm (see Algorithm 1 in Section 4.1), we obtain the MLE of $\nu$ for $A=10,20,50,70,100$, respectively, with the constraint $\nu\in(0,10000]$. Each $\hat{\nu}$ found by the numerical algorithm is $10000$, which matches the conclusion of Theorem 2.2. normal-$\Box$
3 Negative Binomial Estimates for Poisson Sample
In this section, we derive the main results of this paper: (1) Given a random sample from a Poisson distribution, with probability 1, the MLE of $\nu$ is $\nu_{\rm max}$ for negative binomial distributions under the constraint $0<\nu\leq\nu_{\rm max}$, as $n$ goes to $\infty$ (see Theorem 3.4 in Section 3.1); and (2) if we adjust the constraint to be $0<\nu\leq\nu_{\rm max}(n)$, where $\lim_{n\rightarrow\infty}\nu_{\rm max}(n)=\infty$, then the KS test statistic $D_{n}$ for a negative binomial distribution vanishes with probability 1, as $n$ goes to infinity (see Theorem 3.6 in Section 3.2).
3.1 Negative binomial MLE given Poisson sample
Recall from Section 2.2 that the efficient score function $g(\nu)$, expressed in (3) or (4), plays an important role in finding the MLE of $\nu$ for a negative binomial distribution.
[…]
$G_{F}(\nu)$ $=$ $E\Psi(\nu+Y_{1})-\Psi(\nu)-\log\left(1+\frac{\mu}{\nu}\right)$ (6)
$=$ $\sum_{y=1}^{\infty}P(Y_{1}=y)\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\log\left(1+\frac{\mu}{\nu}\right)$ (7)
which is actually the limiting function of $g(\nu)$ as $n$ goes to infinity (see Lemma 3.1 and Theorem 3.1 below).
Lemma 3.1.
Given $\mu=E(Y_{1})<\infty$, $G_{F}(\nu)$ is a continuous function on $\nu\in(0,\infty)$. Furthermore,
$G_{F}(\nu)=\sum_{y=0}^{\infty}\frac{1-F(y)}{\nu+y}-\log\left(1+\frac{\mu}{\nu}\right)$ (8)
where $F(y)=P(Y_{1}\leq y)$ is the CDF of $Y_{1}$
Theorem 3.1.
[…]
Using Kolmogorov’s Strong Law of Large Numbers (see, for example, Theorem 7.5.1 in Resnick (1999)), we provide a rigorous proof of Theorem 3.1 in the Appendix. It is actually applicable for discrete distributions other than Poisson distributions as well, as long as the distribution is on nonnegative integers. For example, in Lemma 3.2, we show that $G_{F}(\nu)\equiv 0$ if $F$ is a negative binomial CDF.
Lemma 3.2.
Suppose $Y_{1},\ldots,Y_{n}$ are iid $\sim$ NB$(\nu,p)$ with $\nu>0$ and $10<p<$ 1. Then $G_{NB}(\nu)=0$, for all $\nu>0$.
Suppose $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson$(\lambda)$ with $\lambda>0$. For $y\in\{0,1,2,\ldots,\}$, we denote $f_{\lambda}(y)=P(Y_{1}=y)=\frac{\lambda^{y}}{y!}e^{-\lambda}$, $F_{\lambda}(y)=P(Y_{1}\leq y)$, and
$G_{\lambda}(\nu)$ $=$ $\sum_{y=1}^{\infty}\frac{\lambda^{y}}{y!}e^{-\lambda}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\log\left(1+\frac{\lambda}{\nu}\right)$ (9)
$=$ $\sum_{y=0}^{\infty}\frac{1-F_{\lambda}(y)}{\nu+y}-\log\left(1+\frac{\lambda}{\nu}\right)$
as the almost sure limit of $g(\nu)$ given a random sample of Poisson$(\lambda)$.
Given $\lambda>0$ and $\nu>0$, we let $p=\nu/(\nu+\lambda)\in(0,1)$. Then the mean of NB$(\nu,p)$ is $\nu(1-p)/p=\nu\lambda/\nu=\lambda$, the same as Poisson$(\lambda)$’s. For such an NB$(\nu,p)$, we denote $f_{NB}(y)$ and $F_{NB}(y)$ as the corresponding probability mass function (PMF) and cumulative distribution function (CDF), respectively. According to Lemma 3.2,
$G_{NB}(\nu)=\sum_{y=0}^{\infty}\frac{1-F_{NB}(y)}{\nu+y}-\log\left(1+\frac{\mu}{\nu}\right)=\sum_{y=0}^{\infty}\frac{1-F_{NB}(y)}{\nu+y}-\log\left(1+\frac{\lambda}{\nu}\right)=0$
Comparing $G_{NB}(\nu)$ with (9), if we can show that
$\sum_{y=0}^{\infty}\frac{1-F_{\lambda}(y)}{\nu+y}>\sum_{y=0}^{\infty}\frac{1-F_{NB}(y)}{\nu+y}$ (10)
then $G_{\lambda}(\nu)>G_{NB}(\nu)=0$, for each $\nu>0$.
To prove (10), we first denote the continuous version of $f_{\lambda}(y)$ and $f_{NB}(y)$ by
$f_{\lambda}(x)$ $=$ $\frac{\lambda^{x}}{\Gamma(x+1)}e^{-\lambda}$ (11)
$f_{NB}(x)$ $=$ $\frac{\Gamma(\nu+x)}{\Gamma(x+1)\Gamma(\nu)}p^{\nu}(1-p)^{x}=\frac{\Gamma(\nu+x)}{\Gamma(x+1)\Gamma(\nu)}\cdot\frac{\nu^{\nu}\lambda^{x}}{(\nu+\lambda)^{\nu+x}}$ (12)
[…]
Lemma 3.3.
We let $r(x)=\log f_{NB}(x)-\log f_{\lambda}(x)$, $x\in[0,\infty)$. Then its first derivative
$r^{\prime}(x)\left\{\begin{array}[]{cl}<0&\mbox{ if }0<x<x_{*}\\ =0&\mbox{ if }x=x_{*}\\ >0&\mbox{ if }x>x_{*}\end{array}\right.$
where $x_{*}=\Psi^{-1}\left(\log(\nu+\lambda)\right)-\nu>\lambda$. Furthermore, $r(0)>0$ and $\lim_{x\rightarrow\infty}r(x)=\infty$.
Lemma 3.3 tells us that $f_{NB}(x)<\log f_{\lambda}(x)$, for $0<x<x_{*}$, and $f_{NB}(x)>\log f_{\lambda}(x)$, for $x>x_{*}$ As the next step, we compare $F_{NB}(y)$ and $F_{\lambda}(y)$ in Theorem 3.2.
Theorem 3.2.
Given $\lambda>0$ and $\nu>0$, we let $p=\nu/(\nu+\lambda)\in(0,1)$, $d(y)=f_{NB}(y)-f_{\lambda}(y)$, and $D(y)=\sum_{k=0}^{y}d(k)=F_{NB}(y)-F_{\lambda}(y)$, $y\in\{0,1,2,\ldots\}$. Then there exist three positive integers $K_{1}$, $K_{*}$, and $K_{2}$, such that, $0<K_{1}<K_{*}<K_{2}<\infty$,
$d(y)\left\{\begin{array}[]{cl}>0&\mbox{ if }0\leq y<K_{1}\mbox{ or }y>K_{2}\\ \geq 0&\mbox{ if }y=K_{1}\mbox{ or }K_{2}\\ <0&\mbox{ if }K_{1}<y<K_{2}\end{array}\right.,\>\>\>D(y)\left\{\begin{array}[]{cl}>0&\mbox{ if }0\leq y<K_{*}\\ \geq 0&\mbox{ if }y=K_{*}\\ <0&\mbox{ if }y>K_{*}\end{array}\right.$
Furthermore, $D(y)$ strictly increases before $K_{1}$ and attains its maximum at $y=K_{1}$ with $D(K_{1})>0$, then strictly decreases between $K_{1}$ and $K_{2}$ and attains its minimum at $y=K_{2}$ with $D(K_{2})<0$, and then strictly increases after $K_{2}$ and approaches to 00 as $y$ goes to $\infty$.
Lemma 3.4.
Given $D(y)$ as defined in Theorem 3.2, we must have $\sum_{y=0}^{\infty}D(y)=0$.
Now we are ready to show (10) and thus $G_{\lambda}(\nu)>0$ for all $\nu>0$ in Theorem 3.3.
Theorem 3.3.
For any $\lambda>0$ and $\nu>0$, we must have $G_{\lambda}(\nu)>0$.
According to Theorem 3.1, given $Y_{1},\ldots,Y_{n}$ iid $\sim$ Poisson($\lambda$), we have $g(\nu)\rightarrow G_{\lambda}(\nu)$ almost surely, as $n$ goes to infinity. Now Theorem 3.3 guarantees that $G_{\lambda}(\nu)>0$ for all $\nu>0$. It is straightforward to derive $g(\nu)>0$ for large enough $n$ as in Lemma 3.5.
Lemma 3.5.
Suppose $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson($\lambda$) with $\lambda>0$. Then for each $\nu>0$, there exists a constant $c>0$, such that, $g(\nu)>c$ almost surely, as $n$ goes to $\infty$.
For each $\nu>0$, Lemma 3.5 guarantees $g(\nu)>c$ for some $c>0$ and large enough $n$. Nevertheless, it does not guarantee that $g(\nu)>0$ for all $\nu>0$ and large enough $n$. To show that $g(\nu)>0$ for all $\nu\in(0,\nu_{\rm max}]$ and large enough $n$, we split $(0,\nu_{\rm max}]$ into $(0,\nu_{0}]$ and $(\nu_{0},\nu_{\rm max}]$, where $\nu_{0}$ is some positive number that is close enough to 00. Then we bound $g(\nu)$ from below for large enough $n$.
First, we find a lower bound for $g(\nu)$ on $\nu\in(0,\nu_{0}]$. Actually, according to (4), if $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson($\lambda$) with $\lambda>0$, then
$g(\nu)$
$=$ $\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right)$
$=$ $\frac{1}{n}\cdot\frac{\sum_{y=1}^{\infty}f_{y}}{\nu}+\frac{1}{n}\sum_{y\in I\setminus\{0,1\}}f_{y}\left(\frac{1}{\nu+1}+\cdots+\frac{1}{\nu+y-1}\right)-\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right)$
$\geq$ $\frac{1}{\nu}\cdot\frac{\#\{i:Y_{i}>0\}}{n}-\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right)$
$\rightarrow$ $\frac{1-e^{-\lambda}}{\nu}-\log\left(1+\frac{\lambda}{\nu}\right)$
almost surely, as $n$ goes to $\infty$.
Lemma 3.6.
Given $\lambda>0$ and $c>0$, there exists a $\nu_{0}>0$, such that,
$h(\nu,a,b):=\frac{a}{\nu}-\log\left(1+\frac{b}{\nu}\right)\geq c$
for all $0<\nu\leq\nu_{0}$, $a\in\left[\frac{1}{2}(1-e^{-\lambda}),\frac{3}{2}(1-e^{-\lambda})\right]$, and $b\in\left[\frac{\lambda}{2},\frac{3\lambda}{2}\right]$.
Note that $\nu_{0}$ in Lemma 3.6 does not depend on $Y_{i}$’s. Since #$\frac{\#\{i:Y_{i}>0\}}{n}\rightarrow 1-e^{-\lambda}$ and $\bar{Y}_{n}\rightarrow\lambda$ almost surely, as $n$ goes to $\infty$, there exists an event
$A=\{\omega\in\Omega\mid\lim_{n\rightarrow\infty}\frac{\#\{i:Y_{i}(\omega)>0\}}{n}=1-e^{-\lambda}\mbox{ and }\lim_{n\rightarrow\infty}\bar{Y}_{n}(\omega)=\lambda\}$ (13)
with probability $1$, where $\Omega$ is the sample space. Given $\omega\in A$, for large enough $n$, #$\frac{\#\{i:Y_{i}>0\}}{n}\in\left[\frac{1}{2}(1-e^{-\lambda}),\frac{3}{2}(1-e^{-\lambda})\right]$, $\bar{Y}_{n}\in\left[\frac{\lambda}{2},\frac{3\lambda}{2}\right]$, and thus $g(\nu)(\omega)\geq c$ for all $\nu\in(0,\nu_{0}]$. That is, with probability $1$, we obtain a common lower bound $c$ of $g(\nu)$, for all $\nu\in(0,\nu_{0}]$, with large enough $n$.
Secondly, we construct a lower bound for $g(\nu)$, $\nu\in(\nu_{0},\nu_{\rm max}]$, with large enough $n$, which relies on Lemma 3.7.
Lemma 3.7.
For any given $\nu_{0}>0$, there exists an $M>0$, which does not depend on the data, such that $|g^{\prime}(\nu)|\leq M$ for all $\nu>\nu_{0}$
With the aids of Theorem 3.3, Lemmas 3.5, 3.6, and 3.7, we are ready to derive the first main result of this paper as described in Theorem 3.4.
Theorem 3.4.
Suppose $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson($\lambda$) with $\lambda>0$, then for any $\nu_{\rm max}>0$, with probability 1, the MLE of $(\nu,p)$ for a negative binomial distribution under the constraint $0<\nu\leq\nu_{\rm max}$ is $\left(\nu_{\rm max},\frac{\nu_{\rm max}}{\nu_{\rm max}+\bar{Y}_{n}}\right)$ as $n$ goes to $\infty$, where $\bar{Y}_{n}=n^{-1}\sum_{i=1}^{n}Y_{i}$
In practice, we may allow $\nu_{\rm max}$ to increase along with $n$. As a direct conclusion of the proof of Theorem 3.4, we have the following theorem.
Theorem 3.5.
Suppose $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson($\lambda$) with $\lambda>0$. Let $\nu_{\rm max}(n)$ be a sequence of positive real numbers such that $\nu_{\rm max}(n)\rightarrow\infty$ as $n$ goes to $\infty$. Denote $(\hat{\nu}_{n},\hat{p}_{n})$ be the MLE of $(\nu,p)$ for a negative binomial distribution under the constraint $0<\nu\leq\nu_{\rm max}(n)$. Then with probability $1$, $\hat{\nu}_{n}\rightarrow\infty$, as $n$ goes to $\infty$.
3.2 KS test statistic
Given $Y_{1},\ldots,Y_{n}$ iid $\sim{\rm Poisson}(\lambda)$, we denote $\hat{\bm{\theta}}=(\hat{\nu}_{n},\hat{p}_{n})^{T}$ be the MLE of $\bm{\theta}=(\nu,p)^{T}$ for a negative binomial distribution under the constraint $0<\nu\leq\nu_{\rm max}(n)$, where $\lim_{n\rightarrow\infty}\nu_{\rm max}(n)=\infty$. With the aid of Theorem 3.5, we obtain the following theorem for the limiting behavior of the KS test statistic $D_{n}$ for negative binomial distributions.
Theorem 3.6.
Given $Y_{1},\ldots,Y_{n}$ iid $\sim{\rm Poisson}(\lambda)$ with $\lambda>0$, we denote $D_{n}=\sup_{y}|F_{n}(y)-F_{\hat{\bm{\theta}}}(y)|$ as the KS test statistic for negative binomial distribution, where $F_{n}(y)$ is the EDF of the data, and $F_{\hat{\bm{\theta}}}(y)$ is the CDF of negative binomial distribution with the MLE $\hat{\bm{\theta}}$ obtained under the constraint $0<\nu\leq\nu_{\rm max}(n)$, where $\lim_{n\rightarrow\infty}\nu_{\rm max}(n)=\infty$. Then $D_{n}\rightarrow 0$ almost surely, as $n\rightarrow\infty$.
Theorem 3.6 theoretically justifies the phenomenon observed in Table 1, which implies that a KS test can hardly distinguish a Poisson distribution from negative binomial distributions, since its test statistic vanishes with more and more observations.
4 Numerical Studies
In this section, we use simulation studies to numerically confirm our findings.
4.1 A new algorithm for negative binomial distributions
In this section, we propose a new algorithm to find the MLE of negative binomial distributions more accurately, by incorporating our theoretical results in Section 2.2.
First of all, instead of the moment estimate $\frac{\bar{Y}_{n}^{2}}{S^{2}-\bar{Y}_{n}}$ widely used for $\nu$ (Fisher, 1941; Bliss and Fisher, 1953; Johnson et al., 2005), we propose the following estimate of $\nu$ as an initial value
$\hat{\nu}^{(0)}=\min\left\{\nu_{\rm max},\>\frac{\bar{Y}_{n}^{2}}{\max\{\varepsilon,\>S^{2}-\bar{Y}_{n}\}}\right\}$ (14)
where $\nu_{\rm max}$ is a predetermined upper bound for $\nu$ (for example, $\nu_{\rm max}=10^{4}$), $\varepsilon$ is a predetermined small positive value (for example, $\varepsilon=10^{-3}$), and $S^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}$ is the sample variance. Note that our initial value (14) allowing $S^{2}-\bar{Y}_{n}\leq 0$, which is more robust in practice.
Due to Theorem 2.1, the likelihood equation $g(\nu)=0$ may not have a solution. Instead of numerically finding a solution for $g(\nu)=0$ as the MLE, we aim to minimize $g^{2}(\nu)$. According to Theorem 2.2, the MLE $\hat{\nu}=\nu_{\rm max}$ if $\min_{\nu\in(0,\nu_{\rm max}]}g^{2}(\nu)>0$.
We describe the proposed algorithm as Algorithm 1.
Algorithm 1 Finding MLE for Negative Binomial Distribution
Input: Data $Y_{1},\ldots,Y_{n}$, predetermined values $\nu_{\rm max}>0$, $\varepsilon>0$
Output: MLE $(\hat{\nu},\hat{p})$ for NB($\nu,p$)
1:Calculate sample mean $\bar{Y}_{n}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}$ and sample variance $S^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}$;
2:If $\bar{Y}_{n}=0$, go to Step $5$ with $\hat{\nu}=1$ and $\hat{p}=1$; otherwise calculate the initial value $\hat{\nu}^{(0)}$ of $\nu$ according to (14);
3:Obtain $\nu_{*}$ by minimizing $g^{2}(\nu)$ (see (4)) with $g^{\prime}(\nu)$ (see (5)) and the starting value $\hat{\nu}^{(0)}$. This can be implemented, for example, by using the R function optim; If $h(\nu_{\rm max})>h(\nu_{*})$ (see (2)), we let $\hat{\nu}=\nu_{\rm max}$; otherwise, $\hat{\nu}=\nu_{*}$;
4:Calculate $\hat{p}=\frac{\hat{\nu}}{\hat{\nu}+\bar{Y}_{n}}$;
5:Report $\hat{\nu}$ and $\hat{p}$.
For illustration purpose, we compare our algorithm with the R package AZIAD (Dousti Mousavi et al., 2023), one of the latest R packages for similar purposes. For a comprehensive simulation study, we consider 15 pairs $(\nu,p)$ of parameters for negative binomial distributions (see Tables 3 and 4). For each pair of $(\nu,p)$, we simulate $B=100$ random samples of size $n=100$ or $1000$, respectively. For each random sample, we use Algorithm 1 and AZIAD (version 0.0.2) to calculate the MLE of $(\nu,p)$, respectively, and use the Relative Error (RE, see Section 3.5 of Abramowitz and Stegun (1970)) to measure their levels of accuracy. The RE for an estimator $\hat{\beta}$ with respect to its true value $\beta\neq 0$ is defined by $|\hat{\beta}-\beta|/|\beta|$.
Table 3: Average Relative Error (RE%) of $\hat{\nu}$ by Algorithm 1 and AZIAD (in brackets)
REof $\hat{\nu}$ with $n=100$
$\bm{\nu}$ p = 0.00001 p = 0.0001 p = 0.001 p = 0.01 p = 0.1
1 10.24 (5643) 11.54 (519.8) 11.41 (10.19) 10.39 (10.39) 13.20 (13.20)
10 12.03 (9396) 11.73 (854.5) 11.69 (8.650) 12.29 (11.91) 13.22 (13.07)
100 11.06 (9838) 11.42 (896.3) 11.15 (7.650) 11.92 (12.11) 12.93 (12.94)
REof $\hat{\nu}$ with $n=1000$
$\bm{\nu}$ p = 0.00001 p = 0.0001 p = 0.001 p = 0.01 p = 0.1
1 2.91 (5601) 3.13 (514.6) 3.08 (2.83) 3.21 (3.21) 3.73 (3.73)
10 3.82 (9427) 3.62 (856.7) 3.59 (2.20) 3.47 (3.45) 4.30 (4.07)
100 3.54 (9858) 3.48 (896.7) 3.46 (1.85) 3.37 (3.39) 3.47 (3.48)
Table 4: Average Relative Error (RE%) of $\hat{p}$ by Algorithm 1 and AZIAD (in brackets)
REof $\hat{p}$ with $n=100$
$\bm{\nu}$ p = 0.00001 p = 0.0001 p = 0.001 p = 0.01 p = 0.1
1 13.24 (9900) 13.98 (900) 13.67 (9.24) 12.09 (12.09) 12.52 (12.52)
10 12.42 (9900) 12.06 (900) 12.09 (8.08) 12.26 (12.03) 11.61 (11.75)
100 11.22 (9900) 11.58 (900) 11.24 (7.37) 11.89 (12.05) 11.58 (11.59)
REof $\hat{p}$ with $n=1000$
$\bm{\nu}$ p = 0.00001 p = 0.0001 p = 0.001 p = 0.01 p = 0.1
1 4.02 (9900) 4.03 (900) 3.77 (2.28) 3.95 (3.95) 3.89 (3.89)
10 3.81 (9900) 3.71 (900) 3.78 (1.85) 3.70 (3.55) 4.01 (3.84)
100 3.55 (9900) 3.49 (900) 3.46 (1.73) 3.34 (3.33) 3.11 (3.10)
In Tables 3 and 4, the average AEs over $B=100$ random samples are displayed for each pair of $(\nu,p)$. When $p$ is not too small, such as $p=0.1$ and $p=0.01$, the average RE of $\hat{\nu}$ (Tables 3) and $\hat{p}$ (Table 4) obtained via Algorithm 1 and AZIAD are comparable; for $p=0.001$, AZIAD’s estimates (in brackets) are slightly better; for $p$ as small as $0.0001$ or $0.00001$, AZIAD’s estimates are not satisfactory, while our estimates still work reasonably well. Overall, we recommend Algoirthm 1, especially when $p$ is small.
In addition, we simulate $B=100$ random samples from Poisson($\lambda$) with $\lambda=1,3,5,10$ and different sample sizes $n$. For each sample, we use Algorithm 1 and AZIAD, respectively, to maximize the likelihood of negative binomial distributions. The average ratios of the maximum likelihood of Algorithm 1’s against AZIAD’s are provided in Table 5. Almost all average ratios are larger than $1$, which indicates that Algorithm 1 outperforms AZIAD in maximizing the likelihood of negative binomial distributions. The average ratio becomes larger with increased $\lambda$ and $n$.
Table 5: Average Ratio of Maximum Likelihood Obtained by Algorithm 1 against AZIAD on $100$ Random Samples from Poisson($\lambda$) of Size $n$
n
$\bm{\lambda}$ 50 500 5000 50000
1 1.00 1.02 1.17 3.78
3 1.01 1.05 1.46 35.2
5 1.01 1.09 5.48 351.5
10 1.03 1.16 3.55 267024
4.2 KS test statistic
To numerically confirm the performance of KS test statistic $D_{n}$, we simulate random samples from Poisson($\lambda$) with various sample sizes $n$. For each random sample, we fit an NB$(\nu,p)$ model and obtain its MLE $(\hat{\nu},\hat{p})$ under the constraints $100000<\nu\leq\nu_{\rm max}=$ 10000 using Algorithm 1, as well as the KS test statistic $D_{n}$ for testing whether the data come from a negative binomial distribution. The results are displayed in Table 6. As the sample size $n$ increases, $\hat{\nu}$ attains $\nu_{\rm max}$ for $n=50000$ or larger, $\hat{p}$ converges to $\nu_{\rm max}/(\nu_{\rm max}+\lambda)$, and $D_{n}$ decreases to zero, which confirm our conclusions in Section 3.
Table 6: MLE and KS Test Statistic for NB$(\nu,p)$ Using Algorithm 1 under $0<\nu\leq 10000$ Given Random Samples of Size $n$ from Poisson($\lambda$)
$\bm{\hat{\nu}}$
$\bm{\lambda}$ n = 50 n = 500 n = 5000 n = 50000 n = 500000
1 10000 54.805 69.710 10000 10000
3 13.057 54.635 91.888 10000 10000
5 13.175 51.383 133.32 10000 10000
10 151.39 10000 2145.2 10000 10000
$\bm{\hat{p}}$
$\bm{\lambda}$ n = 50 n = 500 n = 5000 n = 50000 n = 500000
1 0.9999 0.9822 0.9858 0.9999 0.9999
3 0.8152 0.9485 0.9684 0.9997 0.9997
5 0.7273 0.9122 0.9638 0.9995 0.9995
10 0.9325 0.9990 0.9953 0.9990 0.9990
$\bm{D_{n}}$
$\bm{\lambda}$ n = 50 n = 500 n = 5000 n = 50000 n = 500000
1 0.0215 0.0114 0.0060 0.0009 0.0002
3 0.0671 0.0130 0.0095 0.0013 0.0010
5 0.0937 0.0147 0.0066 0.0009 0.0007
10 0.0555 0.0220 0.0032 0.0013 0.0004
5 Discussion
In this paper, we theoretically verify the phenomenon that given a random sample from a Poisson distribution, the KS test statistic $D_{n}$, for testing whether the data come from a negative binomial distribution, is so small that one can hardly tell that the true distribution is not negative binomial but Poisson (Dousti Mousavi et al., 2023). According to our results, $D_{n}\rightarrow 0$ almost surely, as $n$ goes to infinity. That is saying, with larger and larger sample size, the situation becomes worse and worse.
[…]
Since $\hat{\omega}(n)^{2}\leq\sup_{y}|F_{n}(y)-F_{\hat{\bm{\theta}}}(y)|^{2}=D_{n}^{2}$, then $D_{n}\rightarrow 0$ almost surely implies $\hat{\omega}(n)^{2}\rightarrow 0$ almost surely. We encounter the same issue as with the KS test statistic, which needs further investigation as well.
Appendix: Proofs
Proof of Lemma 2.1: Since $\bar{Y}_{n}>0$, then there exists a $y\in I\setminus\{0\}$, such that, $f_{y}>0$. Then it can be verified that
$\lim_{\nu\rightarrow 0^{+}}\frac{\frac{1}{n\nu}\sum_{y\in I\setminus\{0\}}f_{y}}{g(\nu)}=1$
Since $\lim_{\nu\rightarrow 0^{+}}\frac{1}{n\nu}\sum_{y\in I\setminus\{0\}}f_{y}=\infty$, then $\lim_{\nu\rightarrow 0^{+}}g(\nu)=\infty$. Similarly, since
$\lim_{\nu\rightarrow 0^{+}}\frac{\frac{1}{n\nu^{2}}\sum_{y\in I\setminus\{0\}}f_{y}}{g^{\prime}(\nu)}=-1$
and $\lim_{\nu\rightarrow 0^{+}}\frac{1}{n\nu^{2}}\sum_{y\in I\setminus\{0\}}f_{y}=\infty$, then $\lim_{\nu\rightarrow 0^{+}}g^{\prime}(\nu)=-\infty$.
Note that $\lim_{\nu\rightarrow\infty}g(\nu)=0$ and $\lim_{\nu\rightarrow\infty}g^{\prime}(\nu)=0$ are straightforward. $\Box$
Proof of Theorem 2.2: Let $\nu_{*}={\rm argmin}_{\nu\in(0,\nu_{\rm max}]}g^{2}(\nu)$, then $g(\nu_{*})\neq 0$. According to Lemma 2.1, $\lim_{\nu\rightarrow 0^{+}}g(\nu)=\infty$. With the constraint $\nu\in(0,\nu_{\rm max}]$, we must have $\nu_{*}>0$.
We claim that $g(\nu_{*})>0$. Actually, if $g(\nu_{*})<0$, then there exists a $\nu_{**}\in(0,\nu_{*})$, such that, $g(\nu_{**})=0$, which violates $\min_{\nu\in(0,\nu_{\rm max}]}g^{2}(\nu)>0$.
The fact $g(\nu_{*})>0$ implies $g(\nu)\geq g(\nu_{*})>0$ for all $\nu\in(0,\nu_{\rm max}]$. That is, $h(\nu)$ increases all the time and thus the MLE of $\nu$ is $\nu_{\rm max}$. $\Box$
Proof of Lemma 3.1: For any $\nu_{0}>0$, once $\nu>\nu_{0}/2$,
$\left|\sum_{y=1}^{\infty}P(Y_{1}=y)\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\sum_{y=1}^{\infty}P(Y_{1}=y)\left(\frac{1}{\nu_{0}}+\cdots+\frac{1}{\nu_{0}+y-1}\right)\right|$
$\leq$ $\sum_{y=1}^{\infty}P(Y_{1}=y)\left[\left|\frac{1}{\nu}-\frac{1}{\nu_{0}}\right|+\cdots+\left|\frac{1}{\nu+y-1}-\frac{1}{\nu_{0}+y-1}\right|\right]$
$=$ $\sum_{y=1}^{\infty}P(Y_{1}=y)\left[\frac{|\nu-\nu_{0}|}{\nu\nu_{0}}+\cdots+\frac{|\nu-\nu_{0}|}{(\nu+y-1)(\nu_{0}+y-1)}\right]$
$\leq$ $\sum_{y=1}^{\infty}P(Y_{1}=y)\cdot y\cdot\frac{|\nu-\nu_{0}|}{\nu_{0}^{2}/2}$
$=$ $\frac{2E(Y_{1})}{\nu_{0}^{2}}\cdot|\nu-\nu_{0}|$
$\rightarrow$ 00 0
as $\nu$ goes to $\nu_{0}$ Since $\log(1+\mu/\nu)$ is a continuous function of $\nu$, $G_{F}(\nu)$ is a continuous function of $\nu$ on $(0,\infty)$, according to (7).
Furthermore,
$G_{F}(\nu)$ $=$ $\sum_{y=1}^{\infty}P(Y_{1}=y)\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\log\left(1+\frac{\mu}{\nu}\right)$
$=$ $\sum_{y=1}^{\infty}P(Y_{1}=y)\sum_{k=0}^{y-1}\frac{1}{\nu+k}-\log\left(1+\frac{\mu}{\nu}\right)$
$=$ $\sum_{k=0}^{\infty}\sum_{y=k+1}^{\infty}\frac{P(Y_{1}=y)}{\nu+k}-\log\left(1+\frac{\mu}{\nu}\right)$
$=$ $\sum_{k=0}^{\infty}\frac{1-P(Y_{1}\leq k)}{\nu+k}-\log\left(1+\frac{\mu}{\nu}\right)$
$=$ $\sum_{k=0}^{\infty}\frac{1-F(k)}{\nu+k}-\log\left(1+\frac{\mu}{\nu}\right)$
Thus the conclusion follows. $\Box$
Proof of Theorem 3.1: According to Kolmogorov’s Strong Law of Large Numbers (see, for example, Theorem 7.5.1 in Resnick (1999)), since $E(|Y_{1}|)=E(Y_{1})=\mu<\infty$, then $\bar{Y}_{n}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}\rightarrow\mu$ almost surely, as $n\rightarrow\infty$.
Similarly, for each $y\in\{0,1,2,\ldots\}$, $E(|\mathds{1}_{\{Y_{1}=y\}}|)=P(Y_{1}=y)\leq 1<\infty$. Then
$\frac{f_{y}}{n}$ $=$ #$\frac{\#\{1\leq i\leq n\mid Y_{i}=y\}}{n}\quad=\quad\frac{1}{n}\sum_{i=1}^{n}\mathds{1}_{\{Y_{i}=y\}}$
$\longrightarrow$ $E\mathds{1}_{\{Y_{1}=y\}}=P(Y_{1}=y)$
almost surely as $n\rightarrow\infty$. Furthermore, for each integer $M>0$, $E(|Y_{1}\mathds{1}_{\{Y_{1}>M\}}|)=E(Y_{1}\mathds{1}_{\{Y_{1}>M\}})\leq E(Y_{1})=\mu<\infty$. Then
$\sum_{y=M+1}^{\infty}\frac{yf_{y}}{n}$ $=$ $\sum_{y=M+1}^{\infty}\frac{\#\{1\leq i\leq n\mid Y_{i}=y\}\cdot y}{n}\quad=\quad\frac{1}{n}\sum_{i=1}^{n}Y_{i}\mathds{1}_{\{Y_{i}>M\}}$
$\longrightarrow$ $E(Y_{1}\mathds{1}_{\{Y_{1}>M\}})$
almost surely as $n\rightarrow\infty$. Then there exists an event $A$ with probability $1$, such that, for each $\omega\in A$,
$\frac{f_{y}(\omega)}{n}=\frac{\#\{1\leq i\leq n\mid Y_{i}(\omega)=y\}}{n}\longrightarrow P(Y_{1}=y)$ (15)
for each $y\in\{0,1,2,\ldots,\}$, $\bar{Y}_{n}(\omega)\rightarrow\mu$ and thus $\log\left(1+\frac{\bar{Y}_{n}(\omega)}{\nu}\right)\rightarrow\log\left(1+\frac{\mu}{\nu}\right)$, and
$\sum_{y=M+1}^{\infty}\frac{yf_{y}(\omega)}{n}\rightarrow E(Y_{1}\mathds{1}_{\{Y_{1}>M\}})$ (16)
for each $M>0$ as the convergence of a sequence of real numbers.
Now we only need to show that
$\sum_{y=1}^{\infty}\frac{f_{y}(\omega)}{n}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)\longrightarrow\sum_{y=1}^{\infty}P(Y_{1}=y)\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)$ (17)
for each $\omega\in A$. That is, we want to show that, for any $\epsilon>0$, there exists an $N$, such that,
$\left|\sum_{y=1}^{\infty}\frac{f_{y}(\omega)}{n}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\sum_{y=1}^{\infty}P(Y_{1}=y)\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)\right|<\epsilon$ (18)
for all $n>N$.
First, there exists an $M>0$, such that,
$\sum_{y=M+1}^{\infty}P(Y_{1}=y)\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)<\frac{\epsilon}{3}$ (19)
and
$E(Y_{1}\mathds{1}_{\{Y_{1}>M\}})<\frac{\nu\epsilon}{6}$
Then for each $y=1,\ldots,M$, there exists an $N_{y}>0$, such that,
$\left|\frac{f_{y}(\omega)}{n}-P(Y_{1}=y)\right|<\frac{\epsilon}{3}\cdot\frac{\nu}{M}\cdot\frac{1}{2^{y}}$
as long as $n>N_{y}$. Then for each $n>\max\{N_{1},\ldots,N_{M}\}$,
$\left|\sum_{y=1}^{M}\frac{f_{y}(\omega)}{n}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\sum_{y=1}^{M}P(Y_{1}=y)\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)\right|$ (20)
$\leq$ $\sum_{y=1}^{M}\left|\frac{f_{y}(\omega)}{n}-P(Y_{1}=y)\right|\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)$
$<$ $\sum_{y=1}^{M}\frac{\epsilon}{3}\cdot\frac{\nu}{M}\cdot\frac{1}{2^{y}}\cdot\frac{M}{\nu}$
$<$ $\frac{\epsilon}{3}$
We still need to show that
$\sum_{y=M+1}^{\infty}\frac{f_{y}(\omega)}{n}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)<\frac{\epsilon}{3}$
for large enough $n$. Actually, there exists an $N_{\nu}>0$ such that
$\left|\sum_{y=M+1}^{\infty}\frac{yf_{y}(\omega)}{n}-E(Y_{1}\mathds{1}_{\{Y_{1}>M\}})\right|<\frac{\nu\epsilon}{6}$
for all $n>N_{\nu}$. Then
$\sum_{y=M+1}^{\infty}\frac{f_{y}(\omega)}{n}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)$ (21)
$<$ $\sum_{y=M+1}^{\infty}\frac{f_{y}(\omega)}{n}\frac{y}{\nu}$
$<$ $\frac{1}{\nu}\left(E\left(Y_{1}{\mathbf{1}}_{Y_{1}>M}\right)+\frac{\nu\epsilon}{6}\right)$
$<$ $\frac{1}{\nu}\left(\frac{\nu\epsilon}{6}+\frac{\nu\epsilon}{6}\right)$
$=$ $\frac{\epsilon}{3}$
Combining (19), (20), and (21), we obtain that (18) is true for all $n>\max\{N_{1},\ldots,N_{M},$ $N_{\nu}\}$. That is, (17) is true for all $\omega\in A$. Combining with $\log\left(1+\frac{\bar{Y}_{n}(\omega)}{\nu}\right)\rightarrow\log\left(1+\frac{\mu}{\nu}\right)$ for all $\omega\in A$, we obtain $g(\nu)\rightarrow G_{F}(\nu)$ almost surely. $\Box$
Proof of Lemma 3.2: According to, for example, Example 3 in Aldirawi and Yang (2022), $E\Psi(\nu+Y_{1})=\Psi(\nu)-\log(p)$, if $Y_{1}\sim{\rm NB}(\nu,p)$ (please note that the parameters in their paper are different from ours, with $p$ be replaced with $1-p$). Then according to (6),
$G_{NB}(\nu)$ $=$ $E\Psi(\nu+Y_{1})-\Psi(\nu)-\log\left(1+\frac{\mu}{\nu}\right)$
$=$ $\Psi(\nu)-\log(p)-\Psi(\nu)-\log\left(1+\frac{1-p}{p}\right)$
$=$ 00 0
where $\mu=\nu(1-p)/p$ for NB$(\nu,p)$. In other words, $G_{NB}(\nu)=\sum_{y=0}^{\infty}\frac{1-F_{NB}(y)}{\nu+y}-\log\left(1+\frac{\mu}{\nu}\right)=0$, where $F_{NB}(y)=P(Y_{1}\leq y)$ with $Y_{1}\sim{\rm NB}(\nu,p)$. $\Box$
Proof of Lemma 3.3: Plugging in (11) and (12), we have for $x>0$,
$r(x)$ $=$ $\log\Gamma(\nu+x)-x\log(\nu+\lambda)-\log\Gamma(\nu)+\lambda-\nu\log\left(1+\frac{\lambda}{\nu}\right)$
$r^{\prime}(x)$ $=$ $\Psi(\nu+x)-\log(\nu+\lambda)$
According to, for example, Abramowitz and Stegun (1970) or Alzer and Jameson (2017), $\Psi(x)$ is strictly increasing and differentiable on $(0,\infty)$. Therefore, $r^{\prime}(x)$ is strictly increasing and differentiable on $(0,\infty)$.
According to Alzer (1997), $\log x-\frac{1}{x}<\Psi(x)<\log x-\frac{1}{2x}$ for all $x>0$. Then for $0<x<\lambda$, $r^{\prime}(x)<\Psi(\nu+\lambda)-\log(\nu+\lambda)<0$, and $\lim_{x\rightarrow\infty}r^{\prime}(x)=\infty$. The conclusion about $r^{\prime}(x)$ follows, with $x_{*}=\Psi^{-1}\left(\log(\nu+\lambda)\right)-\nu$ well defined and $x_{*}>\lambda$.
Since $x>\log(1+x)$ for all $x>0$, then $r(0)=\lambda-\nu\log\left(1+\frac{\lambda}{\nu}\right)>\lambda-\nu\cdot\frac{\lambda}{\nu}=0$.
On the other hand, according to Lemma 1 of Minc and Sathre (1964) or Alzer (1997),
$0<\log\Gamma(x)-\left[(x-\frac{1}{2})\log x-x+\frac{1}{2}\log(2\pi)\right]<\frac{1}{x}$
for all $x>1$. It can be verified that $\lim_{x\rightarrow\infty}\frac{r(x)}{x\log(\nu+x)}=1$, which implies $\lim_{x\rightarrow\infty}r(x)=\infty$. $\Box$
Proof of Theorem 3.2: According to Lemma 3.3, $r(x)$ attains its minimum at $x_{*}>\lambda$.
We first show that $r(x_{*})<0$. Otherwise, if $r(x_{*})\geq 0$, then $r(x)\geq 0$ for all $x\geq 0$ and $r(x)>0$ for all $x\neq x_{*}$. Then equivalently, we have $f_{NB}(x)\geq f_{\lambda}(x)$ for all $x\geq 0$ and $f_{NB}(x)>f_{\lambda}(x)$ for all $x\neq x_{*}$, which implies $\sum_{y=0}^{\infty}f_{NB}(y)>\sum_{y=0}^{\infty}f_{\lambda}(y)$. However, $\sum_{y=0}^{\infty}f_{NB}(y)=\sum_{y=0}^{\infty}f_{\lambda}(y)=1$. The contradiction implies that we must have $r(x_{*})<0$.
Note that the sign of $r(y)$ is exactly the same as the sign of $d(y)$ for all $y\in\{0,1,2,\ldots,\}$. We let $K_{1}$ be the largest possible integer such that $d(y)\geq 0$ for all $0\leq y\leq K_{1}$, and $K_{2}$ be the smallest possible integer such that $d(y)\geq 0$ for all $y\geq K_{2}$ Note that we must have $K_{2}-K_{1}\geq 2$ and $d(y)<0$ for each $K_{1}<y<K_{2}$ (otherwise, we have the contradiction $\sum_{y=0}^{\infty}f_{NB}(y)>\sum_{y=0}^{\infty}f_{\lambda}(y)$ again).
According to Lemma 3.3, $r(0)>0$ and thus $d(0)=D(0)>0$. Since $d(y)>0$ for $0\leq y<K_{1}$ and $y>K_{2}$, then $D(y)$ strictly increases on $0\leq y<K_{1}$ and $y>K_{2}$; while $d(y)<0$ for $K_{1}<y<K_{2}$ implies that $D(y)$ strictly decreases on $K_{1}<y<K_{2}$ Since $\lim_{y\rightarrow\infty}D(y)=1-1=0$, then $D(y)$ must first increase and attain its maximum $D(K_{1})>0$, then strictly decrease and attain its minimum $D(K_{2})<0$, and then strictly increase after $K_{2}$ and approach to 00 as $y$ goes to $\infty$. We let $K_{*}$ be the largest possible integer such that $D(y)\geq 0$. Then we must have $K_{1}<K_{*}<K_{2}$, $D(y)>0$ for all $0\leq y<K_{*}$, and $D(y)<0$ for all $y>K_{*}$Box
[…]
$\sum_{y=0}^{\infty}[1-F(y)]$ $=$ $\sum_{y=0}^{\infty}\sum_{k=y+1}^{\infty}P(Y=k)=\sum_{k=1}^{\infty}\sum_{y=0}^{k-1}P(Y=k)$
$=$ $\sum_{k=1}^{\infty}k\cdot P(Y=k)$
$=$ $E(Y)$
Since $p=\nu/(\nu+\lambda)$, the means of NB$(\nu,p)$ and Poisson$(\lambda)$ are the same. That is, $\sum_{y=0}^{\infty}[1-F_{NB}(y)]=\sum_{y=0}^{\infty}[1-F_{\lambda}(y)]=\lambda$. Then
$\sum_{y=0}^{\infty}D(y)=\sum_{y=0}^{\infty}[F_{NB}(y)-F_{\lambda}(y)]=\sum_{y=0}^{\infty}[1-F_{\lambda}(y)]-\sum_{y=0}^{\infty}[1-F_{NB}(y)]=0$
$\Box$
Proof of Theorem 3.3: Given $\lambda>0$ and $\nu>0$, we let $p=\nu/(\nu+\lambda)\in(0,1)$. Then the mean of NB$(\nu,p)$ is $\mu=\nu(1-p)/p=\nu\cdot\lambda/\nu=\lambda$. According to Lemmas 3.1 and 3.2, $G_{NB}(\nu)=\sum_{y=0}^{\infty}\frac{1-F_{NB}(y)}{\nu+y}-\log\left(1+\frac{\lambda}{\nu}\right)=0$. Since $G_{\lambda}(\nu)=\sum_{y=0}^{\infty}\frac{1-F_{\lambda}(y)}{\nu+y}-\log\left(1+\frac{\lambda}{\nu}\right)$, we only need to show that $\sum_{y=0}^{\infty}\frac{1-F_{\lambda}(y)}{\nu+y}>\sum_{y=0}^{\infty}\frac{1-F_{NB}(y)}{\nu+y}$.
According to Theorem 3.2 and Lemma 3.4, we have $D(y)\geq 0$ for $0\leq y\leq K_{*}$, $D(y)<0$ for $y>K_{*}$, and $\sum_{y=0}^{K_{*}}D(y)=-\sum_{y=K_{*}+1}^{\infty}D(y)>0$. Then
$\sum_{y=0}^{\infty}\frac{1-F_{\lambda}(y)}{\nu+y}-\sum_{y=0}^{\infty}\frac{1-F_{NB}(y)}{\nu+y}$ $=$ $\sum_{y=0}^{\infty}\frac{F_{NB}(y)-F_{\lambda}(y)}{\nu+y}=\sum_{y=0}^{\infty}\frac{D(y)}{\nu+y}$
$=$ $\sum_{y=0}^{K_{*}}\frac{D(y)}{\nu+y}+\sum_{y=K_{*}+1}^{\infty}\frac{D(y)}{\nu+y}$
$>$ $\sum_{y=0}^{K_{*}}\frac{D(y)}{\nu+K_{*}}+\sum_{y=K_{*}+1}^{\infty}\frac{D(y)}{\nu+K_{*}+1}$
$=$ $\sum_{y=0}^{K_{*}}D(y)\cdot\left(\frac{1}{\nu+K_{*}}-\frac{1}{\nu+K_{*}+1}\right)$
$=$ $\sum_{y=0}^{K_{*}}D(y)\cdot\frac{1}{(\nu+K_{*})(\nu+K_{*}+1)}$
$>$ 00 0
$\Box$
Proof of Lemma 3.5: According to Theorem 3.1, if $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson$(\lambda)$ with $\lambda>0$, then there exists an event $A$ with probability $1$, such that, for each $\omega\in A$, $g(\nu)\rightarrow G_{\lambda}(\nu)>0$, as $n$ goes to $\infty$. We let $c=G_{\lambda}(\nu)/2>0$. Then for each $\omega\in A$, $|g(\nu)-G_{\lambda}(\nu)|<c$ and thus $g(\nu)>G_{\lambda}(\nu)-c=c$, as $n$ goes to $\infty$. $\Box$
Proof of Lemma 3.6: First of all, for any $a\in\left[\frac{1}{2}(1-e^{-\lambda}),\frac{3}{2}(1-e^{-\lambda})\right]$, $b\in\left[\frac{\lambda}{2},\frac{3\lambda}{2}\right]$,
$h(\nu,a,b)\geq\frac{1-e^{-\lambda}}{2\nu}-\log\left(1+\frac{3\lambda}{2\nu}\right)$
Since $\lim_{\nu\rightarrow 0^{+}}\frac{1-e^{-\lambda}}{2\nu}=\infty$ and
$\lim_{\nu\rightarrow 0^{+}}\frac{\log\left(1+\frac{3\lambda}{2\nu}\right)}{\frac{1-e^{-\lambda}}{2\nu}}=\lim_{\nu\rightarrow 0^{+}}\frac{\frac{3\lambda}{2}}{(1+\frac{3\lambda}{2\nu})\frac{1-e^{-\lambda}}{2}}=0$
then $\lim_{\nu\rightarrow 0^{+}}\frac{1-e^{-\lambda}}{2\nu}-\log\left(1+\frac{3\lambda}{2\nu}\right)=\infty$. Therefore, there exists a $\nu_{0}>0$, such that, $\frac{1-e^{-\lambda}}{2\nu}-\log\left(1+\frac{3\lambda}{2\nu}\right)>c$ for all $0<\nu\leq\nu_{0}$, which leads to the conclusion. $\Box$
Proof of Lemma 3.7: According to (5), for each $\nu>\nu_{0}$,
$|g^{\prime}(\nu)|$ $=$ $\left|-\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\left[\frac{1}{\nu^{2}}+\cdots+\frac{1}{(\nu+y-1)^{2}}\right]+\frac{1}{\nu}-\frac{1}{\nu+\bar{Y}_{n}}\right|$
$\leq$ $\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\left[\frac{1}{\nu^{2}}+\cdots+\frac{1}{(\nu+y-1)^{2}}\right]+\frac{2}{\nu_{0}}$
$\leq$ $\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\cdot\sum_{k=0}^{\infty}\frac{1}{(\nu+k)^{2}}+\frac{2}{\nu_{0}}$
$\leq$ $1\cdot\sum_{k=0}^{\infty}\frac{1}{(\nu_{0}+k)^{2}}+\frac{2}{\nu_{0}}$
$=$ $\frac{1}{{\nu_{0}}^{2}}+\sum_{k=1}^{\infty}\frac{1}{(\nu_{0}+k)^{2}}+\frac{2}{\nu_{0}}$
$\leq$ $\frac{1}{{\nu_{0}}^{2}}+\sum_{k=1}^{\infty}\frac{1}{k^{2}}+\frac{2}{\nu_{0}}$
$=$ $\frac{1}{{\nu_{0}}^{2}}+\frac{\pi}{6}+\frac{2}{\nu_{0}}$
Let $M=\frac{1}{{\nu_{0}}^{2}}+\frac{2}{\nu_{0}}+\frac{\pi}{6}>0$. Then $|g^{\prime}(\nu)|\leq M<\infty$ for all $\nu>\nu_{0}$Box
Proof of Theorem 3.4: Given $\lambda>0$ and $\nu_{\rm max}>0$, we set $c=1$ in Lemma 3.6 and obtain $\nu_{0}\in(0,\nu_{\rm max})$ as in Lemma 3.6 and event $A$ as defined in (13). Then for each $\omega\in A$, let $N_{0}(\omega)$ be a large enough integer, such that for all $n\geq N_{0}(\omega)$, #$\frac{\#\{i:Y_{i}(\omega)>0\}}{n}\in\left[\frac{1}{2}(1-e^{-\lambda}),\frac{3}{2}(1-e^{-\lambda})\right]$, $\bar{Y}_{n}(\omega)\in\left[\frac{\lambda}{2},\frac{3\lambda}{2}\right]$, and thus $g(\nu)(\omega)\geq 1$ for all $0<\nu\leq\nu_{0}$.
According to Lemma 3.1, $G_{\lambda}(\nu)$ is continuous on $(0,\infty)$. According to Theorem 3.3, $G_{\lambda}(\nu)>0$ for each $\nu>0$. Therefore, for the given $0<\nu_{0}<\nu_{\rm max}<\infty$, there exists a $c\in(0,1)$ such that $G_{\lambda}(\nu)>c$ for all $\nu\in[\nu_{0},\nu_{max}]$. On the other hand, according to Lemma 3.7, there exists an $M>0$, which does not depend on $\omega\in\Omega$, such that $|g^{\prime}(\nu)|\leq M<\infty$ for all $\nu>\nu_{0}$ and all $\omega\in\Omega$.
We let $K=4(\nu_{\rm max}-\nu_{0})M/c+1$, then $(\nu_{\rm max}-\nu_{0})M/K<c/4$. We denote the grid points, $\nu_{i}=\nu_{0}+(\nu_{\rm max}-\nu_{0})\cdot i/K$, for $i=1,\ldots,K$. According to Theorem 3.1, for each $i$, $g(\nu_{i})\rightarrow G_{\lambda}(\nu_{i})$ almost surely. We denote
$A_{K}=\{\omega\in A\mid\lim_{n\rightarrow\infty}g(\nu_{i})(\omega)=G(\nu_{i}),i=1,\ldots,K\}$
which has probability $1$. For each $\omega\in A_{K}$, there exist $N_{i}(\omega)$, $i=1,\ldots,K$, such that, $g(\nu_{i})(\omega)\geq\frac{c}{2}$ for all $n\geq N_{i}(\omega)$. We denote $N(\omega)=\max\{N_{0}(\omega),N_{1}(\omega),\ldots,N_{K}(\omega)\}$.
For each $\nu\in(\nu_{0},\nu_{\rm max}]$, there exists a $\nu_{i}$, such that, $\nu\in(\nu_{i-1},\nu_{i}]$. Then
$|g(\nu)(\omega)-g(\nu_{i})(\omega)|\leq M\cdot\frac{\nu_{\rm max}-\nu_{0}}{K}<M\cdot\frac{c}{4M}=\frac{c}{4}$
which implies $g(\nu)(\omega)>g(\nu_{i})(\omega)-\frac{c}{4}\geq\frac{c}{2}-\frac{c}{4}=\frac{c}{4}$, for all $n\geq N(\omega)$.
After all, for each $\omega\in A_{K}$ and each $n\geq N(\omega)$, $g(\nu)(\omega)>\frac{c}{4}>0$ for all $0<\nu\leq\nu_{\rm max}$. According to Theorem 2.2, under the constraint $\nu\in(0,\nu_{\rm max}]$, the MLE of $\nu$ for a negative binomial distribution is $\hat{\nu}(\omega)=\nu_{\rm max}$. The MLE of $p$ is via (1). $\Box$
Proof of Theorem 3.5: According to the proof of Theorem 3.4, for each positive integer $k$, there exists an event $A_{k}$ with probability $1$, such that, for each $\omega\in A_{k}$, there exist an integer $N(\omega,k)$ and a lower bound $c_{k}>0$, such that, $g(\nu)>c_{k}>0$ for all $n>N(\omega,k)$ and all $\nu\in(0,\nu_{\rm max}(k)]$.
We let $A_{0}=\cap_{k=1}^{\infty}A_{k}$ Then $P(A_{0})=1$ as well. Since $\nu_{\rm max}(n)\rightarrow\infty$ as $n$ goes to infinity, then for any $M>0$, there exists an integer $K_{M}>0$, such that, $\nu_{\rm max}(k)>M$ for all $k\geq K_{M}$ For each $\omega\in A_{0}$, given any $M>0$, we let $N_{M}=\max_{1\leq k\leq N_{M}}N(\omega,k)$. Then $g(\nu)(\omega)>\min_{1\leq k\leq K_{M}}c_{k}>0$, for all $\nu\in(0,M]$ and all $n>N_{M}(\omega)$, which implies that the MLE $\hat{\nu}_{n}(\omega)>M$. That is, $\hat{\nu}_{n}(\omega)\rightarrow\infty$, as $n$ goes to infinity. $\Box$
Proof of Theorem 3.6: We denote $F_{\lambda}(y)$ as the CDF of Poisson($\lambda$). For any given $\epsilon>0$, there exists an integer $M>0$, such that, $1-F_{\lambda}(y)\leq\frac{\epsilon}{6}$ for all $y\geq M$. Then
$D_{n}$ $=$ $\sup_{y}|F_{n}(y)-F_{\hat{\bm{\theta}}}(y)|$
$=$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)+F_{\lambda}(y)-F_{\hat{\bm{\theta}}}(y)\right|$
$\leq$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)\right|+\sup_{y<M}\left|F_{\lambda}(y)-F_{\hat{\bm{\theta}}}(y)\right|+\sup_{y\geq M}\left|F_{\lambda}(y)-F_{\hat{\bm{\theta}}}(y)\right|$
$=$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)\right|+\sup_{y<M}\left|F_{\lambda}(y)-F_{\hat{\bm{\theta}}}(y)\right|$
$+\ \sup_{y\geq M}\left|\left[1-F_{\lambda}(y)\right]-\left[1-F_{\hat{\bm{\theta}}}(y)\right]\right|$
$\leq$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)\right|+\sup_{y<M}\left|F_{\lambda}(y)-F_{\hat{\bm{\theta}}}(y)\right|$
$+\ \sup_{y\geq M}\left[1-F_{\lambda}(y)\right]+\sup_{y\geq M}\left[1-F_{\hat{\bm{\theta}}}(y)\right]$
$\leq$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)\right|+\sup_{y<M}\left|F_{\lambda}(y)-F_{\hat{\bm{\theta}}}(y)\right|+\left[1-F_{\lambda}(M)\right]+\left[1-F_{\hat{\bm{\theta}}}(M)\right]$
According to the Glivenko–Cantelli theorem (see, for example, Theorem 7.5.2 in Resnick (1999)), $\sup_{y}|F_{n}(y)-F_{\lambda}(y)|\rightarrow 0$ almost surely, as $n\rightarrow\infty$.
Let $A=\{\omega\in\Omega\mid\lim_{n\rightarrow\infty}\hat{\nu}_{n}(\omega)=\infty\mbox{ and }\lim_{n\rightarrow\infty}\sup_{y}|F_{n}(y)(\omega)-F_{\lambda}(y)|=0\}$. Then according to Theorem 3.5 and the Glivenko-Cantelli theorem, $P(A)=1$.
For each $\omega\in A$ and the given $\epsilon>0$, there exists a constant $N_{e}(\omega)>0$, such that, $\sup_{y}|F_{n}(y)(\omega)-F_{\lambda}(y)|\leq\frac{\epsilon}{3}$ for all $n\geq N_{e}(\omega)$.
According to Theorem 5.13.1 in Fisz (1963), $f_{{\rm NB}(\nu,p(\nu))}(k)\rightarrow f_{\lambda}(k)$ as $\nu\rightarrow\infty$ for each $k$, where $p(\nu)=\frac{\nu}{\nu+\lambda}$. Then, $F_{{\rm NB}(\nu,p(\nu))}(k)\rightarrow F_{\lambda}(k)$ as $\nu\rightarrow\infty$, for each $k$. For each $k=0,1,\ldots,M$, there exists a $V_{k}>0$ such that $|F_{{\rm NB}(\nu,p(\nu))}(k)-F_{\lambda}(k)|<\frac{\epsilon}{6}$ for all $\nu\geq V_{k}$. Let $V_{\rm max}=\max\{V_{0},V_{1},\ldots,V_{M}\}$. Then for all $\nu\geq V_{\rm max}$, $\sup_{k<M}|F_{{\rm NB}(\nu,p(\nu))}(k)-F_{\lambda}(k)|\leq\frac{\epsilon}{6}$ and $|F_{{\rm NB}(\nu,p(\nu))}(M)-F_{\lambda}(M)|\leq\frac{\epsilon}{6}$.
In addition, for each $k$,
$\lim_{n\rightarrow\infty}F_{\hat{\bm{\theta}}(\omega)}(k)=\lim_{n\rightarrow\infty}F_{{\rm NB}(\hat{\nu}_{n}(\omega),\hat{p}_{n}(\hat{\nu}_{n}(\omega)))}(k)=\lim_{\nu\rightarrow\infty}F_{{\rm NB}(\nu,p(\nu))}(k)=F_{\lambda}(k)<\infty$
There exists an $N_{k}(\omega)$ such that for all $n\geq N_{k}(\omega)$, $\hat{\nu}_{n}(\omega)>V_{\rm max}$, $|F_{\hat{\bm{\theta}}(\omega)}(k)-F_{\lambda}(k)|\leq\frac{\epsilon}{6}$. Then for each $n\geq N_{\rm max}(\omega)=\max\{N_{0}(\omega),N_{1}(\omega),\ldots,N_{M}(\omega),N_{e}(\omega)\}$, $\sup_{k<M}|F_{\hat{\bm{\theta}}(\omega)}(k)-F_{\lambda}(k)|\leq\frac{\epsilon}{6}$ and $0<1-F_{\hat{\bm{\theta}}(\omega)}(M)\leq 1-F_{\lambda}(M)+|F_{\hat{\bm{\theta}}(\omega)}(M)-F_{\lambda}(M)|\leq\frac{\epsilon}{6}+\frac{\epsilon}{6}=\frac{\epsilon}{3}$, which leads to $D_{n}(\omega)\leq\epsilon$.
Therefore, $D_{n}(\omega)$ goes to zero as $n$ goes to infinity. $\Box$