### v2

# From Poisson Observations to Fitted Negative Binomial Distribution

Abstract
Negative binomial distribution has been widely used as a more flexible model than Poisson distribution for count data. When the observations come from a Poisson distribution, it is often challenging to rule out the possibility that the data come from a negative binomial distribution with extreme parameter values. To address this phenomenon, we develop a more efficient and accurate algorithm for finding the maximum likelihood estimate of negative binomial parameters, which outperforms the state-of-the-art programs for the same purpose. We also theoretically justify that the negative binomial distribution with parameters estimated from Poisson data converges to the true Poisson distribution with probability one. As a solution to this phenomenon, we extend the negative binomial distributions with a new parameterization, which include Poisson distributions as a special class.
Key words and phrases: Empirical distribution function, extended negative binomial distribution, Kolmogorov-Smirnov (KS) test, maximum likelihood estimate (MLE), Poisson distribution, profile log-likelihood
1 Introduction
Count data arises naturally from many different scientific disciplines. When the events or subjects of interest occur or are positioned randomly and independently in space or time, Poisson distribution is often a good model (Pielou, 1977; Ross and Preece, 1985). It has a probability mass function (PMF)
$f_{\lambda}(y)=\frac{\lambda^{y}}{y!}e^{-\lambda},\>\>\>y=0,1,2,\ldots,$ (1)
where $\lambda>0$ is both the mean and variance of the distribution. However, in practice the phenomenon of over-dispersion, that is, the variance is larger than the mean, is frequently observed (Bliss and Fisher, 1953; Ross and Preece, 1985). For those scenarios, a negative binomial distribution with two parameters is more appropriate.
Following Chapter 32 in Forbes et al. (2011), a negative binomial (NB) distribution with $\bm{\theta}=(\nu,p)^{T}\in(0,\infty)\times(0,1)$ has a PMF
$f_{\bm{\theta}}(y)=\frac{\Gamma(\nu+y)}{\Gamma(\nu)\Gamma(y+1)}p^{\nu}(1-p)^{y},\>\>\>y=0,1,2,\ldots$ (2)
with mean $\nu(1-p)/p$ and variance $\nu(1-p)/p^{2}$. When $\nu$ takes a positive-integer value, the PMF (2) describes the distribution of the number of failures before the $\nu$th success in a sequence of independent Bernoulli trials, where each trial has a success probability of $p$. This is commonly referred to as the Pascal variate (Forbes et al., 2011).
Alternative parameterizations have been used in the literature as well, including $(\nu,\mu=\nu(1-p)/p)$ as in Anscombe (1949) or Bliss and Fisher (1953), $(\nu,P=(1-p)/p)$ as in Johnson et al. (2005), and $(\nu,P=1-p)$ as in Anscombe (1950) and Aldirawi and Yang (2022). For more detailed reviews on the history of negative binomial distribution, please see Ross and Preece (1985), Section 5.1 in Johnson et al. (2005), and references therein.
A natural question is whether we can tell in practice that the data come from a Poisson or a negative binomial distribution. If the true distribution is a negative binomial distribution with some unknown $\nu>0$ and $p\in(0,1)$, one can eventually tell that the true distribution is not Poisson because the sample variance will be significantly larger than the sample mean as the sample size goes to infinity.
However, if the data come from a Poisson distribution with some unknown parameter $\lambda>0$, the answer is essentially negative. We use the Kolmogorov-Smirnov (KS) test (Massey Jr, 1951) as an illustration, which has been widely used as a goodness-of-fit test.
The original KS test was proposed to assess whether a given random sample $\{Y_{1},\ldots,$ $Y_{n}\}$ originates from a continuous cumulative distribution function (CDF) $F_{\bm{\theta}}(y)$, characterized by specific model parameter(s) $\bm{\theta}$. The KS test statistic is defined as $D_{n}=\sup_{y}|F_{n}(y)-F_{\bm{\theta}}(y)|$, where $F_{n}(y)=n^{-1}\sum_{i=1}^{n}{\mathbf{1}}_{(-\infty,y]}(Y_{i})$ represents the empirical distribution function (EDF) of the sample. Dimitrova et al. (2020) extended the applicability of the KS test to encompass a broader range of distributions $F_{\bm{\theta}}(y)$, including discrete and mixed distributions with known $\bm{\theta}$. When $\bm{\theta}$ is unknown, Lilliefors (1967, 1969) suggested replacing $\bm{\theta}$ with its estimate $\hat{\bm{\theta}}$ when calculating $D_{n}$ Adjusted $p$-values or thresholds for test statistics can be estimated by parametric bootstrapping (Henze, 1996), nonparametric bootstrapping (Aldirawi et al., 2019), or nested bootstrapping (Dousti Mousavi et al., 2023) under those circumstances. R packages, such as iZID (Wang et al., 2020) and AZIAD (Dousti Mousavi et al., 2023), have been developed for such purposes.
For illustrative purposes, we adopt the parametric bootstrapping proposed by Henze (1996) to calculate the corresponding $p$-values. More specifically, we simulate random samples from Poisson($\lambda=10$) with various sample sizes. For each random sample of size $n$, we calculate the maximum likelihood estimate (MLE) $\hat{\bm{\theta}}=(\hat{\nu},\hat{p})$, the KS test statistic $D_{n}=\sup_{y}|F_{n}(y)-F_{\hat{\bm{\theta}}}(y)|$, the critical number $d_{n}$ such that $P(D_{n}\geq d_{n})=0.05$ based on the parametric bootstrapping (Henze, 1996), and the empirical power, that is, the ratio of rejecting $H_{0}$ (NB) out of $1,000$ random samples from Poisson($10$). The results are listed in Table 1. There is a clear trend that (i) $D_{n}$ vanishes as $n$ goes to infinity; (ii) the critical number $d_{n}$ also vanishes as $n$ goes to infinity; and (iii) the empirical power is poor (roughly $0.05$ or below) and not getting better as $n$ goes to infinity.
Table 1: KS test for NB ($H_{0}$) against Poisson ($H_{1}$) given $Y_{1},\ldots,Y_{n}$ iid $\sim$ Poisson($\lambda=10$)
Sample Size $n$ $50$ $500$ $5000$ $50000$
$\hat{\nu}$ 36.8779 220.0356 10000 10000
$\hat{p}$ 0.7804 0.9568 0.9990 0.9990
Test Statistic $D_{n}$ 0.0450 0.0154 0.0048 0.0014
Critical Number $d_{n}$ 0.0333 0.0118 0.0036 0.0011
Empirical Power at Poisson($10$) 0.029 0.052 0.039 0.046
According to (1) and (2), Poisson distributions and negative binomial distributions are two non-overlapped classes of distributions. However, as shown in Table 1, it is challenging to determine that the data originate from a Poisson distribution rather than a negative binomial distribution, even with a large sample size. This phenomenon is mainly due to the mathematical fact that a Poisson distribution can be approximated well by a negative binomial distribution. For example, as a direct conclusion of Theorem 5.13.1 in Fisz (1963), $F_{{\rm NB}(\nu,p(\nu))}(k)\rightarrow F_{\lambda}(k)$ as $\nu\rightarrow\infty$ for each $k$, where $p(\nu)=\nu/(\nu+\lambda)$, $F_{{\rm NB}(\nu,p)}$ is the CDF of negative binomial distribution with parameters $\nu>0$ and $p\in(0,1)$, and $F_{\lambda}$ is the CDF of Poisson distribution with parameter $\lambda>0$.
To solve this puzzle, in this paper, we propose a new parameterization of NB distributions with $(\mu=\nu(1-p)/p,\ p)$, and extend the NB family from $(\mu,p)\in(0,\infty)\times(0,1)$ to $[0,\infty)\times(0,1]$. The extended NB family not only contains the regular NB distributions, but also two special cases, Poisson($\lambda$) with $(\mu=\lambda>0,p=1)$, and a singleton distribution $P(Y_{i}=0)=1$ with $\mu=0$. We show theoretically that if $Y_{1},\ldots,Y_{n}$ are iid from Poisson($\lambda$) but we treat it as a NB random sample and find the corresponding MLE $(\hat{\mu}_{n},\hat{p}_{n})$, then with probability $1$, (i) $\hat{\mu}_{n}=n^{-1}\sum_{i=1}^{n}Y_{n}$ and $\lim_{n\to\infty}\hat{p}_{n}=1$ (Theorem 4.5), and (ii) $\lim_{n\rightarrow\infty}\sup_{y}|F_{\hat{\mu}_{n},\hat{p}_{n}}(y)-F_{\lambda}(y)|=0$ (Theorem 4.6), where $F_{\hat{\mu}_{n},\hat{p}_{n}}$ is the CDF of extended NB distribution with estimated parameters.
The rest of this paper is organized as follows. In Section 2, we propose a new algorithm for finding the MLE of NB distribution, which outperforms the state-of-the-art algorithms designed for the same purpose. In Section 3, we theoretically justify the asymptotic behaviors of the MLE for NB distribution given a Poisson random sample. We also show that the NB distribution with estimated parameters converges to the true Poisson distribution with probability 1 (Corollary 1). In Section 4, we provide asymptotic results of MLEs under other parameterizations, including the new $(\mu,p)$ parameterization that we recommend. In Section 5, we numerically justify that our algorithm proposed in Section 2 outperforms the state-of-the-art programs available for users. We conclude in Section 6.
2 Finding MLE for Negative Binomial Distribution
In this section, we consider finding the maximum likelihood estimates (MLE) of $(\nu,p)$ given $Y_{1},\ldots,Y_{n}$ iid $\sim$ NB$(\nu,p)$ under the constraint $\nu\in(0,\nu_{\rm max}]$ and $p\in(0,1]$, where $\nu_{\rm max}>0$ is pre-determined. It is typically the situation when finding the MLE numerically for negative binomial distributions (see, e.g., Dousti Mousavi et al. (2023)).
2.1 Likelihood equation
In this section, we summarize some key results in the literature on finding the MLE for negative binomial parameters taking the form of (2). Then
$\log f_{\bm{\theta}}(y)=\log\Gamma(\nu+y)-\log\Gamma(\nu)-\log\Gamma(y+1)+\nu\log p+y\log(1-p)\ .$
Taking partial derivatives, we have
$\frac{\partial\log f_{\bm{\theta}}(y)}{\partial\nu}$ $=$ $\Psi(\nu+y)-\Psi(\nu)+\log p\ ,$
$\frac{\partial\log f_{\bm{\theta}}(y)}{\partial p}$ $=$ $\frac{\nu}{p}-\frac{y}{1-p}\ ,$
where $\Psi(\cdot)=\Gamma^{\prime}(\cdot)/\Gamma(\cdot)$ is known as the digamma function (see, for example, Section 6.3 in Abramowitz and Stegun (1970)).
Suppose $Y_{1},\ldots,Y_{n}$ are iid $\sim$ NB$(\nu,p)$. The log-likelihood $l(\bm{\theta})=\sum_{i=1}^{n}\log f_{\bm{\theta}}(Y_{i})$ and
$\frac{\partial l(\bm{\theta})}{\partial p}=\frac{n\nu}{p}-\frac{\sum_{i=1}^{n}Y_{i}}{1-p}=\frac{n(\nu+\bar{Y}_{n})}{p(1-p)}\left(\frac{\nu}{\nu+\bar{Y}_{n}}-p\right)\left\{\begin{array}[]{ll}>0&\mbox{ if }p<\frac{\nu}{\nu+\bar{Y}_{n}}\\ =0&\mbox{ if }p=\frac{\nu}{\nu+\bar{Y}_{n}}\\ <0&\mbox{ if }p>\frac{\nu}{\nu+\bar{Y}_{n}}\end{array}\right.\>,$
where $\bar{Y}_{n}=n^{-1}\sum_{i=1}^{n}Y_{i}$ That is, $\partial l(\bm{\theta})/\partial p=0$, known as the likelihood equation of $p$, always implies $\hat{p}=\hat{\nu}/(\hat{\nu}+\bar{Y}_{n})$ for the MLE $(\hat{\nu},\hat{p})$.
Example 2.1.
A degenerated NB distributionThe solution $\hat{p}=\hat{\nu}/(\hat{\nu}+\bar{Y}_{n})$ for the likelihood equation of $p$ suggests that $\hat{p}=1$ if $\bar{Y}_{n}=0$, which is possible in practice. To extend the range of $\bm{\theta}=(\nu,p)^{T}$ from $(0,\infty)\times(0,1)$ to $(0,\infty)\times(0,1]$, we consider the limit of $\log f_{\bm{\theta}}(y)$ as $p\to 1$ given $\nu>0$. Since $\lim_{p\to 1-}\log f_{\bm{\theta}}(y)=0$ for $y=0$; and $\infty$- ∞ for $y>0$, which correspond to $f_{\bm{\theta}}(y)=1$ for $y=0$; and 00 for $y>0$, then the limiting distribution satisfies $P(Y_{i}=0)=1$. In other words, if we regard this singleton distribution, which generates zero with probability $1$, as a degenerated NB distribution as $p\to 1$, then the parameter space for the NB distribution is $\bm{\Theta}=(0,\infty)\times(0,1]$. $\Box$
Example 2.2.
A trivial case of all zerosA special case of the data occurs when $Y_{1}=\cdots=Y_{n}=0$, which is possible in practice especially with a moderate $n$. In this case, $\bar{Y}_{n}=0$ as well. It can be verified that $l(\bm{\theta})=n\nu\log p\leq 0$. To maximize $l(\bm{\theta})$ with $\nu\in(0,\infty)$ and $p\in(0,1]$, we must have $\hat{p}=1$ and $\hat{\nu}$ can be any positive number. $\Box$
As a conclusion, regardless $\bar{Y}_{n}=0$ or $\bar{Y}_{n}>0$, the MLE of $p$ given $\nu>0$ must take the form as follows
$\hat{p}=\hat{p}(\nu)=\frac{\nu}{\nu+\bar{Y}_{n}}$ (3)
(see, e.g., Bliss and Fisher (1953) or Section 5.8.2 in Johnson et al. (2005)).
To avoid the trivial case as in Example 2.2, from now on, we assume $\bar{Y}_{n}>0$, that is, $Y_{i}>0$ for at least one $i$.
Since the best $p$ must be the function $\hat{p}(\nu)$ of $\nu$, maximizing $l(\bm{\theta})=l(\nu,p)$ is equivalent to maximizing the profile log-likelihood function $h(\nu)=l(\nu,\hat{p}(\nu))$ for $\nu$ (Cox and Snell, 1989; Venzon and Moolgavkar, 1988).
$h(\nu)=l(\nu,\hat{p}(\nu))$ (5)
$=$ $\sum_{i=1}^{n}\log\Gamma(\nu+Y_{i})-\sum_{i=1}^{n}\log\Gamma(Y_{i}+1)-n\log\Gamma(\nu)$
$-\ n(\nu+\bar{Y}_{n})\log(\nu+\bar{Y}_{n})+n\nu\log\nu+n\bar{Y}_{n}\log\bar{Y}_{n}$
$=$ $\sum_{y\in I}f_{y}\cdot\log\Gamma(\nu+y)-\sum_{y\in I}f_{y}\cdot\log\Gamma(y+1)-n\log\Gamma(\nu)$
$-\ n(\nu+\bar{Y}_{n})\log(\nu+\bar{Y}_{n})+n\nu\log\nu+n\bar{Y}_{n}\log\bar{Y}_{n}\ .$
Here $I$ in (5) is the range (or the collection of distinct values) of $\{Y_{1},\ldots,Y_{n}\}$, and $f_{y}=\#\{1\leq i\leq n\mid Y_{i}=y\}$ is the frequency of $y\in I$. According to (3), $\partial l(\nu,\hat{p}(\nu))/\partial p=0$. Then the profile score
$h^{\prime}(\nu)$ $=$ $\frac{\partial l}{\partial\nu}(\nu,\hat{p}(\nu))+\frac{\partial l}{\partial p}(\nu,\hat{p}(\nu))\cdot\hat{p}^{\prime}(\nu)=\frac{\partial l}{\partial\nu}(\nu,\hat{p}(\nu))$
$=$ $\sum_{i=1}^{n}\left[\Psi(\nu+Y_{i})-\Psi(\nu)\right]-n\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right)$
$\stackrel{{\scriptstyle\triangle}}{{=}}$ $ng(\nu)$
where (see also Bliss and Fisher (1953))
$g(\nu)$ $=$ $\frac{1}{n}\sum_{i=1}^{n}\left[\Psi(\nu+Y_{i})-\Psi(\nu)\right]-\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right)$ (6)
$=$ $\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right)$ (7)
by using the fact $\Psi(x+1)-\Psi(x)=x^{-1}$ (see, for example, Formula 6.3.5 in Abramowitz and Stegun (1970)). The same technique has also been used for calculating the Fisher information matrices related to negative binomial distributions (Yu et al., 2024).
In the literature, $g(\nu)$, defined by (6) or (7), is known as the efficient score function of $\nu$ up to a constant $n$ (Bliss and Fisher, 1953). Essentially, $g(\nu)=0$ is the likelihood equation of $\nu$ for negative binomial distributions. If $g(\nu)=0$ has a unique solution $\hat{\nu}\in(0,\infty)$, then $\hat{\nu}$ must be the MLE of $\nu$.
In our notation, Anscombe (1950) conjectured that $g(\nu)=0$ has a unique solution if and only if $S_{n}^{2}>\bar{Y}_{n}$, where $S_{n}^{2}=n^{-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}$. Simonsen (1976, 1980) proved Anscombe’s conjecture and obtained the following theorem. For more efforts made in solving this conjecture and relevant problems, please see Wang (1996), Bandara et al. (2019), and reference therein.
Theorem 2.1 (Simonsen (1976, 1980)).
Denote $M=\max\{Y_{1},\ldots,Y_{n}\}$.
  • (1)
    If $M=1$ or $M\geq 2$ and $S_{n}^{2}\leq\bar{Y}_{n}$, then $g(\nu)=0$ has no solution.
  • (2)
    If $M\geq 2$ and $S_{n}^{2}>\bar{Y}_{n}$, then $g(\nu)=0$ has a unique solution in $(0,\infty)$.
Actually, if $M=1$, we always have $\bar{Y}_{n}>0$ and $S_{n}^{2}=\bar{Y}_{n}(1-\bar{Y}_{n})<\bar{Y}_{n}$ The following theorem covers the case of $M=0$.
Theorem 2.2.
Under the constraint $\nu\in(0,\nu_{\rm max}]$ and $p\in(0,1]$, the MLE of $p$ is $\hat{p}=1$ if and only if $M=0$. In this case, $\hat{\nu}$ can be any positive number in $(0,\nu_{\rm max}]$.
The proof of Theorem 2.2, as well as other proofs, is relegated to the Supplementary Material.
If the data come from a Poisson distribution, the probability of $S_{n}^{2}>\bar{Y}_{n}$ is roughly between $0.4$ and $0.5$ (see Table 2). In this case, the MLE of $\nu$ can only be obtained by numerical algorithms.
Table 2: $P(S_{n}^{2}>\bar{Y}_{n})$ estimated by $1000$ random samples of size $n$ from Poisson($\lambda$)
n
$\bm{\lambda}$ 50 500 5000 50000 500000
1 0.408 0.500 0.497 0.476 0.484
3 0.441 0.445 0.461 0.503 0.487
5 0.406 0.495 0.490 0.476 0.501
10 0.408 0.487 0.496 0.496 0.469
2.2 Efficient score function
According to Theorem 2.1, the efficient score function $g(\nu)$ does not have a solution in $(0,\infty)$ if $\bar{Y}_{n}>0$ and $S_{n}^{2}\leq\bar{Y}_{n}$ In this section, we reveal more properties of $g(\nu)$ and also show that the MLE $\hat{\nu}=\nu_{\rm max}$ under the constraint $\nu\in(0,\nu_{\rm max}]$, if $g(\nu)=0$ does not have a solution in $(0,\infty)$.
Due to (6) or (7), the first-order derivative of $g(\nu)$ is
$g^{\prime}(\nu)$ $=$ $\frac{1}{n}\sum_{i=1}^{n}\Psi_{1}(\nu+Y_{i})-\Psi_{1}(\nu)+\frac{1}{\nu}-\frac{1}{\nu+\bar{Y}_{n}}$ (8)
$=$ $-\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\left[\frac{1}{\nu^{2}}+\cdots+\frac{1}{(\nu+y-1)^{2}}\right]+\frac{\bar{Y}_{n}}{\nu(\nu+\bar{Y}_{n})}\ ,$
where $\Psi_{1}(\cdot)=\Psi^{\prime}(\cdot)$ is known as the trigamma function (see, for example, Section 6.4 in Abramowitz and Stegun (1970)).
Lemma 2.1.
Suppose $\bar{Y}_{n}>0$. Then $\lim_{\nu\rightarrow 0^{+}}g(\nu)=\infty$, $\lim_{\nu\rightarrow\infty}g(\nu)=0$, $\lim_{\nu\rightarrow 0^{+}}g^{\prime}(\nu)$ =absent=$\infty$= - ∞, and $\lim_{\nu\rightarrow\infty}g^{\prime}(\nu)=0$. Furthermore, $\lim_{\nu\rightarrow 0^{+}}\nu g(\nu)=1-f_{0}/n>0$, $\lim_{\nu\rightarrow\infty}{\nu}^{2}g(\nu)$ $=\frac{1}{2}\left(\bar{Y}_{n}-S^{2}_{n}\right)$, $\lim_{\nu\rightarrow 0^{+}}\nu^{2}g^{\prime}(\nu)=f_{0}/n-1<0$, and $\lim_{\nu\rightarrow\infty}{\nu}^{3}g^{\prime}(\nu)=S^{2}_{n}-\bar{Y}_{n}$, where $S^{2}_{n}=n^{-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}$.
Example 2.3.
A trivial case of constant observationsSuppose $Y_{1}=\cdots=Y_{n}=c\geq 1$. Then $\bar{Y}_{n}=c$, $I=\{c\}$, $f_{c}=n$, and thus
$g^{\prime}(\nu)$ $=$ $-\frac{1}{\nu^{2}}-\cdots-\frac{1}{(\nu+c-1)^{2}}+\frac{c}{\nu(\nu+c)}$
$<$ $-\frac{1}{\nu(\nu+1)}-\cdots-\frac{1}{(\nu+c-1)(\nu+c)}+\frac{c}{\nu(\nu+c)}$
$=$ $-\frac{1}{\nu}+\frac{1}{\nu+1}-\cdots-\frac{1}{v+c-1}+\frac{1}{\nu+c}+\frac{1}{\nu}-\frac{1}{\nu+c}$
$=$ 00 0
for all $\nu>0$. According to Lemma 2.1, $\lim_{\nu\rightarrow 0^{+}}g(\nu)=\infty$ and $\lim_{\nu\rightarrow\infty}g(\nu)=0$, we must have $g(\nu)>0$ for all $\nu>0$ since $g(\nu)$ is strictly decreasing for all $\nu>0$ in this case. According to the definition of $g(\nu)$, it implies that $\partial l(\bm{\theta})/\partial\nu=ng(\nu)>0$ for all $\nu>0$. Then the MLE of $\nu$ must be the maximum value that $\nu$ is allowed to take. That is, the MLE of $\nu$ under the constraint $\nu\in(0,\nu_{\rm max}]$ is $\hat{\nu}=\nu_{\rm max}$Box
Example 2.3 provides a case when $g(\nu)=0$ does not have a solution. In practice, to find the MLE of $\nu$, instead of solving $g(\nu)=0$, we may minimize $g^{2}(\nu)$, which always has a solution.
Theorem 2.3.
Suppose $\bar{Y}_{n}>0$. If $\min_{\nu\in(0,\nu_{\rm max}]}g^{2}(\nu)>0$, then the MLE of $\nu$ under the constraint $\nu\in(0,\nu_{\rm max}]$ is $\hat{\nu}=\nu_{\rm max}$
Example 2.4.
Mean is greater than varianceIn this example, we simulate a random sample of $n=5,000$ from a Poisson distribution with $\lambda=5$ and then add a constant $A$ to each of the observations. This results in a data set whose sample mean is greater than the sample variance when $A$ is large enough. Using a numerical algorithm (see Algorithm 1 in Section 2.3), we obtain the MLE of $\nu$ for $A=10,20,50,70,100$, respectively, with the constraint $\nu\in(0,10^{4}]$. Each $\hat{\nu}$ found by the numerical algorithm is $10,000$, which matches the conclusion of Theorem 2.3. $\Box$
2.3 A new algorithm for negative binomial distributions
In this section, we propose a new algorithm to find the MLE of negative binomial distributions more accurately, by incorporating the previous theoretical results.
Instead of the moment estimate $\frac{\bar{Y}_{n}^{2}}{S^{2}-\bar{Y}_{n}}$ widely used for $\nu$ (Fisher, 1941; Bliss and Fisher, 1953; Johnson et al., 2005), where $S^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}$ is the sample variance, we propose the following estimate of $\nu$ as an initial value
$\hat{\nu}^{(0)}=\min\left\{\nu_{\rm max},\>\frac{\bar{Y}_{n}^{2}}{\max\{\varepsilon,\>S^{2}-\bar{Y}_{n}\}}\right\}\ ,$ (9)
where $\nu_{\rm max}$ is a predetermined upper bound for $\nu$ (for example, $\nu_{\rm max}=10^{4}$), and $\varepsilon$ is a predetermined small positive value (for example, $\varepsilon=10^{-3}$).
We propose the following algorithm, called an adaptive profile maximization algorithm (APMA, see Algorithm 1), for finding the MLE for a negative binomial distribution.
Algorithm 1 Adaptive Profile Maximization Algorithm
Input: Data $Y_{1},\ldots,Y_{n}\in\{0,1,2,\ldots\}$, predetermined threshold values $\nu_{\rm max}>0$, $\varepsilon>0$, and $\delta\in(0,1)$ (e.g., $\nu_{\rm max}=10^{4}$, $\varepsilon=10^{-3}$, $\delta=0.1$)
Output: MLE $(\hat{\nu},\hat{p})$ for NB($\nu,p$)
1:Calculate the sample mean $\bar{Y}_{n}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}$ and the sample variance $S^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}$;
2:If $\bar{Y}_{n}=0$, go to Step $6$ with $\hat{\nu}=1$ and $\hat{p}=1$; otherwise calculate the initial value $\hat{\nu}^{(0)}$ of $\nu$ according to (9);
3:Obtain $\nu_{*}$ by maximizing $h(\nu)$ with $h^{\prime}(\nu)=ng(\nu)$ and the starting value $\hat{\nu}^{(0)}$. If the ratio of distinct values in the sample $|I|/n<\delta$, use (5) and (7) for computing $h(\nu)$ and $g(\nu)$, respectively; otherwise, use (5) and (6) instead.
4:If $h(\nu_{\rm max})>h(\nu_{*})$ (see (5) or (5)), we let $\hat{\nu}=\nu_{\rm max}$; otherwise, $\hat{\nu}=\nu_{*}$;
5:Calculate $\hat{p}=\frac{\hat{\nu}}{\hat{\nu}+\bar{Y}_{n}}$;
6:Report $\hat{\nu}$ and $\hat{p}$.
3 Negative Binomial Estimates for Poisson Sample
In this section, we derive theoretically three asymptotic results: Given a random sample of size $n$ from a Poisson distribution, as $n$ goes to infinity, with probability $1$, (i) the MLE of $\nu$ is $\nu_{\rm max}$ for negative binomial distributions under the constraint $0<\nu\leq\nu_{\rm max}$ (see Theorem 3.4 in Section 3.1); (ii) if we adjust the constraint to be $0<\nu\leq\nu_{\rm max}(n)$ with $\lim_{n\rightarrow\infty}\nu_{\rm max}(n)=\infty$, then the KS test statistic $D_{n}$ for a negative binomial distribution vanishes (see Theorem 3.6 in Section 3.2); and (iii) the NB distribution with estimated parameters converges to the true Poisson distribution (see Corollary 1 in Section 3.2).
3.1 Negative binomial MLE given Poisson sample
Recall from Section 2.2 that the efficient score function $g(\nu)$, expressed in (6) or (7), plays an important role in finding the MLE of $\nu$ for a negative binomial distribution.
[…]
$G_{F}(\nu)$ $=$ $E\Psi(\nu+Y_{1})-\Psi(\nu)-\log\left(1+\frac{\mu}{\nu}\right)$ (10)
$=$ $\sum_{y=1}^{\infty}P(Y_{1}=y)\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\log\left(1+\frac{\mu}{\nu}\right)\ ,$ (11)
which is actually the limiting function of $g(\nu)$ as $n$ goes to infinity (see Lemma 3.1 and Theorem 3.1 below).
Lemma 3.1.
Given $\mu=E(Y_{1})<\infty$, $G_{F}(\nu)$ is a continuous function on $\nu\in(0,\infty)$. Furthermore,
$G_{F}(\nu)=\sum_{y=0}^{\infty}\frac{1-F(y)}{\nu+y}-\log\left(1+\frac{\mu}{\nu}\right)\ ,$ (12)
where $F(y)=P(Y_{1}\leq y)$ is the CDF of $Y_{1}$
Theorem 3.1.
[…]
Using Kolmogorov’s Strong Law of Large Numbers (see, for example, Theorem 7.5.1 in Resnick (1999)), we provide a rigorous proof of Theorem 3.1 in the Supplementary Material. It is actually applicable for discrete distributions other than Poisson distributions as well, as long as the distribution is on nonnegative integers. For example, in Lemma 3.2, we show that $G_{F}(\nu)\equiv 0$ if $F$ is a negative binomial CDF.
Lemma 3.2.
Suppose $Y_{1},\ldots,Y_{n}$ are iid $\sim$ NB$(\nu,p)$ with $\nu>0$ and $10<p<$ 1. Then $G_{NB(\nu,p)}(\nu)=0$, for all $\nu>0$.
Suppose $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson$(\lambda)$ with $\lambda>0$. For $y\in\{0,1,2,\ldots,\}$, we denote $f_{\lambda}(y)=P(Y_{1}=y)=\frac{\lambda^{y}}{y!}e^{-\lambda}$, $F_{\lambda}(y)=P(Y_{1}\leq y)$, and
$G_{\lambda}(\nu)$ $=$ $\sum_{y=1}^{\infty}\frac{\lambda^{y}}{y!}e^{-\lambda}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\log\left(1+\frac{\lambda}{\nu}\right)$ (13)
$=$ $\sum_{y=0}^{\infty}\frac{1-F_{\lambda}(y)}{\nu+y}-\log\left(1+\frac{\lambda}{\nu}\right)$
as the almost sure limit of $g(\nu)$ given a random sample of Poisson$(\lambda)$.
Given $\lambda>0$ and $\nu>0$, we let $p=\nu/(\nu+\lambda)\in(0,1)$. Then the mean of NB$(\nu,p)$ is $\nu(1-p)/p=\nu\lambda/\nu=\lambda$, the same as Poisson$(\lambda)$’s. For such an NB$(\nu,p)$, we denote $f_{NB(\nu,p)}(y)$ and $F_{NB(\nu,p)}(y)$ as the corresponding PMF and CDF, respectively. According to Lemma 3.2,
$G_{NB(\nu,p)}(\nu)$ $=$ $\sum_{y=0}^{\infty}\frac{1-F_{NB(\nu,p)}(y)}{\nu+y}-\log\left(1+\frac{\mu}{\nu}\right)$
$=$ $\sum_{y=0}^{\infty}\frac{1-F_{NB(\nu,p)}(y)}{\nu+y}-\log\left(1+\frac{\lambda}{\nu}\right)$
$=$ $0$
Comparing $G_{NB(\nu,p)}(\nu)$ with (13), if we can show that
$\sum_{y=0}^{\infty}\frac{1-F_{\lambda}(y)}{\nu+y}>\sum_{y=0}^{\infty}\frac{1-F_{NB(\nu,p)}(y)}{\nu+y}\ ,$ (14)
then $G_{\lambda}(\nu)>G_{NB(\nu,p)}(\nu)=0$, for each $\nu>0$.
To prove (14), we first denote the continuous version of $f_{\lambda}(y)$ and $f_{NB(\nu,p)}(y)$ by
$f_{\lambda}(x)$ $=$ $\frac{\lambda^{x}}{\Gamma(x+1)}e^{-\lambda}\ ,$ (15)
$f_{NB(\nu,p)}(x)$ $=$ $\frac{\Gamma(\nu+x)}{\Gamma(x+1)\Gamma(\nu)}p^{\nu}(1-p)^{x}=\frac{\Gamma(\nu+x)}{\Gamma(x+1)\Gamma(\nu)}\cdot\frac{\nu^{\nu}\lambda^{x}}{(\nu+\lambda)^{\nu+x}}\ ,$ (16)
[…]
Lemma 3.3.
We let $r(x)=\log f_{NB(\nu,p)}(x)-\log f_{\lambda}(x)$, where $x\in[0,\infty)$ and $p=\nu/(\nu+\lambda)$. Then its first derivative
$r^{\prime}(x)\left\{\begin{array}[]{cl}<0&\mbox{ if }0<x<x_{*}\\ =0&\mbox{ if }x=x_{*}\\ >0&\mbox{ if }x>x_{*}\end{array}\right.\>,$
where $x_{*}=\Psi^{-1}\left(\log(\nu+\lambda)\right)-\nu>\lambda$. Furthermore, $r(0)>0$ and $\lim_{x\rightarrow\infty}r(x)=\infty$.
Lemma 3.3 tells us that $f_{NB(\nu,p)}(x)<\log f_{\lambda}(x)$, for $0<x<x_{*}$, and $f_{NB(\nu,p)}(x)>\log f_{\lambda}(x)$, for $x>x_{*}$ As the next step, we compare $F_{NB(\nu,p)}(y)$ and $F_{\lambda}(y)$ in Theorem 3.2.
Theorem 3.2.
Given $\lambda>0$ and $\nu>0$, we let $p=\nu/(\nu+\lambda)\in(0,1)$, $d(y)=f_{NB(\nu,p)}(y)-f_{\lambda}(y)$, and $D(y)=\sum_{k=0}^{y}d(k)=F_{NB(\nu,p)}(y)-F_{\lambda}(y)$, $y\in\{0,1,2,\ldots\}$. Then there exist three positive integers $K_{1}$, $K_{*}$, and $K_{2}$, such that, $0<K_{1}<K_{*}<K_{2}<\infty$,
$d(y)\left\{\begin{array}[]{cl}>0&\mbox{ if }0\leq y<K_{1}\mbox{ or }y>K_{2}\\ \geq 0&\mbox{ if }y=K_{1}\mbox{ or }K_{2}\\ <0&\mbox{ if }K_{1}<y<K_{2}\end{array}\right.,\>\>\>D(y)\left\{\begin{array}[]{cl}>0&\mbox{ if }0\leq y<K_{*}\\ \geq 0&\mbox{ if }y=K_{*}\\ <0&\mbox{ if }y>K_{*}\end{array}\right.\>.$
Furthermore, $D(y)$ strictly increases before $K_{1}$ and attains its maximum at $y=K_{1}$ with $D(K_{1})>0$, then strictly decreases between $K_{1}$ and $K_{2}$ and attains its minimum at $y=K_{2}$ with $D(K_{2})<0$, and then strictly increases after $K_{2}$ and approaches to 00 as $y$ goes to $\infty$.
Lemma 3.4.
Given $D(y)$ as defined in Theorem 3.2, we must have $\sum_{y=0}^{\infty}D(y)=0$.
Now we are ready to show (14) and thus $G_{\lambda}(\nu)>0$ for all $\nu>0$ in Theorem 3.3.
Theorem 3.3.
For any $\lambda>0$ and $\nu>0$, we must have $G_{\lambda}(\nu)>0$.
According to Theorem 3.1, given $Y_{1},\ldots,Y_{n}$ iid $\sim$ Poisson($\lambda$), we have $g(\nu)\rightarrow G_{\lambda}(\nu)$ almost surely, as $n$ goes to infinity. Now Theorem 3.3 guarantees that $G_{\lambda}(\nu)>0$ for all $\nu>0$. It is straightforward to derive $g(\nu)>0$ for large enough $n$ as in Lemma 3.5.
Lemma 3.5.
Suppose $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson($\lambda$) with $\lambda>0$. Then for each $\nu>0$, there exists a constant $c>0$, such that, with probability $1$, $g(\nu)>c$ for all large enough $n$.
For each $\nu>0$, Lemma 3.5 guarantees $g(\nu)>c$ for some $c>0$ and large enough $n$. Nevertheless, it does not guarantee that $g(\nu)>0$ for all $\nu>0$ and large enough $n$. To show that $g(\nu)>0$ for all $\nu\in(0,\nu_{\rm max}]$ and large enough $n$, we split $(0,\nu_{\rm max}]$ into $(0,\nu_{0}]$ and $(\nu_{0},\nu_{\rm max}]$, where $\nu_{0}$ is some positive number that is close enough to 00. Then we bound $g(\nu)$ from below for large enough $n$.
First, we find a lower bound for $g(\nu)$ on $\nu\in(0,\nu_{0}]$. Actually, according to (7), if $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson($\lambda$) with $\lambda>0$, then
$g(\nu)$
$=$ $\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right)$
$=$ $\frac{1}{n}\cdot\frac{\sum_{y=1}^{\infty}f_{y}}{\nu}+\frac{1}{n}\sum_{y\in I\setminus\{0,1\}}f_{y}\left(\frac{1}{\nu+1}+\cdots+\frac{1}{\nu+y-1}\right)-\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right)$
$\geq$ $\frac{1}{\nu}\cdot\frac{\#\{i:Y_{i}>0\}}{n}-\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right)$
$\rightarrow$ $\frac{1-e^{-\lambda}}{\nu}-\log\left(1+\frac{\lambda}{\nu}\right)$
almost surely, as $n$ goes to $\infty$.
Lemma 3.6.
Given $\lambda>0$ and $c>0$, there exists a $\nu_{0}>0$, such that,
$h(\nu,a,b):=\frac{a}{\nu}-\log\left(1+\frac{b}{\nu}\right)\geq c$
for all $0<\nu\leq\nu_{0}$, $a\in\left[\frac{1}{2}(1-e^{-\lambda}),\frac{3}{2}(1-e^{-\lambda})\right]$, and $b\in\left[\frac{\lambda}{2},\frac{3\lambda}{2}\right]$.
Note that $\nu_{0}$ in Lemma 3.6 does not depend on $Y_{i}$’s. Since #$\frac{\#\{i:Y_{i}>0\}}{n}\rightarrow 1-e^{-\lambda}$ and $\bar{Y}_{n}\rightarrow\lambda$ almost surely, as $n$ goes to $\infty$, there exists an event
$A=\{\omega\in\Omega\mid\lim_{n\rightarrow\infty}\frac{\#\{i:Y_{i}(\omega)>0\}}{n}=1-e^{-\lambda}\mbox{ and }\lim_{n\rightarrow\infty}\bar{Y}_{n}(\omega)=\lambda\}$ (17)
with probability $1$, where $\Omega$ is the sample space. Given $\omega\in A$, for large enough $n$, #$\frac{\#\{i:Y_{i}>0\}}{n}\in\left[\frac{1}{2}(1-e^{-\lambda}),\frac{3}{2}(1-e^{-\lambda})\right]$, $\bar{Y}_{n}\in\left[\frac{\lambda}{2},\frac{3\lambda}{2}\right]$, and thus $g(\nu)(\omega)\geq c$ for all $\nu\in(0,\nu_{0}]$. That is, with probability $1$, we obtain a common lower bound $c$ of $g(\nu)$, for all $\nu\in(0,\nu_{0}]$, with large enough $n$.
Secondly, we construct a lower bound for $g(\nu)$, $\nu\in(\nu_{0},\nu_{\rm max}]$, with large enough $n$, which relies on Lemma 3.7.
Lemma 3.7.
For any given $\nu_{0}>0$, there exists an $M>0$, which does not depend on the data, such that $|g^{\prime}(\nu)|\leq M$ for all $\nu>\nu_{0}$
With the aids of Theorem 3.3, Lemmas 3.5, 3.6, and 3.7, we are ready to derive the first major theoretical result of this paper as described in Theorem 3.4.
Theorem 3.4.
Suppose $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson($\lambda$) with $\lambda>0$, then for any $\nu_{\rm max}>0$, with probability 1, the MLE of $(\nu,p)$ for a negative binomial distribution under the constraint $0<\nu\leq\nu_{\rm max}$ is $\left(\nu_{\rm max},\frac{\nu_{\rm max}}{\nu_{\rm max}+\bar{Y}_{n}}\right)$ for all large enough $n$.
In practice, we may allow $\nu_{\rm max}$ to increase along with $n$. As a direct conclusion of the proof of Theorem 3.4, we have the following theorem.
Theorem 3.5.
Suppose $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson($\lambda$) with $\lambda>0$. Let $\nu_{\rm max}(n)$ be a sequence of positive real numbers such that $\nu_{\rm max}(n)\rightarrow\infty$ as $n$ goes to $\infty$. Denote $(\hat{\nu}_{n},\hat{p}_{n})$ as the MLE of $(\nu,p)$ for a negative binomial distribution under the constraint $0<\nu\leq\nu_{\rm max}(n)$. Then with probability $1$, $\hat{\nu}_{n}\rightarrow\infty$, as $n$ goes to $\infty$.
3.2 KS test statistic
Given $Y_{1},\ldots,Y_{n}$ iid $\sim{\rm Poisson}(\lambda)$, we denote $\hat{\bm{\theta}}=(\hat{\nu}_{n},\hat{p}_{n})^{T}$ as the MLE of $\bm{\theta}=(\nu,p)^{T}$ for a negative binomial distribution under the constraint $0<\nu\leq\nu_{\rm max}(n)$, where $\lim_{n\rightarrow\infty}\nu_{\rm max}(n)=\infty$. With the aid of Theorem 3.5, we obtain the following theorem for the limiting behavior of the KS test statistic $D_{n}$ for negative binomial distributions. Recall that for this KS test, the null hypothesis $H_{0}$ is that $Y_{1},\ldots,Y_{n}$ are iid from an NB distribution with unknown parameters, and the alternative hypothesis $H_{1}$ is that $Y_{1},\ldots,Y_{n}$ are iid from a Poisson distribution with unknown parameters.
Theorem 3.6.
Given $Y_{1},\ldots,Y_{n}$ iid $\sim{\rm Poisson}(\lambda)$ with $\lambda>0$, we denote $D_{n}=\sup_{y}|F_{n}(y)$ $-F_{\hat{\bm{\theta}}}(y)|$ as the KS test statistic for negative binomial distribution, where $F_{n}(y)$ is the EDF of the data, and $F_{\hat{\bm{\theta}}}(y)$ is the CDF of negative binomial distribution with the MLE $\hat{\bm{\theta}}$ obtained under the constraint $0<\nu\leq\nu_{\rm max}(n)$, where $\lim_{n\rightarrow\infty}\nu_{\rm max}(n)=\infty$. Then $D_{n}\rightarrow 0$ almost surely, as $n\rightarrow\infty$.
Theorem 3.6 theoretically justifies the phenomenon on $D_{n}$ observed in Table 1.
As a direct conclusion of Theorem 3.6 and the Glivenko-Cantelli theorem (see, for example, Theorem 7.5.2 in Resnick (1999)), we have the following corollary to conclude the convergence of the estimated NB distribution to the true Poisson($\lambda$) distribution.
Corollary 1.
Under the circumstance of Theorem 3.6, we must have $\sup_{y}|F_{\hat{\bm{\theta}}}(y)-F_{\lambda}(y)|\rightarrow 0$ almost surely, as $n\rightarrow\infty$.
4 Other Parameterizations for NB Distributions
In this section, we first review other parameterizations in the literature for negative binomial distributions, and derive the corresponding asymptotic properties of their MLEs as direct conclusions of the invariance property of MLE (Zehna, 1966; Pal and Berry, 1992). We further propose a new parameterization of negative binomial distributions, which extends the NB distributions to cover Poisson distributions as a special class. Using the new parameterization of NB distributions, the extended NB distribution with estimated parameters from a Poisson sample converges with probability 1 to the underlying Poisson distribution (Theorem 4.6).
4.1 Asymptotic results with other parameterizations in the literature
Anscombe (1949) and Bliss and Fisher (1953) adopted $\nu$ and $\mu=\nu(1-p)/p$ as the parameters for a negative binomial distribution with the PMF
$f_{\nu,\mu}(y)=\frac{\Gamma(\nu+y)}{\Gamma(\nu)\Gamma(y+1)}\left(\frac{\nu}{\nu+\mu}\right)^{\nu}\left(\frac{\mu}{\nu+\mu}\right)^{y},\>\>\>y=0,1,2,\ldots\ .$ (18)
Theorem 4.1.
Given nonnegative integers $Y_{1},\ldots,Y_{n}$, let $(\hat{\nu},\hat{\mu})$ denote an MLE of $(\nu,\mu)$ for the negative binomial distribution described as in (18) with constraints $\nu\in(0,\nu_{\rm max}]$ and $\mu\in[0,\infty)$. Then (i) $\hat{\mu}=0$ if and only if $\bar{Y}_{n}=0$; and (ii) if $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson($\lambda$) with $\lambda>0$, then with probability 1, $\hat{\nu}=\nu_{\rm max}$ and $\hat{\mu}=\bar{Y}_{n}$ for all large enough $n$.
Johnson et al. (2005) used $\nu$ and $P=(1-p)/p$ for a negative binomial distribution with the PMF
$f_{\nu,P}(y)=\frac{\Gamma(\nu+y)}{\Gamma(\nu)\Gamma(y+1)}\left(\frac{1}{1+P}\right)^{\nu}\left(\frac{P}{1+P}\right)^{y},\>\>\>y=0,1,2,\ldots\ .$ (19)
Theorem 4.2.
Given nonnegative integers $Y_{1},\ldots,Y_{n}$, let $(\hat{\nu},\hat{P})$ denote an MLE of $(\nu,P)$ for the negative binomial distribution described as in (19) with constraints $\nu\in(0,\nu_{\rm max}]$ and $P\in[0,\infty)$. Then (i) $\hat{P}=0$ if and only if $\bar{Y}_{n}=0$; and (ii) if $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson($\lambda$) with $\lambda>0$, then with probability 1, $\hat{\nu}=\nu_{\rm max}$ and $\hat{P}=\bar{Y}_{n}/\nu_{\rm max}$ for all large enough $n$.
Anscombe (1950) and Aldirawi and Yang (2022) used $\nu$ and $P=1-p$ for a negative binomial distribution with the PMF
$f_{\nu,P}(y)=\frac{\Gamma(\nu+y)}{\Gamma(\nu)\Gamma(y+1)}(1-P)^{\nu}P^{y},\>\>\>y=0,1,2,\ldots\ .$ (20)
Theorem 4.3.
Given nonnegative integers $Y_{1},\ldots,Y_{n}$, let $(\hat{\nu},\hat{P})$ denote an MLE of $(\nu,P)$ for the negative binomial distribution described as in (20) with constraints $\nu\in(0,\nu_{\rm max}]$ and $P\in[0,1)$. Then (i) $\hat{P}=0$ if and only if $\bar{Y}_{n}=0$; and (ii) if $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson($\lambda$) with $\lambda>0$, then with probability 1, $\hat{\nu}=\nu_{\rm max}$ and $\hat{P}=\bar{Y}_{n}/(\nu_{\rm max}+\bar{Y}_{n})$ for all large enough $n$.
So far all the four parameterizations for negative binomial distributions, namely (2), (18), (19), (20), adopted $\nu$, whose original meaning is the given number of successes in a sequence of indendent Bernoulli trials, while the negative binomial random variable is the observed number of failures before the $\nu$th success. Although we knew that a Poisson$(\lambda)$ distribution can be regarded as the limiting distribution as $\nu\to\infty$ and $p\to 1$ with $\nu(1-p)/p=\lambda$, the current parameterizations do not allow us to simply claim that $\hat{\nu}=\infty$, if the data come from a Poisson distribution.
4.2 A new parameterization and extended NB distributions
In this section, we propose a new parameterization, namely $\mu\in[0,\infty)$ and $p\in(0,1]$ for an extended negative binomial distribution. It includes both the degenerated case (Example 2.2) with $\mu=0$ and Poisson distributions with $p=1$. More specifically, (i) for $\mu>0$ and $p\in(0,1)$, its PMF is defined as
$f_{\mu,p}(y)=\frac{\Gamma\left(\frac{\mu p}{1-p}+y\right)}{\Gamma\left(\frac{\mu p}{1-p}\right)\Gamma(y+1)}p^{\frac{\mu p}{1-p}}(1-p)^{y},\>\>\>y=0,1,2,\ldots\ ;$ (21)
(ii) for $p\in(0,1)$, as the limit when $\mu\to 0^{+}$, its PMF is
$f_{0,p}(y)=\lim_{\mu\to 0^{+}}f_{\mu,p}(y)=\left\{\begin{array}[]{cl}1&\mbox{ if }y=0\\ 0&\mbox{ otherwise}\end{array}\right.\ ,$ (22)
namely, the degenerated case described in Example 2.2; (iii) for $\mu>0$, as the limit when $p\to 1^{-}$, its PMF is
$f_{\mu,1}(y)=\lim_{p\to 1^{-}}f_{\mu,p}(y)=\frac{\mu^{y}}{y!}e^{-\mu},\>\>\>y=0,1,2,\ldots\ ,$ (23)
namely, a Poisson$(\mu)$ distribution; and (iv) as the limit when $\mu\to 0^{+}$ and $p\to 1^{-}$, its PMF is
$f_{0,1}(y)=\lim_{\begin{smallmatrix}\mu\to 0^{+}\\ p\to 1^{-}\end{smallmatrix}}f_{\mu,p}(y)=\left\{\begin{array}[]{cl}1&\mbox{ if}y=0\\ 0&\mbox{ otherwise}\end{array}\right.\ ,$ (24)
namely, the degenerated case in Example 2.2 again.
Given a nonnegative-integer-valued dataset $\{Y_{1},\ldots,Y_{n}\}$, the log-likelihood function of $(\mu,p)\in[0,\infty)\times(0,1]$ is $l(\mu,p)=\sum_{i=1}^{n}\log f_{\mu,p}(Y_{i})$. Recall that the function $g(\nu)$ defined in (6) or (7) plays a central role in finding the MLE of $\nu$. It can be verified that for $\mu>0$ and $p\in(0,1)$,
$\frac{\partial l(\mu,p)}{\partial\mu}$ $=$ $\frac{np}{1-p}\left\{g\left(\frac{\mu p}{1-p}\right)+\log\left[p+\frac{\bar{Y}_{n}}{\mu}(1-p)\right]\right\}\ ,$ (25)
$\frac{\partial l(\mu,p)}{\partial p}$ $=$ $\frac{n\mu}{(1-p)^{2}}\left\{g\left(\frac{\mu p}{1-p}\right)+\log\left[p+\frac{\bar{Y}_{n}}{\mu}(1-p)\right]\right\}+\frac{n}{1-p}(\mu-\bar{Y}_{n})\ .$ (26)
According to the definition of $f_{\mu,p}(y)$ under different scenarios, we have the following lemma and theorem for an MLE $(\hat{\mu},\hat{p})\in[0,\infty)\times(0,1]$.
Lemma 4.1.
For each $y\in\{0,1,2,\ldots\}$, $f_{\mu,p}(y)$ as defined by equations (21), (22), (23), and (24) is a continuous function of $\mu\in[0,\infty)$ and $p\in(0,1]$, so does the log-likelihood $l(\mu,p)$.
Theorem 4.4.
There must exist a $\hat{\mu}\in[0,\infty)$ and a $\hat{p}\in(0,1]$ that maximizes $l(\mu,p)$ under the constraint $(\mu,p)\in[0,\infty)\times(0,1]$. If $\bar{Y}_{n}=0$, then we must have $\hat{\mu}=0$ while $\hat{p}$ can be any number in $(0,1]$; if $\bar{Y}_{n}>0$, then we must have $\hat{\mu}=\bar{Y}_{n}$
Theorem 4.5.
Given nonnegative integers $Y_{1},\ldots,Y_{n}$, let $(\hat{\mu},\hat{p})$ denote an MLE of $(\mu,p)$ for the extended negative binomial distribution described via (21), (22), (23), and (24) with parameters $\mu\in[0,\infty)$ and $p\in(0,1]$. Then (i) $\hat{\mu}=0$ if and only if $\bar{Y}_{n}=0$; and (ii) if $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson($\lambda$) with $\lambda>0$, then with probability 1, $\hat{\mu}=\bar{Y}_{n}$ and $\lim_{n\to\infty}\hat{p}=1$.
According to Theorem 4.5, by adopting the extended NB distribution with parameters $(\mu,p)\in[0,\infty)\times(0,1]$, if the data come from a Poisson distribution, then with probability 1, the MLE leads to the true distribution consistently. The major difference between the new parameterization and the old ones is that the new one contains Poisson distributions as a special class.
Similarly to Theorem 3.6 and Corollary 1, we have the following theorem for the extended negative binomial distribution described via (21), (22), (23), and (24) with parameters $\mu\in[0,\infty)$ and $p\in(0,1]$.
Theorem 4.6.
Given $Y_{1},\ldots,Y_{n}$ iid $\sim{\rm Poisson}(\lambda)$ with $\lambda>0$, we denote $D_{n}=\sup_{y}|F_{n}(y)$ $-F_{\hat{\mu},\hat{p}}(y)|$ as the KS test statistic for the extended negative binomial distribution, where $F_{\hat{\mu},\hat{p}}$ is the CDF of extended NB distribution with parameters equal to their MLEs $\hat{\mu},\hat{p}$. Then (i) $D_{n}\rightarrow 0$ almost surely, as $n\rightarrow\infty$; (ii) $\sup_{y}|F_{\hat{\mu},\hat{p}}(y)-F_{\lambda}(y)|\rightarrow 0$ almost surely, as $n\rightarrow\infty$.
Similarly to Algorithm 1, we propose the following Algorithm 2 to find the MLE $(\hat{\mu},\hat{p})$ for the extended negative binomial distribution, which finds $\hat{p}$ with the aid of $\hat{\nu}$.
Algorithm 2 Adaptive Profile Maximization Algorithm for Extended NB($\mu,p$)
Input: Data $Y_{1},\ldots,Y_{n}\in\{0,1,2,\ldots\}$, predetermined threshold values $\nu_{\rm max}>0$, $\varepsilon>0$, and $\delta\in(0,1)$ (e.g., $\nu_{\rm max}=10^{4}$, $\varepsilon=10^{-3}$, $\delta=0.1$)
Output: MLE $(\hat{\mu},\hat{p})$ for NB($\mu,p$)
1:Calculate the sample mean $\bar{Y}_{n}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}$ and the sample variance $S^{2}=\frac{1}{n-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}$;
2:If $\bar{Y}_{n}=0$, go to Step $6$ with $\hat{\mu}=0$ and $\hat{p}=1$; if $\bar{Y}_{n}\geq\frac{n-1}{n}S^{2}$, go to Step $6$ with $\hat{\mu}=\bar{Y}_{n}$ and $\hat{p}=1$; otherwise calculate the initial value $\hat{\nu}^{(0)}$ of $\nu$ according to (9);
3:Obtain $\nu_{*}$ by maximizing $h(\nu)$ with $h^{\prime}(\nu)=ng(\nu)$ and the starting value $\hat{\nu}^{(0)}$. If the ratio of distinct values in the sample $|I|/n<\delta$, use (5) and (7) for computing $h(\nu)$ and $g(\nu)$, respectively; otherwise, use (5) and (6) instead.
4:If $h(\nu_{\rm max})>h(\nu_{*})$ (see (5) or (5)), we let $\hat{\nu}=\nu_{\rm max}$; otherwise, $\hat{\nu}=\nu_{*}$;
5:Calculate $\hat{p}=\frac{\hat{\nu}}{\hat{\nu}+\bar{Y}_{n}}$;
6:Report $\hat{\mu}=\bar{Y}_{n}$ and $\hat{p}$.
5 Numerical Studies
In this section, we use simulation studies to numerically confirm our findings.
5.1 Comparison study on finding the MLE of NB distribution
In this section, we compare the proposed Adaptive Profile Maximization Algorithm (APMA, Algorithm 1) with seven other approaches that are currently available for finding the MLE of a negative binomial distribution, including the R function new.mle in package AZIAD (version 0.0.2, Dousti Mousavi et al. (2023)) recently developed for 27 probability distributions including NB, the R function fitdistr in package MASS (version 7.3-61, Venables and Ripley (2002)), a commonly used approach for finding MLE in R, and the R function fitdist in package fitdistrplus (version 1.2-1, Delignette-Muller and Dutang (2015)) supporting a variety of classical optimization methods, namely “Nelder-Mead” (Nelder and Mead, 1965), “BFGS” (a quasi-Newton method), “CG” (a conjugate gradients method), “L-BFGS-B” (BFGS with box constraints), “SANN” (simulated annealing), and “Brent”. We skip “Brent” in our comparison study since it is restricted to one-dimensional problems and not feasible for NB distributions.
For this comprehensive simulation study, we consider five pairs of parameters $(\nu,p)$ for negative binomial distributions, where $\nu=0.01$ and $p=0.99,0.9,0.5,0.1,0.01$ (in Section S2 of the Supplementary Material, we consider $\nu=0.1,1,10,100$ as well, the conclusions are similar). For each pair of $(\nu,p)$, we simulate $B=100$ NB$(\nu,p)$ random samples of size $n=100$ or $1000$, respectively. For each random sample, we calculate the MLE of $(\nu,p)$ using APMA (Algorithm 1), AZIAD, MASS (fitdistr), as well as Nelder-Mead, BFGS, CG, L-BFGS-B, and SANN by utilizing fitdist in R package fitdistrplus.
Table 3 shows the relative frequency of errors, namely the error rate, reported by each method over 100 simulations. The results show that only APMA and AZIAD are applicable in all combinations of $\nu$ and $p$, while the other five methods struggle to return with estimates, especially when $p=0.99$ or $0.9$. A similar phenomenon occurs even with a large sample size $n=1,000$. Further analysis (not shown here) indicates that the difficulty arises for the last six methods when the sample mean is close to the sample variance.
Table 3: Error rate of calculating the MLE of NB$(\nu,p)$ over 100 simulations
Error rate of calculating the MLE of NB$(\nu=0.01,p)$ with $n=100$
$\bm{p}$ APMA AZIAD MASS Nelder-Mead BFGS CG L-BFGS-B SANN
0.99 0 0 0.99 0.99 0.99 0.99 0.99 0.99
0.9 0 0 0.86 0.86 0.86 0.86 0.86 0.86
0.5 0 0 0.46 0.46 0.46 0.62 0.58 0.46
0.1 0 0 0.10 0.10 0.09 0.77 0.81 0.09
0.01 0 0 0.02 0.02 0.02 0.25 0.97 0.02
Error rate of calculating the MLE of NB$(\nu=0.01,p)$ with $n=1000$
$\bm{p}$ APMA AZIAD MASS Nelder-Mead BFGS CG L-BFGS-B SANN
0.99 0 0 0.99 0.99 0.99 0.99 1 1
0.9 0 0 0.71 0.71 0.74 0.71 0.97 0.83
0.5 0 0 0 0 0.03 0 0.29 0.03
0.1 0 0 0.01 0.01 0 0.74 1 0
0.01 0 0 0 0 0 0.03 1 0
Note: For example, the error rate “0.86” indicates that the corresponding algorithm fails to find the MLE in 86 out of 100 simulations.
To compare the efficiency or accuracy of the algorithms, we adopt the maximum likelihood ratio, that is, $L(\hat{\nu},\hat{p})/L(\hat{\nu}_{\rm APMA},\hat{p}_{\rm APMA})=\exp\{l(\hat{\nu},\hat{p})-l(\hat{\nu}_{\rm APMA},\hat{p}_{\rm APMA})\}$ when available. Table 4 reports the average ratios over the simulations when $(\hat{\nu},\hat{p})$ is available, which demonstrate that the last six methods do not achieve higher maximum likelihood on average when they return parameter estimates.
Table 4: Average ratio of maximum likelihood over 100 simulations with $\nu=0.01$
Average ratio of maximum likelihood with $n=100$
$\bm{p}$ APMA AZIAD MASS Nelder-Mead BFGS CG L-BFGS-B SANN
0.99 1 1 1 1 1 1 1 1
0.9 1 1 1 1 1 1 1 1
0.5 1 1 0.73 0.73 1 0.85 0.82 0.69
0.1 1 1 0.85 0.85 1 0.62 0.58 0.62
0.01 1 1 0.96 0.96 1 0.93 0.38 0.90
Average ratio of maximum likelihood with $n=1000$
$\bm{p}$ APMA AZIAD MASS Nelder-Mead BFGS CG L-BFGS-B SANN
0.99 1 1 1 1 1 1 NA NA
0.9 1 1 0.69 0.69 1 0.69 0.35 0.59
0.5 1 1 0.15 0.15 1 0.15 0.05 0.13
0.1 1 1 0.99 0.99 1 0.46 NA 0.87
0.01 1 1 0.99 0.99 0.91 0.99 NA 0.97
Note: (1) The ratio is calculated by “maximum likelihood by the algorithm when available” over “maximum likelihood by APMA”; (2) “NA” indicates that the corresponding algorithm fails to find the MLE in all 100 simulations.
In terms of computational time (see Table 5), our APMA (Algorithm 1) demonstrates a significant advantage compared to all other methods. Although AZIAD does as well as APMA in terms of error rates and average ratio of maximum likelihood, it costs much more computational time than APMA. Overall, we recommend our APMA (Algoirthm 1), especially when $p$ is close to $1$.
Table 5: Average computational time (sec) over 100 simulations with $\nu=0.01$, $n=1000$
$\bm{p}$ APMA AZIAD MASS Nelder-Mead BFGS CG L-BFGS-B SANN
0.99 $<0.0001$ 0.0048 0.0005 0.0006 0.0111 0.0002 0.0014 2.2171
0.9 0.0039 0.0582 0.0034 0.0089 0.0324 0.0164 0.0128 2.4634
0.5 0.0041 0.0368 0.0096 0.0132 0.0389 0.0488 0.0295 1.3228
0.1 0.0011 0.0044 0.0209 0.0255 0.0150 0.0776 0.0028 1.1146
0.01 0.0005 0.0234 0.0435 0.0490 0.0366 0.0873 0.0086 2.4210
5.2 Further comparison between APMA and AZIAD
From Table 4, we observe that APMA and AZIAD outperform others in terms of average maximum likelihood ratios, when the data are simulated from negative binomial distributions. In this section, we further compare the efficiency of APMA and AZIAD when they encounter Poisson samples.
Table 6: Average ratio of maximum likelihood obtained by AZIAD against APMA on $100$ random samples from Poisson($\lambda$) of size $n$
n
$\bm{\lambda}$ 50 500 5000 50000
1 0.995 0.979 0.884 0.461
3 0.990 0.948 0.763 0.178
5 0.984 0.934 0.655 0.062
10 0.976 0.886 0.471 0.004
From Table 6, we can see that APMA outperforms AZIAD in terms of finding higher maximum likelihood when the data come from a Poisson distribution. The advantage of APMA over AZIAD is greater as the sample size $n$ or the sample mean becomes more pronounced.
6 Discussion
In this paper, we investigate the asymptotic behaviors of the MLE of NB distribution and the resulting distribution when the data actually come from a Poisson distribution. As a solution, we provide a more efficient algorithm for finding the MLE of NB distributions and a new parameterization of NB distributions to include Poisson distributions as a special class.
In Table 1 of the Introduction section, we use the KS test to illustrate that a statistical test may have a poor power to detect that the data actually come from a Poisson distribution, not an NB distribution.
[…]
Since $\hat{\omega}(n)^{2}\leq\sup_{y}|F_{n}(y)-F_{\hat{\bm{\theta}}}(y)|^{2}=D_{n}^{2}$, then $D_{n}\rightarrow 0$ almost surely implies $\hat{\omega}(n)^{2}\rightarrow 0$ almost surely. We expect that the Cramer-von-Mises test may have a poor power for testing if the data come from an NB distribution ($H_{0}$) or a Poisson distribution ($H_{1}$) as well.
Supplementary Material
S1 Proofs
Proof of Theorem 2.2: If $M=0$, as mentioned in Example 2.2, $l(\nu,p)=n\nu\log p\leq 0$. To maximize $l(\nu,p)$ with $\nu\in(0,\nu_{\rm max}]$ and $p\in(0,1]$, we must have $\hat{p}=1$. In this case, $\hat{\nu}$ can be any positive number in $(0,\nu_{\rm max}]$.
On the other hand, if $M>0$, then $\bar{Y}_{n}>0$. For any fixed $\nu>0$, the best $p$ must take the form of $\nu/(\nu+\bar{Y}_{n})$ as in (3). Then $\hat{p}\leq\nu_{\rm max}/(\nu_{\rm max}+\bar{Y}_{n})<1$. $\Box$
Proof of Lemma 2.1: Since $\bar{Y}_{n}>0$, then there exists a $y\in I\setminus\{0\}$, such that, $f_{y}>0$, or equivalently, $f_{0}<n$. It can be verified that
$\nu g(\nu)=\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}+\frac{\nu}{n}\sum_{y\in I\setminus\{0,1\}}f_{y}\left(\frac{1}{\nu+1}+\cdots+\frac{1}{\nu+y-1}\right)-\nu\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right).$
Then $\lim_{\nu\to 0^{+}}\nu g(\nu)=\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}=1-\frac{f_{0}}{n}>0$. As a direct conclusion, $\lim_{\nu\rightarrow 0^{+}}g(\nu)=\infty$. Similarly, since
$\nu^{2}g^{\prime}(\nu)=-\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}-\frac{\nu^{2}}{n}\sum_{y\in I\setminus\{0,1\}}f_{y}\left[\frac{1}{(\nu+1)^{2}}+\cdots+\frac{1}{(\nu+y-1)^{2}}\right]+\frac{\nu\bar{Y}_{n}}{\nu+\bar{Y}_{n}},$
and thus $\lim_{\nu\rightarrow 0^{+}}\nu^{2}g^{\prime}(\nu)=-\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}=\frac{f_{0}}{n}-1<0$, then $\lim_{\nu\rightarrow 0^{+}}g^{\prime}(\nu)=-\infty$.
On the other hand, as $\nu\rightarrow\infty$, by applying Taylor’s theorem with the Peano form of the reminder (see, for example, Section 7.9 in Apostol (1967)), we have
$\frac{1}{\nu+k}$ $=$ $\frac{\frac{1}{\nu}}{1+\frac{k}{\nu}}=\frac{1}{\nu}-\frac{k}{\nu^{2}}+o(\nu^{-2})$
$\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right)$ $=$ $\frac{\bar{Y}_{n}}{\nu}-\frac{\bar{Y}_{n}^{2}}{2\nu^{2}}+o(\nu^{-2})$
$\frac{1}{(\nu+k)^{2}}$ $=$ $\frac{\frac{1}{\nu^{2}}}{\left(1+\frac{k}{\nu}\right)^{2}}=\frac{1}{\nu^{2}}-\frac{2k}{\nu^{3}}+o(\nu^{-3})$
$\frac{\bar{Y}_{n}}{\nu(\nu+\bar{Y}_{n})}$ $=$ $\frac{\bar{Y}_{n}}{\nu^{2}}-\frac{\bar{Y}_{n}^{2}}{\nu^{3}}+o(\nu^{-3})$
Then, as $\nu\rightarrow\infty$, by letting $S^{2}_{n}=n^{-1}\sum_{i=1}^{n}(Y_{i}-\bar{Y}_{n})^{2}=n^{-1}\sum_{i=1}^{n}Y_{i}^{2}-\bar{Y}_{n}^{2}$,
$g(\nu)$ $=$ $\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\log\left(1+\frac{\bar{Y}_{n}}{\nu}\right)$
$=$ $\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\left[\frac{y}{\nu}-\frac{y(y-1)}{2\nu^{2}}\right]-\frac{\bar{Y}_{n}}{\nu}+\frac{\bar{Y}_{n}^{2}}{2\nu^{2}}+o(\nu^{-2})$
$=$ $\frac{1}{n}\sum_{y\in I\setminus\{0\}}\frac{yf_{y}}{\nu}-\frac{1}{n}\sum_{y\in I\setminus\{0\}}\frac{y^{2}f_{y}}{2\nu^{2}}+\frac{1}{n}\sum_{y\in I\setminus\{0\}}\frac{yf_{y}}{2\nu^{2}}-\frac{\bar{Y}_{n}}{\nu}+\frac{\bar{Y}_{n}^{2}}{2\nu^{2}}+o(\nu^{-2})$
$=$ $\frac{\bar{Y}_{n}}{\nu}-\frac{1}{2\nu^{2}}\left(S^{2}_{n}+\bar{Y}_{n}^{2}\right)+\frac{\bar{Y}_{n}}{2\nu^{2}}-\frac{\bar{Y}_{n}}{\nu}+\frac{\bar{Y}_{n}^{2}}{2\nu^{2}}+o(\nu^{-2})$
$=$ $\frac{1}{2\nu^{2}}\left(\bar{Y}_{n}-S^{2}_{n}\right)+o(\nu^{-2}),$
$g^{\prime}(\nu)$ $=$ $-\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\left[\frac{1}{\nu^{2}}+\cdots+\frac{1}{(\nu+y-1)^{2}}\right]+\frac{\bar{Y}_{n}}{\nu(\nu+\bar{Y}_{n})}$
$=$ $-\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\left[\frac{y}{\nu^{2}}-\frac{y(y-1)}{\nu^{3}}\right]+\frac{\bar{Y}_{n}}{\nu^{2}}-\frac{\bar{Y}_{n}^{2}}{\nu^{3}}+o(\nu^{-3})$
$=$ $-\frac{1}{\nu^{2}}\sum_{y\in I\setminus\{0\}}\frac{yf_{y}}{n}+\frac{1}{\nu^{3}}\sum_{y\in I\setminus\{0\}}\frac{y^{2}f_{y}}{n}-\frac{1}{\nu^{3}}\sum_{y\in I\setminus\{0\}}\frac{yf_{y}}{n}+\frac{\bar{Y}_{n}}{\nu^{2}}-\frac{\bar{Y}_{n}^{2}}{\nu^{3}}+o(\nu^{-3})$
$=$ $-\frac{\bar{Y}_{n}}{\nu^{2}}+\frac{1}{\nu^{3}}\left(S^{2}_{n}+\bar{Y}_{n}^{2}\right)-\frac{\bar{Y}_{n}}{\nu^{3}}+\frac{\bar{Y}_{n}}{\nu^{2}}-\frac{\bar{Y}_{n}^{2}}{\nu^{3}}+o(\nu^{-3})$
$=$ $\frac{1}{\nu^{3}}\left(S^{2}_{n}-\bar{Y}_{n}\right)+o(\nu^{-3}).$
In other words, We obtain $\lim_{\nu\rightarrow\infty}{\nu}^{2}g(\nu)=\frac{1}{2}\left(\bar{Y}_{n}-S^{2}_{n}\right)$ and $\lim_{\nu\rightarrow\infty}{\nu}^{3}g^{\prime}(\nu)=S^{2}_{n}-\bar{Y}_{n}$, which imply $\lim_{\nu\rightarrow\infty}g(\nu)=0$ and $\lim_{\nu\rightarrow\infty}g^{\prime}(\nu)=0$, respectively. $\Box$
Proof of Theorem 2.3: Let $\nu_{*}={\rm argmin}_{\nu\in(0,\nu_{\rm max}]}g^{2}(\nu)$, then $g(\nu_{*})\neq 0$. According to Lemma 2.1, $\lim_{\nu\rightarrow 0^{+}}g(\nu)=\infty$. With the constraint $\nu\in(0,\nu_{\rm max}]$, we must have $\nu_{*}>0$.
We claim that $g(\nu_{*})>0$. Actually, if $g(\nu_{*})<0$, then there exists a $\nu_{**}\in(0,\nu_{*})$, such that, $g(\nu_{**})=0$, which violates $\min_{\nu\in(0,\nu_{\rm max}]}g^{2}(\nu)>0$.
The fact $g(\nu_{*})>0$ implies $g(\nu)\geq g(\nu_{*})>0$ for all $\nu\in(0,\nu_{\rm max}]$. That is, $h(\nu)$ increases all the time and thus the MLE of $\nu$ is $\nu_{\rm max}$. $\Box$
Proof of Lemma 3.1: For any $\nu_{0}>0$, once $\nu>\nu_{0}/2$,
$\left|\sum_{y=1}^{\infty}P(Y_{1}=y)\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\sum_{y=1}^{\infty}P(Y_{1}=y)\left(\frac{1}{\nu_{0}}+\cdots+\frac{1}{\nu_{0}+y-1}\right)\right|$
$\leq$ $\sum_{y=1}^{\infty}P(Y_{1}=y)\left[\left|\frac{1}{\nu}-\frac{1}{\nu_{0}}\right|+\cdots+\left|\frac{1}{\nu+y-1}-\frac{1}{\nu_{0}+y-1}\right|\right]$
$=$ $\sum_{y=1}^{\infty}P(Y_{1}=y)\left[\frac{|\nu-\nu_{0}|}{\nu\nu_{0}}+\cdots+\frac{|\nu-\nu_{0}|}{(\nu+y-1)(\nu_{0}+y-1)}\right]$
$\leq$ $\sum_{y=1}^{\infty}P(Y_{1}=y)\cdot y\cdot\frac{|\nu-\nu_{0}|}{\nu_{0}^{2}/2}$
$=$ $\frac{2E(Y_{1})}{\nu_{0}^{2}}\cdot|\nu-\nu_{0}|$
$\rightarrow$ 00 0
as $\nu$ goes to $\nu_{0}$ Since $\log(1+\mu/\nu)$ is a continuous function of $\nu$, $G_{F}(\nu)$ is a continuous function of $\nu$ on $(0,\infty)$, according to (11).
Furthermore,
$G_{F}(\nu)$ $=$ $\sum_{y=1}^{\infty}P(Y_{1}=y)\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\log\left(1+\frac{\mu}{\nu}\right)$
$=$ $\sum_{y=1}^{\infty}P(Y_{1}=y)\sum_{k=0}^{y-1}\frac{1}{\nu+k}-\log\left(1+\frac{\mu}{\nu}\right)$
$=$ $\sum_{k=0}^{\infty}\sum_{y=k+1}^{\infty}\frac{P(Y_{1}=y)}{\nu+k}-\log\left(1+\frac{\mu}{\nu}\right)$
$=$ $\sum_{k=0}^{\infty}\frac{1-P(Y_{1}\leq k)}{\nu+k}-\log\left(1+\frac{\mu}{\nu}\right)$
$=$ $\sum_{k=0}^{\infty}\frac{1-F(k)}{\nu+k}-\log\left(1+\frac{\mu}{\nu}\right)$
Thus the conclusion follows. $\Box$
Proof of Theorem 3.1: According to Kolmogorov’s Strong Law of Large Numbers (see, for example, Theorem 7.5.1 in Resnick (1999)), since $E(|Y_{1}|)=E(Y_{1})=\mu<\infty$, then $\bar{Y}_{n}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}\rightarrow\mu$ almost surely, as $n\rightarrow\infty$.
Similarly, for each $y\in\{0,1,2,\ldots\}$, $E(|\mathds{1}_{\{Y_{1}=y\}}|)=P(Y_{1}=y)\leq 1<\infty$. Then
$\frac{f_{y}}{n}$ $=$ #$\frac{\#\{1\leq i\leq n\mid Y_{i}=y\}}{n}\quad=\quad\frac{1}{n}\sum_{i=1}^{n}\mathds{1}_{\{Y_{i}=y\}}$
$\longrightarrow$ $E\mathds{1}_{\{Y_{1}=y\}}=P(Y_{1}=y)$
almost surely as $n\rightarrow\infty$. Furthermore, for each integer $M>0$, $E(|Y_{1}\mathds{1}_{\{Y_{1}>M\}}|)=E(Y_{1}\mathds{1}_{\{Y_{1}>M\}})\leq E(Y_{1})=\mu<\infty$. Then
$\sum_{y=M+1}^{\infty}\frac{yf_{y}}{n}$ $=$ $\sum_{y=M+1}^{\infty}\frac{\#\{1\leq i\leq n\mid Y_{i}=y\}\cdot y}{n}\quad=\quad\frac{1}{n}\sum_{i=1}^{n}Y_{i}\mathds{1}_{\{Y_{i}>M\}}$
$\longrightarrow$ $E(Y_{1}\mathds{1}_{\{Y_{1}>M\}})$
almost surely as $n\rightarrow\infty$. Then there exists an event $A$ with probability $1$, such that, for each $\omega\in A$,
$\frac{f_{y}(\omega)}{n}=\frac{\#\{1\leq i\leq n\mid Y_{i}(\omega)=y\}}{n}\longrightarrow P(Y_{1}=y)$ (S.1)
for each $y\in\{0,1,2,\ldots,\}$, $\bar{Y}_{n}(\omega)\rightarrow\mu$ and thus $\log\left(1+\frac{\bar{Y}_{n}(\omega)}{\nu}\right)\rightarrow\log\left(1+\frac{\mu}{\nu}\right)$, and
$\sum_{y=M+1}^{\infty}\frac{yf_{y}(\omega)}{n}\rightarrow E(Y_{1}\mathds{1}_{\{Y_{1}>M\}})$ (S.2)
for each $M>0$ as the convergence of a sequence of real numbers.
Now we only need to show that
$\sum_{y=1}^{\infty}\frac{f_{y}(\omega)}{n}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)\longrightarrow\sum_{y=1}^{\infty}P(Y_{1}=y)\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)$ (S.3)
for each $\omega\in A$. That is, we want to show that, for any $\epsilon>0$, there exists an $N$, such that,
$\left|\sum_{y=1}^{\infty}\frac{f_{y}(\omega)}{n}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\sum_{y=1}^{\infty}P(Y_{1}=y)\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)\right|<\epsilon$ (S.4)
for all $n>N$.
First, there exists an $M>0$, such that,
$\sum_{y=M+1}^{\infty}P(Y_{1}=y)\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)<\frac{\epsilon}{3}$ (S.5)
and
$E(Y_{1}\mathds{1}_{\{Y_{1}>M\}})<\frac{\nu\epsilon}{6}$
Then for each $y=1,\ldots,M$, there exists an $N_{y}>0$, such that,
$\left|\frac{f_{y}(\omega)}{n}-P(Y_{1}=y)\right|<\frac{\epsilon}{3}\cdot\frac{\nu}{M}\cdot\frac{1}{2^{y}}$
as long as $n>N_{y}$. Then for each $n>\max\{N_{1},\ldots,N_{M}\}$,
$\left|\sum_{y=1}^{M}\frac{f_{y}(\omega)}{n}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)-\sum_{y=1}^{M}P(Y_{1}=y)\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)\right|$ (S.6)
$\leq$ $\sum_{y=1}^{M}\left|\frac{f_{y}(\omega)}{n}-P(Y_{1}=y)\right|\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)$
$<$ $\sum_{y=1}^{M}\frac{\epsilon}{3}\cdot\frac{\nu}{M}\cdot\frac{1}{2^{y}}\cdot\frac{M}{\nu}$
$<$ $\frac{\epsilon}{3}$
We still need to show that
$\sum_{y=M+1}^{\infty}\frac{f_{y}(\omega)}{n}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)<\frac{\epsilon}{3}$
for large enough $n$. Actually, there exists an $N_{\nu}>0$ such that
$\left|\sum_{y=M+1}^{\infty}\frac{yf_{y}(\omega)}{n}-E(Y_{1}\mathds{1}_{\{Y_{1}>M\}})\right|<\frac{\nu\epsilon}{6}$
for all $n>N_{\nu}$. Then
$\sum_{y=M+1}^{\infty}\frac{f_{y}(\omega)}{n}\left(\frac{1}{\nu}+\cdots+\frac{1}{\nu+y-1}\right)$ (S.7)
$<$ $\sum_{y=M+1}^{\infty}\frac{f_{y}(\omega)}{n}\frac{y}{\nu}$
$<$ $\frac{1}{\nu}\left(E\left(Y_{1}{\mathbf{1}}_{Y_{1}>M}\right)+\frac{\nu\epsilon}{6}\right)$
$<$ $\frac{1}{\nu}\left(\frac{\nu\epsilon}{6}+\frac{\nu\epsilon}{6}\right)$
$=$ $\frac{\epsilon}{3}$
Combining (S.5), (S.6), and (S.7), we obtain that (S.4) is true for all $n>\max\{N_{1},\ldots,N_{M},$ $N_{\nu}\}$. That is, (S.3) is true for all $\omega\in A$. Combining with $\log\left(1+\frac{\bar{Y}_{n}(\omega)}{\nu}\right)\rightarrow\log\left(1+\frac{\mu}{\nu}\right)$ for all $\omega\in A$, we obtain $g(\nu)\rightarrow G_{F}(\nu)$ almost surely. $\Box$
Proof of Lemma 3.2: According to, for example, Example 3 in Aldirawi and Yang (2022), $E\Psi(\nu+Y_{1})=\Psi(\nu)-\log(p)$, if $Y_{1}\sim{\rm NB}(\nu,p)$ (please note that the parameters in their paper are different from ours, with $p$ be replaced with $1-p$). Then according to (10),
$G_{NB(\nu,p)}(\nu)$ $=$ $E\Psi(\nu+Y_{1})-\Psi(\nu)-\log\left(1+\frac{\mu}{\nu}\right)$
$=$ $\Psi(\nu)-\log(p)-\Psi(\nu)-\log\left(1+\frac{1-p}{p}\right)$
$=$ 00 0
where $\mu=\nu(1-p)/p$ for NB$(\nu,p)$. In other words,
$G_{NB(\nu,p)}(\nu)=\sum_{y=0}^{\infty}\frac{1-F_{NB(\nu,p)}(y)}{\nu+y}-\log\left(1+\frac{\mu}{\nu}\right)=0,$
where $F_{NB(\nu,p)}(y)=P(Y_{1}\leq y)$ with $Y_{1}\sim{\rm NB}(\nu,p)$. $\Box$
Proof of Lemma 3.3: Plugging in (15) and (16), we have for $x>0$,
$r(x)$ $=$ $\log\Gamma(\nu+x)-x\log(\nu+\lambda)-\log\Gamma(\nu)+\lambda-\nu\log\left(1+\frac{\lambda}{\nu}\right)$
$r^{\prime}(x)$ $=$ $\Psi(\nu+x)-\log(\nu+\lambda)$
According to, for example, Abramowitz and Stegun (1970) or Alzer and Jameson (2017), $\Psi(x)$ is strictly increasing and differentiable on $(0,\infty)$. Therefore, $r^{\prime}(x)$ is strictly increasing and differentiable on $(0,\infty)$.
According to Alzer (1997), $\log x-\frac{1}{x}<\Psi(x)<\log x-\frac{1}{2x}$ for all $x>0$. Then for $0<x<\lambda$, $r^{\prime}(x)<\Psi(\nu+\lambda)-\log(\nu+\lambda)<0$, and $\lim_{x\rightarrow\infty}r^{\prime}(x)=\infty$. The conclusion about $r^{\prime}(x)$ follows, with $x_{*}=\Psi^{-1}\left(\log(\nu+\lambda)\right)-\nu$ well defined and $x_{*}>\lambda$.
Since $x>\log(1+x)$ for all $x>0$, then $r(0)=\lambda-\nu\log\left(1+\frac{\lambda}{\nu}\right)>\lambda-\nu\cdot\frac{\lambda}{\nu}=0$.
On the other hand, according to Lemma 1 of Minc and Sathre (1964) or Alzer (1997),
$0<\log\Gamma(x)-\left[(x-\frac{1}{2})\log x-x+\frac{1}{2}\log(2\pi)\right]<\frac{1}{x}$
for all $x>1$. It can be verified that $\lim_{x\rightarrow\infty}\frac{r(x)}{x\log(\nu+x)}=1$, which implies $\lim_{x\rightarrow\infty}r(x)=\infty$. $\Box$
Proof of Theorem 3.2: According to Lemma 3.3, $r(x)$ attains its minimum at $x_{*}>\lambda$.
We first show that $r(x_{*})<0$. Otherwise, if $r(x_{*})\geq 0$, then $r(x)\geq 0$ for all $x\geq 0$ and $r(x)>0$ for all $x\neq x_{*}$. Then equivalently, we have $f_{NB(\nu,p)}(x)\geq f_{\lambda}(x)$ for all $x\geq 0$ and $f_{NB(\nu,p)}(x)>f_{\lambda}(x)$ for all $x\neq x_{*}$, which implies $\sum_{y=0}^{\infty}f_{NB(\nu,p)}(y)>\sum_{y=0}^{\infty}f_{\lambda}(y)$. However, $\sum_{y=0}^{\infty}f_{NB(\nu,p)}(y)=\sum_{y=0}^{\infty}f_{\lambda}(y)=1$. The contradiction implies that we must have $r(x_{*})<0$.
Note that the sign of $r(y)$ is exactly the same as the sign of $d(y)$ for all $y\in\{0,1,2,\ldots,\}$. We let $K_{1}$ be the largest possible integer such that $d(y)\geq 0$ for all $0\leq y\leq K_{1}$, and $K_{2}$ be the smallest possible integer such that $d(y)\geq 0$ for all $y\geq K_{2}$ Note that we must have $K_{2}-K_{1}\geq 2$ and $d(y)<0$ for each $K_{1}<y<K_{2}$ (otherwise, we have the contradiction $\sum_{y=0}^{\infty}f_{NB(\nu,p)}(y)>\sum_{y=0}^{\infty}f_{\lambda}(y)$ again).
According to Lemma 3.3, $r(0)>0$ and thus $d(0)=D(0)>0$. Since $d(y)>0$ for $0\leq y<K_{1}$ and $y>K_{2}$, then $D(y)$ strictly increases on $0\leq y<K_{1}$ and $y>K_{2}$; while $d(y)<0$ for $K_{1}<y<K_{2}$ implies that $D(y)$ strictly decreases on $K_{1}<y<K_{2}$ Since $\lim_{y\rightarrow\infty}D(y)=1-1=0$, then $D(y)$ must first increase and attain its maximum $D(K_{1})>0$, then strictly decrease and attain its minimum $D(K_{2})<0$, and then strictly increase after $K_{2}$ and approach to 00 as $y$ goes to $\infty$. We let $K_{*}$ be the largest possible integer such that $D(y)\geq 0$. Then we must have $K_{1}<K_{*}<K_{2}$, $D(y)>0$ for all $0\leq y<K_{*}$, and $D(y)<0$ for all $y>K_{*}$Box
[…]
$\sum_{y=0}^{\infty}[1-F(y)]$ $=$ $\sum_{y=0}^{\infty}\sum_{k=y+1}^{\infty}P(Y=k)=\sum_{k=1}^{\infty}\sum_{y=0}^{k-1}P(Y=k)$
$=$ $\sum_{k=1}^{\infty}k\cdot P(Y=k)$
$=$ $E(Y)$
Since $p=\nu/(\nu+\lambda)$, the means of NB$(\nu,p)$ and Poisson$(\lambda)$ are the same. That is, $\sum_{y=0}^{\infty}[1-F_{NB(\nu,p)}(y)]=\sum_{y=0}^{\infty}[1-F_{\lambda}(y)]=\lambda$. Then
$\sum_{y=0}^{\infty}D(y)=\sum_{y=0}^{\infty}[F_{NB(\nu,p)}(y)-F_{\lambda}(y)]=\sum_{y=0}^{\infty}[1-F_{\lambda}(y)]-\sum_{y=0}^{\infty}[1-F_{NB(\nu,p)}(y)]=0$
$\Box$
Proof of Theorem 3.3: Given $\lambda>0$ and $\nu>0$, we let $p=\nu/(\nu+\lambda)\in(0,1)$. Then the mean of NB$(\nu,p)$ is $\mu=\nu(1-p)/p=\nu\cdot\lambda/\nu=\lambda$. According to Lemmas 3.1 and 3.2, $G_{NB(\nu,p)}(\nu)=\sum_{y=0}^{\infty}\frac{1-F_{NB(\nu,p)}(y)}{\nu+y}-\log\left(1+\frac{\lambda}{\nu}\right)=0$. Since $G_{\lambda}(\nu)=\sum_{y=0}^{\infty}\frac{1-F_{\lambda}(y)}{\nu+y}-\log\left(1+\frac{\lambda}{\nu}\right)$, we only need to show that $\sum_{y=0}^{\infty}\frac{1-F_{\lambda}(y)}{\nu+y}>\sum_{y=0}^{\infty}\frac{1-F_{NB(\nu,p)}(y)}{\nu+y}$.
According to Theorem 3.2 and Lemma 3.4, we have $D(y)\geq 0$ for $0\leq y\leq K_{*}$, $D(y)<0$ for $y>K_{*}$, and $\sum_{y=0}^{K_{*}}D(y)=-\sum_{y=K_{*}+1}^{\infty}D(y)>0$. Then
$\sum_{y=0}^{\infty}\frac{1-F_{\lambda}(y)}{\nu+y}-\sum_{y=0}^{\infty}\frac{1-F_{NB(\nu,p)}(y)}{\nu+y}$ $=$ $\sum_{y=0}^{\infty}\frac{F_{NB(\nu,p)}(y)-F_{\lambda}(y)}{\nu+y}=\sum_{y=0}^{\infty}\frac{D(y)}{\nu+y}$
$=$ $\sum_{y=0}^{K_{*}}\frac{D(y)}{\nu+y}+\sum_{y=K_{*}+1}^{\infty}\frac{D(y)}{\nu+y}$
$>$ $\sum_{y=0}^{K_{*}}\frac{D(y)}{\nu+K_{*}}+\sum_{y=K_{*}+1}^{\infty}\frac{D(y)}{\nu+K_{*}+1}$
$=$ $\sum_{y=0}^{K_{*}}D(y)\cdot\left(\frac{1}{\nu+K_{*}}-\frac{1}{\nu+K_{*}+1}\right)$
$=$ $\sum_{y=0}^{K_{*}}D(y)\cdot\frac{1}{(\nu+K_{*})(\nu+K_{*}+1)}$
$>$ 00 0
$\Box$
Proof of Lemma 3.5: According to Theorem 3.1, if $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson$(\lambda)$ with $\lambda>0$, then there exists an event $A$ with probability $1$, such that, for each $\omega\in A$, $g(\nu)\rightarrow G_{\lambda}(\nu)>0$, as $n$ goes to $\infty$. We let $c=G_{\lambda}(\nu)/2>0$. Then for each $\omega\in A$, there exists an $N(\omega)>0$, such that, $|g(\nu)-G_{\lambda}(\nu)|<c$ and thus $g(\nu)>G_{\lambda}(\nu)-c=c$, for all $n\geq N(\omega)$. $\Box$
Proof of Lemma 3.6: First of all, for any $a\in\left[\frac{1}{2}(1-e^{-\lambda}),\frac{3}{2}(1-e^{-\lambda})\right]$, $b\in\left[\frac{\lambda}{2},\frac{3\lambda}{2}\right]$,
$h(\nu,a,b)\geq\frac{1-e^{-\lambda}}{2\nu}-\log\left(1+\frac{3\lambda}{2\nu}\right)$
Since $\lim_{\nu\rightarrow 0^{+}}\frac{1-e^{-\lambda}}{2\nu}=\infty$ and
$\lim_{\nu\rightarrow 0^{+}}\frac{\log\left(1+\frac{3\lambda}{2\nu}\right)}{\frac{1-e^{-\lambda}}{2\nu}}=\lim_{\nu\rightarrow 0^{+}}\frac{\frac{3\lambda}{2}}{(1+\frac{3\lambda}{2\nu})\frac{1-e^{-\lambda}}{2}}=0$
then $\lim_{\nu\rightarrow 0^{+}}\frac{1-e^{-\lambda}}{2\nu}-\log\left(1+\frac{3\lambda}{2\nu}\right)=\infty$. Therefore, there exists a $\nu_{0}>0$, such that, $\frac{1-e^{-\lambda}}{2\nu}-\log\left(1+\frac{3\lambda}{2\nu}\right)>c$ for all $0<\nu\leq\nu_{0}$, which leads to the conclusion. $\Box$
Proof of Lemma 3.7: According to (8), for each $\nu>\nu_{0}$,
$|g^{\prime}(\nu)|$ $=$ $\left|-\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\left[\frac{1}{\nu^{2}}+\cdots+\frac{1}{(\nu+y-1)^{2}}\right]+\frac{1}{\nu}-\frac{1}{\nu+\bar{Y}_{n}}\right|$
$\leq$ $\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\left[\frac{1}{\nu^{2}}+\cdots+\frac{1}{(\nu+y-1)^{2}}\right]+\frac{2}{\nu_{0}}$
$\leq$ $\frac{1}{n}\sum_{y\in I\setminus\{0\}}f_{y}\cdot\sum_{k=0}^{\infty}\frac{1}{(\nu+k)^{2}}+\frac{2}{\nu_{0}}$
$\leq$ $1\cdot\sum_{k=0}^{\infty}\frac{1}{(\nu_{0}+k)^{2}}+\frac{2}{\nu_{0}}$
$=$ $\frac{1}{{\nu_{0}}^{2}}+\sum_{k=1}^{\infty}\frac{1}{(\nu_{0}+k)^{2}}+\frac{2}{\nu_{0}}$
$\leq$ $\frac{1}{{\nu_{0}}^{2}}+\sum_{k=1}^{\infty}\frac{1}{k^{2}}+\frac{2}{\nu_{0}}$
$=$ $\frac{1}{{\nu_{0}}^{2}}+\frac{\pi}{6}+\frac{2}{\nu_{0}}$
Let $M=\frac{1}{{\nu_{0}}^{2}}+\frac{2}{\nu_{0}}+\frac{\pi}{6}>0$. Then $|g^{\prime}(\nu)|\leq M<\infty$ for all $\nu>\nu_{0}$Box
Proof of Theorem 3.4: Given $\lambda>0$ and $\nu_{\rm max}>0$, we set $c=1$ in Lemma 3.6 and obtain $\nu_{0}\in(0,\nu_{\rm max})$ as in Lemma 3.6 and event $A$ as defined in (17). Then for each $\omega\in A$, let $N_{0}(\omega)$ be a large enough integer, such that for all $n\geq N_{0}(\omega)$, #$\frac{\#\{i:Y_{i}(\omega)>0\}}{n}\in\left[\frac{1}{2}(1-e^{-\lambda}),\frac{3}{2}(1-e^{-\lambda})\right]$, $\bar{Y}_{n}(\omega)\in\left[\frac{\lambda}{2},\frac{3\lambda}{2}\right]$, and thus $g(\nu)(\omega)\geq 1$ for all $0<\nu\leq\nu_{0}$.
According to Lemma 3.1, $G_{\lambda}(\nu)$ is continuous on $(0,\infty)$. According to Theorem 3.3, $G_{\lambda}(\nu)>0$ for each $\nu>0$. Therefore, for the given $0<\nu_{0}<\nu_{\rm max}<\infty$, there exists a $c\in(0,1)$ such that $G_{\lambda}(\nu)>c$ for all $\nu\in[\nu_{0},\nu_{max}]$. On the other hand, according to Lemma 3.7, there exists an $M>0$, which does not depend on $\omega\in\Omega$, such that $|g^{\prime}(\nu)|\leq M<\infty$ for all $\nu>\nu_{0}$ and all $\omega\in\Omega$.
We let $K=4(\nu_{\rm max}-\nu_{0})M/c+1$, then $(\nu_{\rm max}-\nu_{0})M/K<c/4$. We denote the grid points, $\nu_{i}=\nu_{0}+(\nu_{\rm max}-\nu_{0})\cdot i/K$, for $i=1,\ldots,K$. According to Theorem 3.1, for each $i$, $g(\nu_{i})\rightarrow G_{\lambda}(\nu_{i})$ almost surely. We denote
$A_{K}=\{\omega\in A\mid\lim_{n\rightarrow\infty}g(\nu_{i})(\omega)=G(\nu_{i}),i=1,\ldots,K\}$
which has probability $1$. For each $\omega\in A_{K}$, there exist $N_{i}(\omega)$, $i=1,\ldots,K$, such that, $g(\nu_{i})(\omega)\geq\frac{c}{2}$ for all $n\geq N_{i}(\omega)$. We denote $N(\omega)=\max\{N_{0}(\omega),N_{1}(\omega),\ldots,N_{K}(\omega)\}$.
For each $\nu\in(\nu_{0},\nu_{\rm max}]$, there exists a $\nu_{i}$, such that, $\nu\in(\nu_{i-1},\nu_{i}]$. Then
$|g(\nu)(\omega)-g(\nu_{i})(\omega)|\leq M\cdot\frac{\nu_{\rm max}-\nu_{0}}{K}<M\cdot\frac{c}{4M}=\frac{c}{4}$
which implies $g(\nu)(\omega)>g(\nu_{i})(\omega)-\frac{c}{4}\geq\frac{c}{2}-\frac{c}{4}=\frac{c}{4}$, for all $n\geq N(\omega)$.
After all, for each $\omega\in A_{K}$ and each $n\geq N(\omega)$, $g(\nu)(\omega)>\frac{c}{4}>0$ for all $0<\nu\leq\nu_{\rm max}$ According to Theorem 2.3, under the constraint $\nu\in(0,\nu_{\rm max}]$, the MLE of $\nu$ for a negative binomial distribution is $\hat{\nu}(\omega)=\nu_{\rm max}$ The MLE of $p$ is via (3). $\Box$
Proof of Theorem 3.5: According to the proof of Theorem 3.4, for each positive integer $k$, there exists an event $A_{k}$ with probability $1$, such that, for each $\omega\in A_{k}$, there exist an integer $N(\omega,k)$ and a lower bound $c_{k}>0$, such that, $g(\nu)>c_{k}>0$ for all $n>N(\omega,k)$ and all $\nu\in(0,\nu_{\rm max}(k)]$.
We let $A_{0}=\cap_{k=1}^{\infty}A_{k}$ Then $P(A_{0})=1$ as well. Since $\nu_{\rm max}(n)\rightarrow\infty$ as $n$ goes to infinity, then for any $M>0$, there exists an integer $K_{M}>0$, such that, $\nu_{\rm max}(k)>M$ for all $k\geq K_{M}$ For each $\omega\in A_{0}$, given any $M>0$, we let $N_{M}=\max_{1\leq k\leq N_{M}}N(\omega,k)$. Then $g(\nu)(\omega)>\min_{1\leq k\leq K_{M}}c_{k}>0$, for all $\nu\in(0,M]$ and all $n>N_{M}(\omega)$, which implies that the MLE $\hat{\nu}_{n}(\omega)>M$. That is, $\hat{\nu}_{n}(\omega)\rightarrow\infty$, as $n$ goes to infinity. $\Box$
Proof of Theorem 3.6: We denote $F_{\lambda}(y)$ as the CDF of Poisson($\lambda$). For any given $\epsilon>0$, there exists an integer $M>0$, such that, $1-F_{\lambda}(y)\leq\frac{\epsilon}{6}$ for all $y\geq M$. Then
$D_{n}$ $=$ $\sup_{y}|F_{n}(y)-F_{\hat{\bm{\theta}}}(y)|$
$=$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)+F_{\lambda}(y)-F_{\hat{\bm{\theta}}}(y)\right|$
$\leq$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)\right|+\sup_{y<M}\left|F_{\lambda}(y)-F_{\hat{\bm{\theta}}}(y)\right|+\sup_{y\geq M}\left|F_{\lambda}(y)-F_{\hat{\bm{\theta}}}(y)\right|$
$=$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)\right|+\sup_{y<M}\left|F_{\lambda}(y)-F_{\hat{\bm{\theta}}}(y)\right|$
$+\ \sup_{y\geq M}\left|\left[1-F_{\lambda}(y)\right]-\left[1-F_{\hat{\bm{\theta}}}(y)\right]\right|$
$\leq$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)\right|+\sup_{y<M}\left|F_{\lambda}(y)-F_{\hat{\bm{\theta}}}(y)\right|$
$+\ \sup_{y\geq M}\left[1-F_{\lambda}(y)\right]+\sup_{y\geq M}\left[1-F_{\hat{\bm{\theta}}}(y)\right]$
$\leq$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)\right|+\sup_{y<M}\left|F_{\lambda}(y)-F_{\hat{\bm{\theta}}}(y)\right|+\left[1-F_{\lambda}(M)\right]+\left[1-F_{\hat{\bm{\theta}}}(M)\right]$
According to the Glivenko-Cantelli theorem (see, for example, Theorem 7.5.2 in Resnick (1999)), $\sup_{y}|F_{n}(y)-F_{\lambda}(y)|\rightarrow 0$ almost surely, as $n\rightarrow\infty$.
Let $A=\{\omega\in\Omega\mid\lim_{n\rightarrow\infty}\hat{\nu}_{n}(\omega)=\infty\mbox{ and }\lim_{n\rightarrow\infty}\sup_{y}|F_{n}(y)(\omega)-F_{\lambda}(y)|=0\}$. Then according to Theorem 3.5 and the Glivenko-Cantelli theorem, $P(A)=1$.
For each $\omega\in A$ and the given $\epsilon>0$, there exists a constant $N_{e}(\omega)>0$, such that, $\sup_{y}|F_{n}(y)(\omega)-F_{\lambda}(y)|\leq\frac{\epsilon}{3}$ for all $n\geq N_{e}(\omega)$.
According to Theorem 5.13.1 in Fisz (1963), $f_{{\rm NB}(\nu,p(\nu))}(k)\rightarrow f_{\lambda}(k)$ as $\nu\rightarrow\infty$ for each $k$, where $p(\nu)=\frac{\nu}{\nu+\lambda}$. Then, $F_{{\rm NB}(\nu,p(\nu))}(k)\rightarrow F_{\lambda}(k)$ as $\nu\rightarrow\infty$, for each $k$. For each $k=0,1,\ldots,M$, there exists a $V_{k}>0$ such that $|F_{{\rm NB}(\nu,p(\nu))}(k)-F_{\lambda}(k)|<\frac{\epsilon}{6}$ for all $\nu\geq V_{k}$. Let $V_{\rm max}=\max\{V_{0},V_{1},\ldots,V_{M}\}$. Then for all $\nu\geq V_{\rm max}$, $\sup_{k<M}|F_{{\rm NB}(\nu,p(\nu))}(k)-F_{\lambda}(k)|\leq\frac{\epsilon}{6}$ and $|F_{{\rm NB}(\nu,p(\nu))}(M)-F_{\lambda}(M)|\leq\frac{\epsilon}{6}$.
In addition, for each $k$,
$\lim_{n\rightarrow\infty}F_{\hat{\bm{\theta}}(\omega)}(k)=\lim_{n\rightarrow\infty}F_{{\rm NB}(\hat{\nu}_{n}(\omega),\hat{p}_{n}(\hat{\nu}_{n}(\omega)))}(k)=\lim_{\nu\rightarrow\infty}F_{{\rm NB}(\nu,p(\nu))}(k)=F_{\lambda}(k)<\infty$
There exists an $N_{k}(\omega)$ such that for all $n\geq N_{k}(\omega)$, $\hat{\nu}_{n}(\omega)>V_{\rm max}$, $|F_{\hat{\bm{\theta}}(\omega)}(k)-F_{\lambda}(k)|\leq\frac{\epsilon}{6}$. Then for each $n\geq N_{\rm max}(\omega)=\max\{N_{0}(\omega),N_{1}(\omega),\ldots,N_{M}(\omega),N_{e}(\omega)\}$, $\sup_{k<M}|F_{\hat{\bm{\theta}}(\omega)}(k)-F_{\lambda}(k)|\leq\frac{\epsilon}{6}$ and $0<1-F_{\hat{\bm{\theta}}(\omega)}(M)\leq 1-F_{\lambda}(M)+|F_{\hat{\bm{\theta}}(\omega)}(M)-F_{\lambda}(M)|\leq\frac{\epsilon}{6}+\frac{\epsilon}{6}=\frac{\epsilon}{3}$, which leads to $D_{n}(\omega)\leq\epsilon$.
Therefore, $D_{n}(\omega)$ goes to zero as $n$ goes to infinity. $\Box$
Proof of Corollary 1: Under the circumstance of Theorem 3.6, we have $\sup_{y}|F_{n}(y)-F_{\hat{\bm{\theta}}}(y)|\rightarrow 0$ almost surely, as $n\rightarrow\infty$. According to the Glivenko-Cantelli theorem (see, for example, Theorem 7.5.2 in Resnick (1999)), $\sup_{y}|F_{n}(y)-F_{\lambda}(y)|\rightarrow 0$ almost surely, as $n\rightarrow\infty$. We conclude because $|F_{\hat{\bm{\theta}}}(y)-F_{\lambda}(y)|\leq|F_{n}(y)-F_{\hat{\bm{\theta}}}(y)|+|F_{n}(y)-F_{\lambda}(y)|$. $\Box$
Proof of Theorem 4.1: If $\bar{Y}_{n}=0$, then $Y_{1}=\cdots=Y_{n}=0$ and the log-likelihood $l(\nu,\mu)=n\nu\log\frac{\nu}{\nu+\mu}\leq 0$, which is maximized only when $\mu=0$. That is, $\hat{\mu}=0$ while $\hat{\nu}$ can be any number in $(0,\nu_{\rm max}]$.
If $\bar{Y}_{n}>0$, then $Y_{i}>0$ for at least one $i$. It can be verified that $\lim_{\mu\to 0^{+}}l(\nu,\mu)=-\infty$ for any fixed $\nu>0$. Thus we must have $\hat{\mu}>0$. Actually, it can be verified that
$\frac{\partial l(\nu,\mu)}{\partial\mu}=\frac{n\nu}{\mu(\nu+\mu)}\cdot(\bar{Y}_{n}-\mu),$
which implies $\hat{\mu}=\bar{Y}_{n}>0$. Combining the cases of $\bar{Y}_{n}=0$ and $\bar{Y}_{n}>0$, we conclude that $\hat{\mu}=0$ if and only if $\bar{Y}_{n}=0$.
Recall the invariance property of MLE (see, e.g., Zehna (1966) or Pal and Berry (1992)), that is, given a function $\tau$ and an MLE $\hat{\bm{\theta}}$ of $\bm{\theta}$, $\tau(\hat{\bm{\theta}})$ is always an MLE of $\tau(\bm{\theta})$. In this case, $\mu=\nu(1-p)/p$ builds up a one-to-one correspondence between $\{(\nu,p)\mid\nu\in(0,\nu_{\rm max}],p\in(0,1)\}$ to $\{(\nu,\mu)\mid\nu\in(0,\nu_{\rm max}],\mu\in(0,\infty)\}$, which will pass along the uniqueness of MLE. According to the proof of Theorem 3.4, there exists an event $A$ of probability $1$, such that, for each $\omega\in A$, there exists an $N(\omega)>0$, $\bar{Y}_{n}(\omega)\geq\lambda/2>0$, $(\hat{\nu}=\nu_{\rm max},\ \hat{p}=\nu_{\rm max}/(\nu_{\rm max}+\bar{Y}_{n}))$ is the unique MLE of $(\nu,p)$ for all $n\geq N(\omega)$. Therefore, $(\hat{\nu}=\nu_{\rm max},\ \hat{\mu}=\hat{\nu}(1-\hat{p})/\hat{p}=\bar{Y}_{n})$ is the unique MLE of $(\nu,\mu)$ for all $n\geq N(\omega)$. $\Box$
Proof of Theorem 4.2: If $\bar{Y}_{n}=0$, then $Y_{1}=\cdots=Y_{n}=0$ and the log-likelihood $l(\nu,P)=-n\nu\log(1+P)\leq 0$, which is maximized only when $P=0$. That is, $\hat{P}=0$ while $\hat{\nu}$ can be any number in $(0,\nu_{\rm max}]$.
If $\bar{Y}_{n}>0$, it can be verified that $\lim_{P\to 0^{+}}l(\nu,P)=-\infty$ for any fixed $\nu>0$. Thus we must have $\hat{P}>0$. Actually, it can be verified that
$\frac{\partial l(\nu,P)}{\partial P}=\frac{n\nu}{P(1+P)}\left(\frac{\bar{Y}_{n}}{\nu}-P\right),$
which implies $\hat{P}=\bar{Y}_{n}/\hat{\nu}>0$. Combining the cases of $\bar{Y}_{n}=0$ and $\bar{Y}_{n}>0$, we conclude that $\hat{P}=0$ if and only if $\bar{Y}_{n}=0$.
Note that in this case, $P=(1-p)/p$ builds up a one-to-one correspondence between $\{(\nu,p)\mid\nu\in(0,\nu_{\rm max}],p\in(0,1)\}$ to $\{(\nu,P)\mid\nu\in(0,\nu_{\rm max}],P\in(0,\infty)\}$. According to the proof of Theorem 3.4, there exists an event $A$ of probability $1$, such that, for each $\omega\in A$, there exists an $N(\omega)>0$, $\bar{Y}_{n}(\omega)\geq\lambda/2>0$, $(\hat{\nu}=\nu_{\rm max},\ \hat{p}=\nu_{\rm max}/(\nu_{\rm max}+\bar{Y}_{n}))$ is the unique MLE of $(\nu,p)$ for all $n\geq N(\omega)$. Therefore, $(\hat{\nu}=\nu_{\rm max},\ \hat{P}=(1-\hat{p})/\hat{p}=\bar{Y}_{n}/\nu_{\rm max})$ is the unique MLE of $(\nu,P)$ for all $n\geq N(\omega)$. $\Box$
Proof of Theorem 4.3: If $\bar{Y}_{n}=0$, then $Y_{1}=\cdots=Y_{n}=0$ and the log-likelihood $l(\nu,P)=n\nu\log(1-P)\leq 0$, which is maximized only when $P=0$. That is, $\hat{P}=0$ while $\hat{\nu}$ can be any number in $(0,\nu_{\rm max}]$.
If $\bar{Y}_{n}>0$, it can be verified that $\lim_{P\to 0^{+}}l(\nu,P)=-\infty$ for any fixed $\nu>0$. Thus we must have $\hat{P}>0$. Actually, it can be verified that
$\frac{\partial l(\nu,P)}{\partial P}=\frac{n(\nu+\bar{Y}_{n})}{P(1-P)}\left(\frac{\bar{Y}_{n}}{\nu+\bar{Y}_{n}}-P\right),$
which implies $\hat{P}=\bar{Y}_{n}/(\hat{\nu}+\bar{Y}_{n})>0$. Combining the cases of $\bar{Y}_{n}=0$ and $\bar{Y}_{n}>0$, we conclude that $\hat{P}=0$ if and only if $\bar{Y}_{n}=0$.
In this case, $P=1-p)$ builds up a one-to-one correspondence between $\{(\nu,p)\mid\nu\in(0,\nu_{\rm max}],p\in(0,1)\}$ to $\{(\nu,P)\mid\nu\in(0,\nu_{\rm max}],P\in(0,1)\}$. According to the proof of Theorem 3.4, there exists an event $A$ of probability $1$, such that, for each $\omega\in A$, there exists an $N(\omega)>0$, $\bar{Y}_{n}(\omega)\geq\lambda/2>0$, $(\hat{\nu}=\nu_{\rm max},\ \hat{p}=\nu_{\rm max}/(\nu_{\rm max}+\bar{Y}_{n}))$ is the unique MLE of $(\nu,p)$ for all $n\geq N(\omega)$. Therefore, $(\hat{\nu}=\nu_{\rm max},\ \hat{P}=1-\hat{p}=\bar{Y}_{n}/(\nu_{\rm max}+\bar{Y}_{n})$ is the unique MLE of $(\nu,P)$ for all $n\geq N(\omega)$. $\Box$
Proof of Lemma 4.1: We only need to confirm the definitions in (22), (23), and (24) as the corresponding limits of $f_{\mu,p}(y)$ defined in (21) with more detail.
To verify (22), we fix a $p\in(0,1)$. According to (21), for $\mu>0$, $f_{\mu,p}(0)=p^{\frac{\mu p}{1-p}}$, which leads to $\lim_{\mu\to 0^{+}}f_{\mu,p}(0)=1$. For each $y>0$, $\lim_{\mu\to 0^{+}}f_{\mu,p}(y)=0$, as $\lim_{x\to 0^{+}}\Gamma(x)=\infty$.
To verify (23), we fix a $\mu\in(0,\infty)$. It can be verified that $\log f_{\mu,p}(0)=\frac{\mu p}{1-p}\log p=\frac{\mu p}{1-p}\log[1-(1-p)]$ goes to $-\mu$, as $p\to 1^{-}$. That is, $\lim_{p\to 1^{-}}f_{\mu,p}(0)=e^{-\mu}$. For each $y>0$, it can be verified that, as $p\to 1^{-}$,
$f_{\mu,p}(y)=\frac{1}{y!}\cdot\prod_{j=0}^{y-1}\left(\frac{\mu p}{1-p}+j\right)(1-p)^{y}\cdot p^{\frac{\mu p}{1-p}}\longrightarrow\frac{1}{y!}\cdot\mu^{y}\cdot e^{-\mu}\ .$
As for (24), when $\mu\to 0^{+}$ and $p\to 1^{-}$, it can be verified that $f_{\mu,p}(0)=p^{\frac{\mu p}{1-p}}=\exp\left\{\frac{\mu p}{1-p}\log[1-(1-p)]\right\}\to 1$. For each $y>0$, it can be verified that $f_{\mu,p}(y)=\frac{1}{y!}\cdot\prod_{j=0}^{y-1}[\mu p+j(1-p)]\cdot p^{\frac{\mu p}{1-p}}\to\frac{1}{y!}\cdot 0\cdot 1=0$. $\Box$
Proof of Theorem 4.4: Due to Lemma 4.1, we only need to show that the maximum of $l(\mu,p)$ (1) won’t occur when $\mu\to\infty$; and (2) can always occur at some $p\in(0,1]$.
Case (i): $\bar{Y}_{n}=0$, which implies $Y_{1}=\cdots=Y_{n}=0$. Fixing any $p\in(0,1)$, $l(\mu,p)=\frac{n\mu p}{1-p}\log p<0$ when $\mu>0$, and $l(0,p)\equiv 0$, which implies that the best $\mu$ is 00, while $p$ can be any number in $(0,1)$. Fixing $p=1$, $l(\mu,1)=-n\mu<0$ when $\mu>0$, and $l(0,1)=0$, which implies that the best $\mu$ is 00 again. Combining the cases of $p\in(0,1)$ and $p=1$, we conclude that $\bar{Y}_{n}=0$ implies that the best $\mu$ must be 00, while $p$ can be any number in $(0,1]$.
Case (ii): $\bar{Y}_{n}>0$, which implies that $Y_{i}>0$ for at least one $i$. Fixing any $p\in(0,1)$, it can be verified that $l(0,p)=\lim_{\mu\to 0^{+}}l(\mu,p)=-\infty$ and $\lim_{\mu\to\infty}l(\mu,p)=-\infty$, which imply that the best $\mu$ exists in $(0,\infty)$ in this case. When $p=1$, $l(0,1)=\lim_{\mu\to 0^{+}}l(\mu,1)=-\infty$ and $\lim_{\mu\to\infty}l(\mu,1)=-\infty$, which imply that the best $\mu$ exists in $(0,\infty)$ as well. On the other hand, fixing any $\mu\in(0,\infty)$, $\lim_{p\to 0^{+}}l(\mu,p)=-\infty$ as $\lim_{x\to 0^{+}}\Gamma(x)=\infty$. Then the best $p$ exists in $(0,1]$. As a conclusion, $\bar{Y}_{n}>0$ implies that an MLE $(\hat{\mu},\hat{p})$ must exist in $(0,\infty)\times(0,1]$. If $\hat{p}=1$, it can be verified that $l(\mu,1)$ is maximized at $\hat{\mu}=\bar{Y}_{n}$ only. If $\hat{p}<1$, then we must have $\partial l/\partial\mu=\partial l/\partial p=0$ at $(\hat{\mu},\hat{p})$. From (25) and (26), we obtain $\hat{\mu}=\bar{Y}_{n}$ as well. Combining the cases of $\hat{p}=1$ and $\hat{p}<1$, we always have $\hat{\mu}=\bar{Y}_{n}$Box
Proof of Theorem 4.5: As a direct conclusion of Theorem 4.4, $\hat{\mu}=0$ if and only if $\bar{Y_{n}}=0$.
Suppose $Y_{1},\ldots,Y_{n}$ are iid $\sim$ Poisson($\lambda$) with $\lambda>0$. Fixing an $c>0$, for any $M>0$, according to the proof of Theorem 3.4, there exists an event $A_{M}$ of probability $1$, such that, for any $\omega\in A_{M}$, (1) there exists an $N_{M}(\omega)>0$; (2) $\bar{Y}_{n}(\omega)\in\left[\frac{\lambda}{2},\ \frac{3}{2}\lambda\right]$ for all $n\geq N_{M}(\omega)$; (3) $g(\nu)(\omega)>c/4>0$ for all $0<\nu\leq M$ and all $n\geq N_{M}(\omega)$.
According to Theorem 4.4, we always have $\hat{\mu}=\bar{Y}_{n}$ For all $n\geq N_{M}(\omega)$, $\hat{\mu}(\omega)=\bar{Y}_{n}(\omega)\geq\lambda/2>0$, which implies
$\frac{\partial l}{\partial p}\left(\bar{Y}_{n}(\omega),p\right)=\frac{n\bar{Y}_{n}(\omega)}{(1-p)^{2}}\cdot g\left(\frac{\bar{Y}_{n}(\omega)p}{1-p}\right)>0$
for all $0<p\leq\frac{M}{M+\bar{Y}_{n}(\omega)}$ We then conclude $\hat{p}(\omega)\geq\frac{M}{M+\bar{Y}_{n}(\omega)}\geq\frac{M}{M+\frac{3}{2}\lambda}$ for all $n\geq N_{M}(\omega)$.
We let $A_{\infty}=\cap_{M=1}^{\infty}A_{M}$, which also has probability $1$. Then for each $\omega\in A_{\infty}$, $\hat{p}\to 1$, as $n\to\infty$. $\Box$
Proof of Theorem 4.6: Similarly to the proof of Theorem 3.6, we denote $F_{\lambda}(y)$ as the CDF of Poisson($\lambda$). For any given $\epsilon>0$, there exists an integer $M>0$, such that, $1-F_{\lambda}(y)\leq\frac{\epsilon}{6}$ for all $y\geq M$. Then
$D_{n}$ $=$ $\sup_{y}|F_{n}(y)-F_{\hat{\mu},\hat{p}}(y)|$
$=$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)+F_{\lambda}(y)-F_{\hat{\mu},\hat{p}}(y)\right|$
$\leq$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)\right|+\sup_{y<M}\left|F_{\lambda}(y)-F_{\hat{\mu},\hat{p}}(y)\right|+\sup_{y\geq M}\left|F_{\lambda}(y)-F_{\hat{\mu},\hat{p}}(y)\right|$
$=$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)\right|+\sup_{y<M}\left|F_{\lambda}(y)-F_{\hat{\mu},\hat{p}}(y)\right|$
$+\ \sup_{y\geq M}\left|\left[1-F_{\lambda}(y)\right]-\left[1-F_{\hat{\mu},\hat{p}}(y)\right]\right|$
$\leq$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)\right|+\sup_{y<M}\left|F_{\lambda}(y)-F_{\hat{\mu},\hat{p}}(y)\right|$
$+\ \sup_{y\geq M}\left[1-F_{\lambda}(y)\right]+\sup_{y\geq M}\left[1-F_{\hat{\mu},\hat{p}}(y)\right]$
$\leq$ $\sup_{y}\left|F_{n}(y)-F_{\lambda}(y)\right|+\sup_{y<M}\left|F_{\lambda}(y)-F_{\hat{\mu},\hat{p}}(y)\right|+\left[1-F_{\lambda}(M)\right]+\left[1-F_{\hat{\mu},\hat{p}}(M)\right]$
According to the Glivenko-Cantelli theorem (see, for example, Theorem 7.5.2 in Resnick (1999)), $\sup_{y}|F_{n}(y)-F_{\lambda}(y)|\rightarrow 0$ almost surely, as $n\rightarrow\infty$.
Let $A=\{\omega\in\Omega\mid\hat{\mu}_{n}(\omega)=\bar{Y}_{n}(\omega),\lim_{n\rightarrow\infty}\bar{Y}_{n}(\omega)=\lambda,\lim_{n\rightarrow\infty}\hat{p}_{n}(\omega)=1,\lim_{n\rightarrow\infty}\sup_{y}|F_{n}(y)(\omega)-F_{\lambda}(y)|=0\}$. Then according to Theorem 4.5, Kolmogorov’s Strong Law of Large Numbers, and the Glivenko-Cantelli theorem, $P(A)=1$.
For each $\omega\in A$ and the given $\epsilon>0$, there exists a constant $N_{e}(\omega)>0$, such that, $\sup_{y}|F_{n}(y)(\omega)-F_{\lambda}(y)|\leq\frac{\epsilon}{3}$ for all $n\geq N_{e}(\omega)$.
According to Theorem 5.13.1 in Fisz (1963), $f_{{\rm NB}(\nu,p(\nu))}(k)\rightarrow f_{\lambda}(k)$ as $\nu\rightarrow\infty$ for each $k$, where $p(\nu)=\frac{\nu}{\nu+\lambda}$. Then, $F_{{\rm NB}(\nu,p(\nu))}(k)\rightarrow F_{\lambda}(k)$ as $\nu\rightarrow\infty$, for each $k$. For each $k=0,1,\ldots,M$, there exists a $V_{k}>0$ such that $|F_{{\rm NB}(\nu,p(\nu))}(k)-F_{\lambda}(k)|<\frac{\epsilon}{6}$ for all $\nu\geq V_{k}$ Let $V_{\rm max}=\max\{V_{0},V_{1},\ldots,V_{M}\}$. Then for all $\nu\geq V_{\rm max}$, $\sup_{k<M}|F_{{\rm NB}(\nu,p(\nu))}(k)-F_{\lambda}(k)|\leq\frac{\epsilon}{6}$ and $|F_{{\rm NB}(\nu,p(\nu))}(M)-F_{\lambda}(M)|\leq\frac{\epsilon}{6}$.
In addition, for each $k\in\{0,1,\ldots,M\}$,
$\left|F_{\hat{\mu}_{n}(\omega),\hat{p}_{n}(\omega)}(k)-F_{\lambda}(k)\right|$ $=$ $\left\{\begin{array}[]{cl}\left|F_{{\rm NB}(\hat{\nu}_{n}(\omega),\hat{p}_{n}(\omega))}(k)-F_{\lambda}(k)\right|\ ,&\mbox{ if }\hat{p}_{n}(\omega)<1;\\ \left|F_{\hat{\lambda}=\bar{Y}_{n}(\omega)}(k)-F_{\lambda}(k)\right|\ ,&\mbox{if }\hat{p}_{n}(\omega)=1.\end{array}\right.$
$\longrightarrow$ $0\ ,$
as $n$ goes to infinity, where $\hat{\nu}_{n}(\omega)=\bar{Y}_{n}(\omega)\hat{p}_{n}(\omega)/[1-\hat{p}_{n}(\omega)]$. Then there exists an $N_{k}(\omega)$ such that for all $n\geq N_{k}(\omega)$, if we still have $\hat{p}_{n}(\omega)<1$, then we must have $\hat{\nu}_{n}(\omega)\geq V_{\rm max}$ and $|F_{{\rm NB}(\hat{\nu}_{n}(\omega),\hat{p}_{n}(\omega))}(k)-F_{\lambda}(k)|\leq\frac{\epsilon}{6}$; if $\hat{p}_{n}(\omega)=1$, we still have $\left|F_{\hat{\lambda}=\bar{Y}_{n}(\omega)}(k)-F_{\lambda}(k)\right|\leq\frac{\epsilon}{6}$ Then for each $n\geq N_{\rm max}(\omega)=\max\{N_{0}(\omega),N_{1}(\omega),\ldots,N_{M}(\omega),N_{e}(\omega)\}$, $\sup_{k<M}|F_{\hat{\mu}_{n}(\omega),\hat{p}_{n}(\omega)}(k)-F_{\lambda}(k)|\leq\frac{\epsilon}{6}$ and $0<1-F_{\hat{\mu}_{n}(\omega),\hat{p}_{n}(\omega)}(M)\leq 1-F_{\lambda}(M)+|F_{\hat{\mu}_{n}(\omega),\hat{p}_{n}(\omega)}(M)-F_{\lambda}(M)|\leq\frac{\epsilon}{6}+\frac{\epsilon}{6}=\frac{\epsilon}{3}$, which leads to $D_{n}(\omega)\leq\epsilon$.
Therefore, $D_{n}(\omega)$ goes to zero as $n$ goes to infinity.
Similarly to the proof for Corollary 1, we obtain $\sup_{y}|F_{\hat{\mu},\hat{p}}(y)-F_{\lambda}(y)|\rightarrow 0$ almost surely, as $n\rightarrow\infty$. $\Box$
S2 More tables for simulation studies
In this section, we provide more tables for supporting the conclusions described in Section 5.1. More specifically, we extend the comparison analysis by considering 25 parameter pairs $(\nu,p)$ for negative binomial distributions, including $\nu=0.01,0.1,1,10,100$ and $p=0.99,0.9,0.5,0.1,0.01$.
For each pair $(\nu,p)$, we generate $B=100$ random samples of size $n=100$ and $1000$, respectively. Table 7 and Table 8 display the error rate over 100 simulations for size $n=100$ and $1000$, respectively. Table 9 and Table 10 report the average computational time in seconds over 100 simulations for size $n=100$ and $1000$, respectively. Table 11 and Table 12 present the average ratio of maximum likelihood over 100 simulations for size $n=100$ and $1000$, respectively.
Table 7: Error rate of calculating the MLE of $NB(\nu,p)$ over 100 simulations with $n=100$
APMA AZIAD
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 0 0 0 0 0 0 0 0 0 0
0.1 0 0 0 0 0 0 0 0 0 0
1 0 0 0 0 0 0 0 0 0 0
10 0 0 0 0 0 0 0 0 0 0
100 0 0 0 0 0 0 0 0 0 0
MASS Nelder-Mead
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 0.99 0.86 0.46 0.10 0.02 0.99 0.86 0.46 0.09 0.02
0.1 0.95 0.34 0 0 0 0.95 0.34 0 0 0
1 0.35 0 0 0 0 0.35 0 0 0 0
10 0 0.23 0 0 0 0 0 0 0 0
100 0.56 0.20 0 0 0 0 0 0 0 0
BFGS CG
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 0.99 0.86 0.46 0.10 0.02 0.99 0.86 0.62 0.77 0.25
0.1 0.95 0.34 0 0 0 0.95 0.55 0.91 0.53 0
1 0.35 0 0 0 0 0.51 0.81 0.66 0 0
10 0 0.23 0 0 0 0.82 0.99 0.57 0 0
100 0.56 0.20 0 0 0 1 1 1 1 1
L-BFGS-B SANN
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 0.99 0.86 0.58 0.81 0.97 0.99 0.86 0.46 0.09 0.02
0.1 0.95 0.34 0.74 1 1 0.95 0.34 0 0 0
1 0.35 0.39 0.99 1 1 0.35 0 0 0 0
10 0.29 0.55 0 0 0 0 0 0 0 0
100 0.57 0.16 0 0 0 0 0 0 0 0
Table 8: Error rate of calculating the MLE of $NB(\nu,p)$ over 100 simulations with $n=1000$
APMA AZIAD
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 0 0 0 0 0 0 0 0 0 0
0.1 0 0 0 0 0 0 0 0 0 0
1 0 0 0 0 0 0 0 0 0 0
10 0 0 0 0 0 0 0 0 0 0
100 0 0 0 0 0 0 0 0 0 0
MASS Nelder-Mead
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 0.99 0.71 0 0.01 0 0.99 0.74 0.03 0 0
0.1 0.76 0 0 0 0 0.76 0 0 0 0
1 0 0 0 0 0 0 0 0 0 0
10 0 0 0 0 0 0 0 0 0 0
100 0.01 0.46 0 0 0 0 0 0 0 0
BFGS CG
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 0.99 0.71 0 0.01 0 0.99 0.71 0 0.74 0.03
0.1 0.76 0 0 0 0 0.77 0 0.98 0.22 0
1 0 0 0 0 0 0 0.87 0.18 0 0
10 0 0 0 0 0 0.40 0.99 0.29 0 0
100 0.01 0.46 0 0 0 1 1 1 1 1
L-BFGS-B SANN
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 1 0.97 0.29 1 1 1 0.83 0.03 0 0
0.1 1 0.18 0.99 1 1 0.97 0.01 0 0 0
1 0.12 0.28 1 1 1 0 0 0 0 0
10 0 0.19 0 0 0 0 0 0 0 0
100 0.40 0.03 0 0 0 0 0 0 0 0
Table 9: Average computational time (in secs) over 100 simulations with $n=100$
APMA AZIAD
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 0.0003 0.0002 0.0001 0.0002 $<0.0001$ 0.0008 0.0003 0.0019 0.0014 0.0015
0.1 0.0002 0.0005 0.0004 0.0004 0.0009 0.0003 0.0027 0.0012 0.0008 0.0007
1 0.0004 0.0004 0.0004 0.0001 0.0004 0.0032 0.0018 0.0007 0.0010 0.0007
10 0.0003 0.0005 0.0003 0.0002 0.0001 0.0027 0.0011 0.0006 0.0005 0.0007
100 0.0002 0.0004 0.0005 0.0001 0.0003 0.0022 0.0019 0.0006 0.0005 0.0007
MASS Nelder-Mead
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 $<0.0001$ 0.0004 0.0006 0.0027 0.0025 0.0034 0.0022 0.0025 0.0045 0.0038
0.1 0.0002 0.0004 0.0025 0.0028 0.0025 0.0036 0.0036 0.0036 0.0036 0.0039
1 0.0004 0.0019 0.0039 0.0029 0.0023 0.0030 0.0029 0.0043 0.0043 0.0041
10 0.0009 0.0042 0.0019 0.0012 0.0012 0.0035 0.0032 0.0029 0.0031 0.0031
100 0.0062 0.0055 0.0019 0.0017 0.0012 0.0034 0.0040 0.0030 0.0029 0.0030
BFGS CG
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 0.0008 0.0008 0.0019 0.0048 0.0041 0.0010 0.0016 0.0036 0.0113 0.0076
0.1 0.0015 0.0019 0.0039 0.0041 0.0040 0.0016 0.0083 0.0131 0.0150 0.0082
1 0.0014 0.0030 0.0054 0.0040 0.0039 0.0067 0.0122 0.0191 0.0107 0.0122
10 0.0020 0.0059 0.0033 0.0024 0.0025 0.0111 0.0193 0.0186 0.0050 0.0038
100 0.0083 0.0075 0.0031 0.0031 0.0031 0.0191 0.0165 0.0141 0.0148 0.0142
L-BFGS-B SANN
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 0.0010 0.0012 0.0016 0.0006 0.0019 0.2086 0.2478 0.2487 0.2295 0.2476
0.1 0.0012 0.0025 0.0012 0.0017 0.0006 0.2572 0.2481 0.2022 0.2258 0.2678
1 0.0017 0.0025 0.0019 0.0015 0.0018 0.2466 0.2022 0.2659 0.3644 0.3313
10 0.0023 0.0019 0.0042 0.0026 0.0026 0.2006 0.2617 0.3111 0.2819 0.2769
100 0.0023 0.0046 0.0035 0.0033 0.0031 0.2527 0.2389 0.2094 0.2114 0.2094
Table 10: Average computational time (in secs) over 100 simulations with $n=1000$
APMA AZIAD
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 $<0.0001$ 0.0039 0.0041 0.0011 0.0005 0.0048 0.0582 0.0368 0.0044 0.0234
0.1 0.0043 0.0064 0.0038 0.0058 0.0022 0.0622 0.0714 0.0173 0.0203 0.0062
1 0.0021 0.0011 0.0023 0.0031 0.0029 0.0352 0.0092 0.0063 0.0045 0.0046
10 0.0021 0.0015 0.0019 0.0013 0.0024 0.0217 0.0096 0.0027 0.0023 0.0036
100 0.0008 0.0001 0.0008 0.0003 0.0017 0.0203 0.0071 0.0042 0.0034 0.0053
MASS Nelder-Mead
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 0.0005 0.0034 0.0096 0.0209 0.0435 0.0111 0.0324 0.0389 0.0150 0.0366
0.1 0.0035 0.0099 0.0494 0.0479 0.0175 0.0333 0.0346 0.0374 0.0449 0.0193
1 0.0031 0.0049 0.0231 0.0247 0.0234 0.0127 0.0123 0.0229 0.0264 0.0214
10 0.0031 0.0134 0.0117 0.0094 0.0070 0.0114 0.0139 0.0149 0.0157 0.0151
100 0.0128 0.0350 0.0115 0.0083 0.0098 0.0169 0.0095 0.0133 0.0106 0.0128
BFGS CG
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 0.0006 0.0089 0.0132 0.0255 0.049 0.0002 0.0164 0.0488 0.0776 0.0873
0.1 0.0068 0.0119 0.0509 0.0500 0.0229 0.0142 0.0563 0.2376 0.2220 0.0568
1 0.0050 0.0058 0.0241 0.0269 0.0235 0.0179 0.0751 0.1288 0.0875 0.0882
10 0.0041 0.0145 0.0131 0.0101 0.0091 0.0448 0.1633 0.1425 0.0168 0.0163
100 0.0141 0.0423 0.0135 0.0116 0.0113 0.1548 0.0639 0.0936 0.0966 0.0936
L-BFGS-B SANN
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 0.0014 0.0128 0.0295 0.0028 0.0086 2.2171 2.4634 1.3228 1.1146 2.4210
0.1 0.0108 0.0303 0.0166 0.0094 0.0034 2.4467 2.5228 2.7569 1.4900 1.6591
1 0.0111 0.0055 0.0034 0.0050 0.0052 1.0991 1.0751 2.2402 3.1270 2.7250
10 0.0053 0.0113 0.0134 0.0086 0.0097 1.0688 2.0337 2.5173 2.1250 1.9603
100 0.0113 0.0030 0.0156 0.0136 0.0130 1.9328 1.4881 1.4335 1.4712 1.4626
Table 11: Average ratio of maximum likelihood over 100 simulations with $n=100$
APMA AZIAD
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 1 1 1 1 1 1 1 1 1 1
0.1 1 1 1 1 1 1 1 1 1 1
1 1 1 1 1 1 1 1 1 1 1
10 1 1 1 1 1 1 1 1 1 1
100 1 1 1 1 1 0.99 0.99 1 1 1
MASS Nelder-Mead
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 1 1 0.73 0.85 0.96 1 1 1 1 1
0.1 1 0.91 0.92 1 1 1 1 1 1 0.99
1 0.99 0.94 1 1 1 1 1 1 1 1
10 0.91 1 1 1 1 1 1 1 1 1
100 1 0.97 1 1 1 1 1 1 1 1
BFGS CG
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 1 1 0.73 0.85 0.96 1 1 0.85 0.62 0.93
0.1 1 0.91 0.92 1 1 1 0.88 0.90 1 1
1 0.99 0.94 1 1 1 0.98 0.99 1 1 1
10 0.91 1 1 1 1 0.98 1 1 1 1
100 1 0.97 1 1 1 NA NA NA NA NA
L-BFGS-B SANN
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 1 1 0.82 0.58 0.38 1 1 0.69 0.62 0.90
0.1 1 0.91 0.80 NA NA 1 0.91 0.45 0.97 0.99
1 0.99 0.95 1 NA NA 0.99 0.82 0.86 1 1
10 0.92 1 1 1 1 0.88 0.92 1 1 1
100 0.99 1 1 1 1 0.88 0.50 0.08 0.03 0.04
Note: “NA” indicates that the corresponding algorithm fails to find the MLE in all 100 simulations.
Table 12: Average ratio of maximum likelihood over 100 simulations with $n=1000$
APMA AZIAD
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 1 1 1 1 1 1 1 1 1 1
0.1 1 1 1 1 1 1 1 1 1 1
1 1 1 1 1 1 1 1 1 1 1
10 1 1 1 1 1 0.99 1 1 1 1
100 1 1 1 1 1 0.96 0.99 1 1 0.99
MASS Nelder-Mead
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 1 0.69 0.15 0.99 0.99 1 1 1 1 0.91
0.1 1 0.70 0.97 1 1 1 1 1 1 0.99
1 0.93 0.46 1 1 1 1 1 1 1 1
10 0.82 0.98 1 1 1 1 1 1 1 1
100 0.83 0.99 1 1 1 1 1 1 1 1
BFGS CG
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 1 0.69 0.15 0.99 0.99 1 0.69 0.15 0.46 0.99
0.1 1 0.70 0.97 1 1 1 0.70 1 1 1
1 0.93 0.46 1 1 1 0.93 0.93 1 1 1
10 0.82 0.98 1 1 1 0.94 1 1 1 1
100 0.83 0.99 1 1 1 NA NA NA NA NA
L-BFGS-B SANN
$\bm{\nu\ \backslash\ p}$ 0.99 0.9 0.5 0.1 0.01 0.99 0.9 0.5 0.1 0.01
0.01 NA 0.35 0.05 NA NA NA 0.59 0.13 0.87 0.97
0.1 NA 0.64 1 NA NA 1 0.70 0.41 1 0.99
1 0.94 0.47 NA NA NA 0.94 0.46 1 1 0.99
10 0.82 0.99 1 1 1 0.85 0.95 1 1 1
100 0.83 1 1 1 1 0.57 0.19 0.36 0.77 0.86
Note: “NA” indicates that the corresponding algorithm fails to find the MLE in all 100 simulations.